### v1

# Stochastic Approximation with Block Coordinate Optimal Stepsizes

Abstract
We consider stochastic approximation with block-coordinate stepsizes and propose adaptive stepsize rules that aim to minimize the expected distance from the next iterate to an optimal point. These stepsize rules employ online estimates of the second moment of the search direction along each block coordinate. The popular Adam algorithm can be interpreted as a particular heuristic for such estimation. By leveraging a simple conditional estimator, we derive a new method that obtains comparable performance as Adam but requires less memory and fewer hyper-parameters. We prove that this family of methods converges almost surely to a small neighborhood of the optimal point, and the radius of the neighborhood depends on the bias and variance of the second-moment estimator. Our analysis relies on a simple aiming condition that assumes neither convexity nor smoothness, thus has broad applicability.
1 Introduction
We consider unconstrained stochastic optimization problems of the form
$\mathop{\mbox{minimize}{}}_{x\in\mathbf{R}^{n}}~{}F(x):=\mathbf{E}_{\xi}[f(x,\xi)],$ (1)
where $x\in\mathbf{R}^{n}$ is the decision variable, $\xi$ is a random variable, and $f$ is the loss function. In the context of statistical machine learning (e.g., hastie2009elements, ), $x$ represents the parameters of a prediction model, $\xi$ represents randomly sampled data from some (unknown) probability distribution, and $f(x,\xi)$ is the error in making predictions about $\xi$ using the model parametrized by $x$. We assume that $F$ is bounded below and there exists an optimal solution $x_{*}$.
Suppose that for any pair $x$ and $\xi$, we can evaluate the gradient of $f$ with respect to $x$, denoted as $\nabla f(x,\xi)$. Starting with an initial point $x_{0}\in\mathbf{R}^{n}$, the classical stochastic approximation method RobbinsMonro51 generates a sequence $\{x_{1},x_{2},\ldots\}$ with the update rule
$x_{t+1}=x_{t}-\alpha_{t}\nabla f(x_{t},\xi_{t}),$ (2)
where $\alpha_{t}$ is the stepsize (often called the learning rate in the machine learning literature). The convergence properties of this method are well studied in the stochastic approximation literature (e.g., RobbinsMonro51, ; Wasan69book, ; KushnerYin03book, ). In particular, the stepsizes $\alpha_{t}$ should be nonnegative and unsummable, i.e., $\sum_{t=0}^{\infty}\alpha_{t}=\infty$. Stronger guarantees such as almost sure convergence also require $\alpha_{t}$ to be square-summable, i.e., $\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty$ (e.g., RobbinsMonro51, ; Blum1954multidim, ; chung1954, ; ROBBINS1971, ).
Despite the rich literature on their convergence theory, stochastic approximation methods in practice often require heuristics and trial and error in choosing the stepsize sequence $\{\alpha_{t}\}$. Adaptive rules that can adjust stepsizes on the fly have been developed in both the optimization literature (e.g., Kesten58, ; MirzoakhmedovUryasev83, ; RuszczynskiSyski83stats, ; RuszczynskiSyski84IFAC, ; DelyonJuditsky93, ) and by the machine learning community (e.g., Jacobs88DBD, ; Sutton92IDBD, ; Schraudolph1999, ; MahmoodSutton12TuningFree, ). More recently, adaptive algorithms that use coordinate-wise stepsizes have become very popular following the seminal works of AdaGrad DuchiHazanSinger2011adagrad and Adam KingmaBa2014adam , and they demonstrate significantly better performance in training large-scale deep neural networks. In this paper, we propose a family of stochastic approximation methods with block-coordinate stepsizes and present a general framework for their convergence analysis.
1.1 Stochastic approximation with block-coordinate stepsizes
We focus on stochastic approximation methods of the form
$x_{t+1}=x_{t}-s_{t}\odot d_{t},$ (3)
where $d_{t}\in\mathbf{R}^{n}$ is a stochastic search direction, $s_{t}\in\mathbf{R}^{n}$ is a vector of coordinate-wise stepsizes, and $\odot$ denotes element-wise product (Hadamard product) of two vectors. Other than using a different stepsize for each single coordinate, a block structure can be imposed where coordinates within each block share a common stepsize (see Section 1.3).
The two most common choices for the search direction are: the stochastic gradient, i.e., $d_{t}=\nabla f(x_{t},\xi_{t})$, and its exponential moving average (EMA). Letting $g_{t}=\nabla f(x_{t},\xi_{t})$, the EMA of stochastic gradient can be expressed as
$d_{t}=\beta_{1}d_{t-1}+(1-\beta_{1})g_{t},$ (4)
where $\beta_{1}\in[0,1)$ is a smoothing factor. This is often called stochastic momentum.
The Adam algorithm KingmaBa2014adam uses the stochastic momentum in (4) as its search direction and sets the coordinate-wise stepsizes as
$s_{t,i}=\frac{\alpha_{t}}{\sqrt{v_{t,i}}+\epsilon},\qquad i=1,\ldots,n,$ (5)
where $\alpha_{t}\in\mathbf{R}$ is a common stepsize schedule for all coordinates, $v_{t,i}$ is the EMA of $g_{t,i}^{2}$, and $\epsilon>0$ is a small constant to improve numerical stability when $v_{t,i}$ becomes very close to zero. Here, computing $v_{t,i}$ requires a second smoothing factor $\beta_{2}\in[0,1]$, i.e.,
$v_{t,i}=\beta_{2}v_{t-1,i}+(1-\beta_{2})g^{2}_{t,i},\qquad i=1,\ldots,n.$ (6)
For Adam to perform well, the second smoothing factor $\beta_{2}$ is often set to be larger than $\beta_{1}$. For example, a pair of commonly used choices is $\beta_{1}=0.9$ and $\beta_{2}=0.99$.
Adam KingmaBa2014adam and its variant AdamW Loshchilov2019AdamW (which implements decoupled weight decay) have been very successful in training large-scale deep neural networks. However, theoretical understanding of their superior performance is still incomplete despite many recent efforts (e.g., reddi2019convergence, ; bock2018improvement, ; balles2018dissecting, ; zou2019sufficient, ; defossez2020simple, ; Luo2022neurips, ; kunstner2024heavy, ). On the other hand, there are also many works that propose new variants or alternatives to Adam and AdamW, either starting from fundamental optimization principles (e.g., yao2021adahessian, ; gomes2024adafisher, ; hwang2024fadam, ; lin2024can, ; jordan2024muon, ; pethick2025LMOs, ) or based on empirical algorithmic search (e.g., chen2023symbolic, ; zhang2024adam, ). But all have limited success. Adam and especially AdamW are still the most widely used algorithms for training large deep learning models, and their effectiveness remains a myth.
1.2 Contributions and outline
In this paper, we propose a family of block-coordinate optimal stepsize (BCOS) rules for stochastic approximation. BCOS gives a novel interpretation of Adam and AdamW and provides their convergence analysis as special cases of a general framework. Moreover, we derive variants of BCOS that obtain competitive performance against Adam(W) but require less memory (optimizer states) and fewer hyper-parameters. More specifically:
  • In Section 2, we derive BCOS by minimizing the expected distance from the next iterate to an optimal point. While the optimal stepsizes cannot be computed exactly, we make several simplifications and approximate the second moments of the search direction with simple EMA estimators.
  • In Section 3, we instantiate BCOS with specific search directions. In particular, we show that RMSProp rmsprop2012lecture and Adam KingmaBa2014adam can be interpreted as special cases of BCOS. With the stochastic momentum as search direction, we leverage a simple conditional estimator to derive a variant that requires fewer optimizer states and hyper-parameters than Adam. Integrating with decoupled weight decay Loshchilov2019AdamW gives the BCOSW variants that enjoy comparable performance as AdamW.
  • In Section 4, we present numerical experiments to compare several variants of BCOS(W) with Adam(W) on several deep learning tasks. We observe that BCOSW performs better than BCOS, and BCOSW with the conditional estimator achieves top performance with fewer optimizer states and less hyper-parameter tuning.
  • In Section 5, we present the convergence analysis of BCOS(W) based on a simple aiming condition, which assumes neither convexity nor smoothness. We prove that the BCOS family of methods (including Adam and AdamW) converge almost surely to a small neighborhood of a target point, and the radius of the neighborhood depends on the bias and variance of the second-moment estimator.
  • We conclude the paper in Section 6.
Generality
The general form of stochastic approximation (e.g., RobbinsMonro51, ; Wasan69book, ; KushnerYin03book, ) concerns solving a nonlinear system of equations $G(x)=0$ where the map $G:\mathbf{R}^{n}\to\mathbf{R}^{n}$ only admits noisy evaluations in the form of $g(\cdot,\xi)$, where $\xi$ is a random variable, under the assumption that $G(x)=\mathbf{E}_{\xi}[g(x,\xi)]$. In this case, stochastic approximation takes the form
$x_{t+1}=x_{t}-\alpha_{t}\,g(x_{t},\xi_{t}).$ (7)
Apparently, the stochastic gradient method (2) is a special case with $g(x,\xi)=\nabla f(x,\xi)$. Although our main motivation is for solving the stochastic optimization problem (1) in the context of machine learning, the algorithms and convergence analysis developed in this paper apply to the general case of stochastic approximation with block-coordinate stepsizes.
1.3 Notations
Let $\mathcal{I}_{1},\ldots,\mathcal{I}_{m}$ be a non-overlapping partition of the index set $\{1,\ldots,n\}$, each with cardinality $n_{k}=|\mathcal{I}_{k}|$. Correspondingly, we partition the vectors $x_{t}$, $s_{t}$ and $d_{t}$ into blocks $x_{t,k}$, $s_{t,k}$ and $d_{t,k}$ in $\mathbf{R}^{n_{k}}$ for $k=1,\ldots,m$. We use a common stepsize $\gamma_{t,k}\in\mathbf{R}$ within each block, i.e., $s_{t,k}=\gamma_{t,k}\mathbf{1}_{n_{k}}$, where $\mathbf{1}_{n_{k}}$ denotes the vector of all ones of dimension $n_{k}$. As a result, the block-coordinate update in (3) can be written as
$x_{t+1,k}=x_{t,k}-s_{t,k}\odot d_{t,k}=x_{t,k}-\gamma_{t,k}d_{t,k},\qquad k=1,\ldots,m.$
Notice that $\gamma_{t,k}$ is always a scalar and $\gamma_{t}$ is a vector in $\mathbf{R}^{m}$ instead of $\mathbf{R}^{n}$ (unless $m=n$).
[…]
The $\textup{sign}(\cdot)$ function is defined as $\textup{sign}(\alpha)=1$ if $\alpha>0$, $-1$ if $\alpha<0$ and 00 if $\alpha=0$.
2 Derivation of BCOS
In this section, we first derive the ideal optimal stepsizes for block-coordinate update, which unfortunately is not computable in practice; then we make several simplifications and approximations to derive a practical algorithm.
2.1 Block-coordinate optimal stepsizes
We consider the distance of the next iterate $x_{t+1}$ to an optimal point $x_{*}$. Specifically, in terms of the squared Euclidean distance,
$\|x_{t+1}-x_{*}\|^{2}$ $=\|x_{t}-s_{t}\odot d_{t}-x_{*}\|^{2}$
$=\|x_{t}-x_{*}\|^{2}-2\langle x_{t}-x_{*},s_{t}\odot d_{t}\rangle+\|s_{t}\odot d_{t}\|^{2}.$
Exploiting the block partitions of $x_{t}$, $s_{t}$ and $d_{t}$ (Section 1.3) and $s_{t,k}=\gamma_{t,k}\mathbf{1}_{n_{k}}$, we obtain
$\|x_{t+1}-x_{*}\|^{2}$ $=\|x_{t}-x_{*}\|^{2}+\sum_{k=1}^{m}\left(-2\gamma_{t,k}\langle x_{t,k}-x_{*,k},\,d_{t,k}\rangle+\gamma_{t,k}^{2}\|d_{t,k}\|^{2}\right).$
Taking expectation conditioned on the realization of all random variables up to $x_{t}$, i.e.,
$\mathbf{E}_{t}[\cdot]:=\mathbf{E}[\cdot|x_{0},d_{0},x_{1},d_{1},\ldots,x_{t}],$ (8)
we have
$\mathbf{E}_{t}\!\left[\|x_{t+1}-x_{*}\|^{2}\right]=\|x_{t}-x_{*}\|^{2}+\sum_{k=1}^{m}\Bigl{(}-2\gamma_{t,k}\bigl{\langle}x_{t,k}-x_{*,k},\,\mathbf{E}_{t}[d_{t,k}]\bigr{\rangle}+\gamma_{t,k}^{2}\mathbf{E}_{t}\!\left[\|d_{t,k}\|^{2}\right]\Bigr{)}.$ (9)
In order to minimize the expected distance from $x_{t+1}$ to $x_{*}$, we can minimize the right-hand side of (9) over the stepsizes $\{\gamma_{t,k}\}_{k=1}^{m}$. This can be done separately for each block by minimizing a simple quadratic, resulting in the optimal stepsizes
$\widehat{\gamma}_{t,k}=\frac{\langle x_{t,k}-x_{*,k},\,\mathbf{E}_{t}[d_{t,k}]\rangle}{\mathbf{E}_{t}[\|d_{t,k}\|^{2}]},\qquad k=1,\ldots,m.$ (10)
Notice that these optimal stepsizes can be positive or negative, depending on the sign of the inner product in the numerator.
Plugging the optimal stepsizes into (9), we have
$\mathbf{E}_{t}\!\left[\|x_{t+1}-x_{*}\|^{2}\right]=\|x_{t}-x_{*}\|^{2}-\sum_{k=1}^{m}\frac{\left\langle x_{t,k}-x_{*,k},\,\mathbf{E}_{t}[d_{t,k}]\right\rangle^{2}}{\mathbf{E}_{t}[\|d_{t,k}\|^{2}]}.$
Therefore, we have a contraction of expected distance to $x_{*}$ unless
$\langle x_{t,k}-x_{*,k},\,\mathbf{E}_{t}[d_{t,k}]\rangle=0,\qquad k=1,\ldots,m,$
which means either $x_{t,k}=x_{*,k}$ or $\mathbf{E}_{t}[d_{t,k}]$ is orthogonal to $x_{t,k}-x_{*,k}$ simultaneously for all $k$. The latter case means that the random direction $d_{t}$ contains no information that can help reduce the distance.
Apparently, the optimal stepsizes in (10) are not computable in practice, because we do not have access to $x_{*}$ and cannot evaluate the expectations precisely. In the next section, we explain how to make approximations and derive a practical stepsize rule.
2.2 Simplification and approximation
We first try to avoid the direct reliance on $x_{*}$. To this end, rewrite the numerator in (10) as
$\bigl{\langle}x_{t,k}-x_{*,k},\,\mathbf{E}_{t}[d_{t,k}]\bigr{\rangle}=\|x_{t,k}-x_{*,k}\|\|\mathbf{E}_{t}[d_{t,k}]\|\cos\theta_{t,k},$
where $\theta_{t,k}$ is the angle between the two vectors $x_{t,k}-x_{*,k}$ and $\mathbf{E}_{t}[d_{t,k}]$. We absorb the quantities related to $x_{*}$ into a tunable parameter
$\alpha_{t,k}=\|x_{t,k}-x_{*,k}\|\cos\theta_{t,k},$ (11)
which leads to the stepsizes
$\widetilde{\gamma}_{t,k}=\frac{\alpha_{t,k}\|\mathbf{E}_{t}[d_{t,k}]\|}{\mathbf{E}_{t}[\|d_{t,k}\|^{2}]},\qquad k=1,\ldots,m.$ (12)
Notice that any $\alpha_{t,k}$ we choose in practice may only be a (very rough) approximation of $\|x_{t,k}-x_{*,k}\|\cos\theta_{t,k}$. In particular, while the optimal stepsizes $\widehat{\gamma}_{t,k}$ can be positive or negative, it is very hard to even correctly estimate the sign of the inner product $\langle x_{t,k}-x_{*,k},\,\mathbf{E}_{t}[d_{t,k}]\rangle$. We take the pragmatic approach of restricting $\alpha_{t,k}>0$, effectively being optimistic that the expected direction $-\mathbf{E}_{t}[d_{t,k}]$ is always pointing towards $x_{*,k}$ (i.e., assuming $\cos\theta_{t,k}>0$).
A further simplification is to use a common stepsize schedule $\alpha_{t}$ across all blocks, i.e., $\alpha_{t,k}=\alpha_{t}$ for all $k=1,\ldots,m$. This can be a reasonable choice for training deep neural networks, where the model parameters are initialized randomly coordinate-wise such that $\mathbf{E}[\|x_{0,k}\|]$ is constant for each coordinate $k$ HeZhangRenSun2015 ; RobertsYaidaHanin2022 ; DinanYaidaZhang2023 . This brings us to
$\widetilde{\gamma}_{t,k}=\frac{\alpha_{t}\|\mathbf{E}_{t}[d_{t,k}]\|}{\mathbf{E}_{t}[\|d_{t,k}\|^{2}]},\qquad k=1,\ldots,m.$ (13)
We note that with some abuse of notation, here $\alpha_{t}$ denotes a scalar, not a vector of $(\alpha_{t,1},\ldots,\alpha_{t,k})$. This simplification motivates us to link $\alpha_{t}$ to the average of $\alpha_{t,k}$ in (11) and further relate to the overall distance $\|x_{t}-x_{*}\|$. Therefore, we expect $\alpha_{t}$ to decrease as $\|x_{t}-x_{*}\|$ gradually shrinks. A simple strategy is to use a monotonic stepsize schedule on $\alpha_{t}$, such as the popular cosine decay (loshchilov2016sgdr, ) or linear decay (defazio2024optimallineardecaylearning, ).
Next, we need to replace the conditional expectations $\mathbf{E}_{t}[d_{t,k}]$ and $\mathbf{E}_{t}[\|d_{t,k}\|^{2}]$ in (13) with computable approximations. We adopt the conventional approach of exponential moving average (EMA):
$u_{t,k}$ $=\beta u_{t-1,k}+(1-\beta)d_{t,k},$ (14)
$v_{t,k}$ $=\beta v_{t-1,k}+(1-\beta)\|d_{t,k}\|^{2},$
where $\beta\in[0,1)$ is the smoothing factor of EMA. Notice that $u_{t,k}\in\mathbf{R}^{n_{k}}$ and $v_{t,k}\in\mathbf{R}$, which have the same dimensions as $\mathbf{E}_{t}[d_{t,k}]$ and $\mathbf{E}_{t}[\|d_{t,k}\|^{2}]$ respectively. This leads to a set of practical stepsizes:
$\gamma_{t,k}=\alpha_{t}\frac{\|u_{t,k}\|}{v_{t,k}+\epsilon},\qquad k=1,\ldots,m,$ (15)
where we added a small constant $\epsilon>0$ in the denominator to improve numerical stability (preventing overflow in case $v_{t,k}$ is too small).
2.3 BCOS with one EMA estimator
The BCOS stepsizes in (15) are easy to implement in practice, but they are computed through the ratios of two online estimators $\|u_{t,k}\|$ and $v_{t,k}$. The ratio of two random quantities may be susceptible to large fluctuations if the numerator and denominator drift in different directions (i.e., when one is much larger than its mean while the other is much smaller than its mean). To avoid this problem, we further simplify BCOS to use only one EMA estimator.
First, recall the mean-variance decomposition of the conditional second moment:
$\mathbf{E}_{t}[\|d_{t,k}\|^{2}]=\|\mathbf{E}_{t}[d_{t,k}]\|^{2}+\mathbf{E}_{t}[\|d_{t,k}-\mathbf{E}_{t}[d_{t,k}]\|^{2}]=\|\mathbf{E}_{t}[d_{t,k}]\|^{2}+\textup{Var}_{t}(d_{t,k}),$
where $\textup{Var}_{t}(d_{t,k})$ denotes the conditional variance of $d_{t,k}$. We interpret $\|\mathbf{E}_{t}[d_{t,k}]\|^{2}$ as the signal power and $\textup{Var}_{t}(d_{t,k})$ as the noise power, and define the signal fraction (SiF)
$\rho_{t,k}=\frac{\|\mathbf{E}_{t}[d_{t,k}]\|^{2}}{\mathbf{E}_{t}[\|d_{t,k}\|^{2}]}=\frac{\|\mathbf{E}_{t}[d_{t,k}]\|^{2}}{\|\mathbf{E}_{t}[d_{t,k}]\|^{2}+\textup{Var}_{t}(d_{t,k})}.$ (16)
Apparently we have $\rho_{t,k}\in[0,1]$, and the two boundary values of 00 and $1$ correspond to the cases of zero signal and full signal (zero noise) respectively.
With the definition of SiF, we can decompose the stepsizes in (12) as
$\widetilde{\gamma}_{t,k}=\alpha_{t,k}\frac{\|\mathbf{E}_{t}[d_{t,k}]\|}{\mathbf{E}_{t}[\|d_{t,k}\|^{2}]}=\alpha_{t,k}\sqrt{\frac{\|\mathbf{E}_{t}[d_{t,k}]\|^{2}}{\mathbf{E}_{t}[\|d_{t,k}\|^{2}]}}\frac{1}{\sqrt{\mathbf{E}_{t}[\|d_{t,k}\|^{2}]}}=\frac{\alpha_{t,k}\sqrt{\rho_{t,k}}}{\sqrt{\mathbf{E}_{t}[\|d_{t,k}\|^{2}]}}.$
Now we can merge $\sqrt{\rho_{t,k}}\in[0,1]$ into the tunable parameters $\alpha_{t,k}$ and define
$\alpha^{\prime}_{t,k}:=\alpha_{t,k}\sqrt{\rho_{t,k}}.$
Then, following the same arguments as in Section 2.2, we choose to use a common stepsize schedule $\alpha^{\prime}_{t}$ and obtain the practical stepsize rule
$\gamma_{t,k}=\alpha^{\prime}_{t}\frac{1}{\sqrt{v_{t,k}}+\epsilon},\qquad k=1,\ldots,m,$ (17)
where $v_{t,k}$ is the EMA estimator of $\mathbf{E}_{t}[\|d_{t,k}\|^{2}]$ given in (14).
Another way to avoid using two estimators is to assimilate $\|\mathbf{E}_{t}[d_{t,k}]\|$, instead of $\sqrt{\rho_{t,k}}$, directly into the tunable parameter $\alpha^{\prime}_{t,k}$ and ultimately $\alpha^{\prime}_{t}$. This would result in
$\gamma_{t,k}=\alpha^{\prime}_{t}\,\frac{1}{v_{t,k}+\epsilon},\qquad k=1,\ldots,m.$ (18)
There is a major advantage of (17) over (18). Specifically, since the SiF $\rho_{t,k}$ is a dimensionless ratio, the per-iteration displacements,
$x_{t+1,k}-x_{t,k}=\alpha^{\prime}_{t}\frac{d_{t,k}}{\sqrt{v_{t,k}}+\epsilon},$
are invariant of scaling of the search directions $d_{t,k}$, which makes the tuning of $\alpha^{\prime}_{t}$ much easier in practice. This invariance property is also shared by Adam KingmaBa2014adam , as seen from (5) and (6). The similarity of Adam and the simplified BCOS in (17) is apparent, and we will explain their connection in detail in the next section.
3 Instantiations of BCOS
The derivation of BCOS in Section 2 is carried out with a general search direction $d_{t}$. In this section, we instantiate BCOS with two common choices of the search direction: the stochastic gradient and its EMA, also known as stochastic momentum.
To simplify presentation, we focus on the case of single-coordinate blocks, i.e., $m=n$ and $\mathcal{I}_{k}=\{k\}$ for $k=1,\ldots,n$. Then we can write the coordinate-wise EMA estimators $v_{t,k}$ collectively in a vector form:
$v_{t}=\beta v_{t-1}+(1-\beta)d_{t}^{2},$ (19)
where $d_{t}^{2}$ denotes the element-wise squared vector $d_{t}\odot d_{t}$. We also have $s_{t}=\gamma_{t}\in\mathbf{R}^{n}$ and
$x_{t+1}=x_{t}-\gamma_{t}\odot d_{t}$
where the vector of coordinate-wise stepsizes can be expressed as
$\gamma_{t}=\alpha_{t}\frac{1}{\sqrt{v_{t}}+\epsilon}$ (20)
Here $v_{t}+\epsilon$ means element-wise addition of $\epsilon$, $\sqrt{v_{t}}+\epsilon$ denotes element-wise square root, and the fraction represents element-wise division or reciprocal. We no longer distinguish between $\alpha_{t}$ and $\alpha^{\prime}_{t}$ because they are both tunable hyper-parameters.
Algorithm 1 BCOS-g (RMSprop)
input: $x_{0}$, $\{\alpha_{t}\}_{t\geq 0}$, $\beta\in[0,1)$, $\epsilon>0$
$v_{-1}=\nabla f(x_{0},\xi_{0})^{2}$
for $t=0,1,2,\ldots$ do
     $g_{t}=\nabla f(x_{t},\xi_{t})$
     $v_{t}=\beta v_{t-1}+(1-\beta)g_{t}^{2}$
     $x_{t+1}=x_{t}-\alpha_{t}\frac{g_{t}}{\sqrt{v_{t}}+\epsilon}$
Algorithm 2 BCOS-m
input: $x_{0}$, $\{\alpha_{t}\}$, $\beta_{1},\beta_{2}\in[0,1)$, $\epsilon>0$
$m_{-1}=\nabla f(x_{0},\xi_{0})$
$v_{-1}=m_{-1}^{2}$
for $t=0,1,2,\ldots$ do
     $g_{t}=\nabla f(x_{t},\xi_{t})$
     $m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})g_{t}$
     $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2}){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{t}^{2}}$
     $x_{t+1}=x_{t}-\alpha_{t}\frac{m_{t}}{\sqrt{v_{t}}+\epsilon}$
3.1 BCOS with EMA estimator
We first present instantiations of BCOS with an EMA estimator (derived in Section 2.3).
BCOS-g
Algorithm 1 is the instantiation of BCOS using $g_{t}=\nabla f(x_{t},\xi_{t})$ as the search direction, where $g_{t}^{2}$ denotes element-wise square and $\frac{g_{t}}{\sqrt{v_{t}}+\epsilon}$ means element-wise division. We call it BCOS-g to signify the use of gradient as search direction.
We immediately recognize that BCOS-g is exactly the RMSprop algorithm rmsprop2012lecture , which is one of the first effective algorithms for training deep neural networks. Our BCOS framework gives a novel interpretation of RMSprop and its effectiveness. In the special case with $\beta=0$ and $\epsilon=0$, we have $v_{t}=g_{t}^{2}$, and BCOS-g becomes the sign gradient method
$x_{t+1}=x_{t}-\alpha_{t}\,\textup{sign}(g_{t}),$ (21)
which also received significant attention in the literature Polyak1973pseudo ; bernstein2018signsgd ; safaryan2021signdescent ; jiang2024efficient .
BCOS-m
Using the stochastic momentum as search direction has a long history in stochastic approximation (e.g., GupalBazhenov72, ; Polyak77comparison, ; RuszczynskiSyski83stats, ). It has become the default option for modern deep learning due to its superior performance compared with using plain stochastic gradients. In Algorithm 2, we use $m_{t}$ to denote the momentum, which is a standard notation in the machine learning literature. We call it BCOS-m to signify the use of momentum as the search direction. BCOS-m employs a second smoothing factor $\beta_{2}$ to calculate the EMA of $m_{t}^{2}$. The two smoothing factors $\beta_{1}$ and $\beta_{2}$ do not need to be the same and can be chosen independently in practice. Similar to BCOS-g, the special case of BCOS-m with $\beta_{2}=0$ and $\epsilon=0$ corresponds to the sign-momentum method
$x_{t+1}=x_{t}-\alpha_{t}\,\textup{sign}(m_{t}).$ (22)
We notice that BCOS-m is very similar to Adam as given in (5) and (6). The difference is that in Adam, $v_{t}$ is the EMA of $g_{t}^{2}$ instead of $m_{t}^{2}$. From the perspective of the BCOS framework, Adam has a mismatch between the search direction $m_{t}$ and the second moment estimator based on $g_{t}^{2}$, which must be compensated for by using a larger smoothing factor $\beta_{2}$ (because $m_{t}$ itself is a smoothed version of $g_{t}$). For BCOS-m, using $\beta_{2}=\beta_{1}$ produces as good performance as Adam with the best tuned $\beta_{2}$ (see numerical experiments in Section 4).
In both BCOS-g and BCOS-m, we initialize $v_{-1}=g_{0}^{2}$, therefore $v_{0}=g_{0}^{2}$ as well. This avoids the initial bias caused by setting $v_{-1}=0$. An alternative way for bias correction is to initialize $m_{-1}$ and $v_{-1}$ as the zero vector, and then multiply $m_{t}$ and $v_{t}$ by $1/(1-\beta^{t+1})$, which quickly converges to $1$, as done in Adam KingmaBa2014adam . Our numerical experiments show that these two choices do not have significant differences in practice.
3.2 BCOS with conditional estimator
Recall that the optimal stepsizes $\widehat{\gamma}_{t,k}$ in (10) and their simplifications $\widetilde{\gamma}_{t,k}$ in (13) are all based on conditional expectation. In Section 3.1, we used coordinate-wise EMA of $d_{t}^{2}$ to approximate the conditional expectation $\mathbf{E}_{t}[d_{t}^{2}]$, i.e., $v_{t}$ as estimator of $\mathbf{E}_{t}[d_{t}^{2}]$ in BCOS-g and of $\mathbf{E}_{t}[m_{t}^{2}]$ in BCOS-m, respectively. In this section, we show that with $m_{t}$ as the search direction, we can exploit its update form to derive effective conditional estimators that can avoid using EMA.
Algorithm 3 BCOS-c
input: $x_{0}$, $\{\alpha_{t}\}$, $\beta\in[0,1)$, $\epsilon>0$
$m_{-1}=\nabla f(x_{0},\xi_{0})$
for $t=0,1,2,\ldots$ do
     $g_{t}=\nabla f(x_{t},\xi_{t})$
     $m_{t}=\beta m_{t-1}+(1-\beta)g_{t}$
     $v_{t}=\bigl{(}1-(1-\beta)^{2}\bigr{)}m^{2}_{t-1}+(1-\beta)^{2}g^{2}_{t}$
     $x_{t+1}=x_{t}-\alpha_{t}\frac{m_{t}}{\sqrt{v_{t}}+\epsilon}$
Algorithm 4 BCOSW-c
input: $x_{0}$, $\{\alpha_{t}\}$, $\beta\in[0,1)$, $\lambda\geq 0$, $\epsilon>0$
$m_{-1}=\nabla f(x_{0},\xi_{0})$
for $t=0,1,2,\ldots$ do
     $g_{t}=\nabla f(x_{t},\xi_{t})$
     $m_{t}=\beta m_{t-1}+(1-\beta)g_{t}$
     $v_{t}=\bigl{(}1-(1-\beta)^{2}\bigr{)}m^{2}_{t-1}+(1-\beta)^{2}g^{2}_{t}$
     $x_{t+1}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\alpha_{t}\lambda)}x_{t}-\alpha_{t}\frac{m_{t}}{\sqrt{v_{t}}+\epsilon}$
We first repeat the definition of momentum with a smoothing factor $\beta\in[0,1)$, i.e.,
$m_{t}=\beta m_{t-1}+(1-\beta)g_{t}.$ (23)
To derive an estimator of $\mathbf{E}_{t}[m_{t}^{2}]$, we expand the square and take the expectation of each term:
$\mathbf{E}_{t}\bigl{[}m_{t}^{2}\bigr{]}$ $=\mathbf{E}_{t}\bigl{[}(\beta m_{t-1}+(1-\beta)g_{t})^{2}\bigr{]}$
$=\beta^{2}\mathbf{E}_{t}\bigl{[}m_{t-1}^{2}\bigr{]}+2\beta(1-\beta)\mathbf{E}_{t}[m_{t-1}\odot g_{t}]+(1-\beta)^{2}\mathbf{E}_{t}\bigl{[}g_{t}^{2}\bigr{]}$
$=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}\odot\mathbf{E}_{t}[g_{t}]+(1-\beta)^{2}\mathbf{E}_{t}\bigl{[}g_{t}^{2}\bigr{]}.$ (24)
where we used the fact $\mathbf{E}_{t}[m_{t-1}^{2}]=m_{t-1}^{2}$ and $\mathbf{E}_{t}[m_{t-1}]=m_{t-1}$ thanks to the definition of $\mathbf{E}_{t}[\cdot]$ in (8). It remains to approximate $\mathbf{E}_{t}[g_{t}]$ and $\mathbf{E}_{t}[g_{t}^{2}]$.
Clearly a good estimator for $\mathbf{E}_{t}[g_{t}]$ is $m_{t}$. To approximate $\mathbf{E}_{t}[g_{t}^{2}]$, we could use a separate EMA estimator $v^{\prime}_{t}=\beta^{\prime}v^{\prime}_{t-1}+(1-\beta^{\prime})g_{t}^{2}$, but this introduces another algorithm state $v^{\prime}_{t}$ and a second smoothing factor $\beta^{\prime}$. Meanwhile, we observe that the factor $(1-\beta)^{2}$ multiplying $\mathbf{E}_{t}[g_{t}^{2}]$ is usually very small, especially for $\beta$ close to 1. As a result, any error in approximating $\mathbf{E}_{t}[g_{t}^{2}]$ is attenuated by a very small factor and may not cause much difference. Therefore, for simplicity, we choose to approximate $\mathbf{E}_{t}[g_{t}^{2}]$ with $g_{t}^{2}$ itself. Combining with approximating $\mathbf{E}_{t}[g_{t}]$ with $m_{t}$, we arrive at the following conditional estimator for $\mathbf{E}_{t}[m_{t}^{2}]$:
$v_{t}=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}\odot m_{t}+(1-\beta)^{2}g_{t}^{2}.$ (25)
Our numerical experiments confirmed that this indeed is a good estimator which obtains performance comparable to Adam.
While effective in practice, the estimator in (25) looks complicated and does not carry much intuition. We are motivated to derive another one that is much simpler and as effective. The key is to approximate $\mathbf{E}[g_{t}]$ in (24) with $m_{t-1}$ instead of $m_{t}$, which results in
$v_{t}$ $=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$
$=\left(1-(1-\beta)^{2}\right)m^{2}_{t-1}+(1-\beta)^{2}g^{2}_{t}.$ (26)
This estimator is much simpler than (25). Moreover, it resembles the standard EMA estimator in Adam, shown in (6), with an effective smoothing factor
$\beta^{\prime}=1-(1-\beta)^{2},$
but with $v_{t-1}$ replaced by $m^{2}_{t-1}$. As a result, the estimator in (26) does not need to store $v_{t-1}$, thus requiring fewer optimizer states (less memory). This also explains that the second smoothing factor in Adam, $\beta_{2}$, corresponding to $\beta^{\prime}$ here, should be much larger or closer to 1 than $\beta$. Specifically, $\beta=0.9$ roughly corresponds to $\beta^{\prime}=0.99$. The estimator in (26) eliminates $\beta_{2}$ as a second hyper-parameter.
Finally, replacing $v_{t}$ in BCOS-m with the one in (26) produces Algorithm 3. We call it BCOS-c to signify the conditional estimator. It has fewer optimizer states and fewer hyper-parameters to tune than BCOS-m and Adam. We note the slight abuse of notation in using the symbol $v_{t}$ for both the EMA estimator in BCOS-g and -m and the conditional estimator here. In general, $v_{t}$ denotes an online estimator of $\mathbf{E}_{t}[d_{t}^{2}]$. The corresponding choice of $d_{t}$ and type of estimator should be clear from the context.
Figure 1: Comparing AdamW and BCOSW-c with different momentum parameters.
3.3 BCOS with decoupled weight decay
Weight decay is a common practice in training deep learning models to obtain better generalization performance. It can be understood as adding an $L_{2}$ regularization to the loss function, i.e., minimizing the regularized loss
$F(x)=\mathbf{E}_{\xi}[f(x,\xi)]+\frac{\lambda}{2}\|x\|^{2},$ (27)
where $\lambda>0$ is a regularization parameter. Effectively, the stochastic gradient at $x_{t}$ becomes $\nabla f(x_{t},\xi_{t})+\lambda x_{t}$. We can apply the BCOS family of algorithms to incorporate the regularization term by simply replacing $g_{t}=\nabla f(x_{t},\xi_{t})$ with $g_{t}=\nabla f(x_{t},\xi_{t})+\lambda x_{t}$.
A more effective way in practice is to use decoupled weight decay as proposed in the AdamW algorithm Loshchilov2019AdamW . Specifically, we apply weight decay separately in the BCOS update:
$x_{t+1}=x_{t}-\gamma_{t}\odot d_{t}-\alpha_{t}\lambda x_{t}=(1-\alpha_{t}\lambda)x_{t}-\gamma_{t}\odot d_{t}.$
We call the resulting method BCOSW following the naming convention of AdamW.
Algorithm 4 shows BCOSW-c, which uses the conditional estimator. Other variants, namely BCOSW-g and BCOSW-m, can be obtained similarly. A PyTorch implementation of all variants of BCOS and BCOSW is given in Appendix A. The code is also available at https://github.com/facebookresearch/bcos.
4 Numerical experiments
Figure 2: Comparing AdamW and three variants of BCOSW. Left: the first 10k iterations; Right: all 100k iterations.
In this section, we present preliminary experiments to compare BCOS with Adam, specifically their variants with decoupled weight decay. Among the BCOSW family, we focus on BCOSW-c (Algorithm 4).
Our first set of experiments is conducted on training a small GPT2 model with 124 million parameters radford2019language on the OpenWebText dataset Gokaslan2019OpenWeb . We use global batch size 512 and run all experiments for 100k iterations. We use linear warmup on the stepsize schedule $\{\alpha_{t}\}$ to reach $\alpha_{\text{max}}$ in the first 2k iterations, and then use cosine decay to toward the final stepsize $\alpha_{\text{min}}=0.01\alpha_{\text{max}}$. The default hyper-parameters are chosen (based on a coarse sweep) as: peak stepsize $\alpha_{\text{max}}=0.002$, $\epsilon=10^{-6}$ and weight decay $\lambda=0.1$.
Figure 1 (left) shows the test loss of AdamW on the GPT2 task with different combinations of $\beta_{1}$ and $\beta_{2}$. For each value of $\beta_{1}\in\{0.8,0.9,0.95\}$, we choose the best $\beta_{2}$ after sweeping $\beta_{2}\in\{0.8,0.9,0.95,0.975,0.99\}$. Their final losses achieved are all very close, around the value $2.82$. For most $(\beta_{1},\beta_{2})$ combinations, we observe some loss spikes, especially at the beginning of the training (as shown in the inset). In contrast, Figure 1 (right) shows that BCOSW-c obtains the same final loss but with very smooth loss curves.
Figure 3: Left: Adam(W) with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. Right: BCOS(W)-c with $\beta=0.9$.
Figure 4: Train and test loss by varying max value of stepsize schedule.
Figure 5: Left: ResNet-20 on CIFAR10. Right: Vision Transformer on ImageNet.
Figure 2 compares AdamW against the three variants BCOSW-g, -m, and -c. We observe that BCOSW-g is significantly worse than the momentum-based methods. The loss curves for the momentum-based methods are all very close, but with spikes for both AdamW and BCOSW-m, while BCOSW-c has very smooth loss curves.
Figure 3 illustrates the difference between algorithms with and without decoupled weight decay. BCOS-c converges to much higher loss than BCOSW-c, and different values of $\lambda$ (weight decay) make a dramatic difference for BCOS-c but cause little change to BCOSW-c. The same phenomenon happens for Adam versus AdamW, and we again observe more spikes from the loss curves of Adam(W) than BCOS(W)-c.
Figure 4 shows how the final training and validation losses (after 100k iterations) vary with $\alpha_{\text{max}}$. For each stepsize schedule, we repeat the training three times with different random seeds and plot the standard deviation as error bars, which are mostly invisible due to small variations among different runs. We observe that AdamW and BCOSW obtain the lowest losses at similar values of $\alpha_{\text{max}}$. BCOSW has slightly better performance for smaller values of $\alpha_{\text{max}}$, but becomes more sensitive for larger values of $\alpha_{\text{max}}$.
[…]
For training ViT on ImageNet, although the best tuned stepsize schedules are similar between AdamW and BCOSW, their training and test curves look quite different. Figure 5 (right) shows that the test precision curves for BCOSW-c rise slowly but reach slightly higher precision at the end.
These preliminary experiments demonstrate that BCOSW-c can obtain competitive performance compared with the state-of-the-art method AdamW, but with fewer optimizer states and fewer hyper-parameters to tune. More comprehensive empirical studies are needed to fully understand its potential in training large deep learning models.
5 Convergence analysis
In this section, we present convergence analysis for BCOS with a general search direction and decoupled weight decay (see Section 3.3). To simplify presentation, we stick with the setting of single-coordinate blocks. Specifically, we analyze the algorithm
$x_{t+1}=(1-\alpha_{t}\lambda)x_{t}-\gamma_{t}\odot d_{t},\quad\text{where}\quad\gamma_{t}=\frac{\alpha_{t}}{\sqrt{v_{t}+\epsilon}}.$ (28)
Here, stepsize $\gamma_{t}$ takes a slightly different form from the one used throughout Section 3, where $\epsilon$ is outside the square-root in the denominator. The resulting difference between the two forms is negligible for corresponding choices of $\epsilon$, with the one in (28) being squared (e.g., being $10^{-12}$ if the one in Section 3 is $10^{-6}$). We choose the form in (28) for ease of presenting the analysis.
Recall that $v_{t}\in\mathbf{R}_{+}^{n}$ is an online estimator of $\mathbf{E}_{t}[d_{t}^{2}]$. Our analysis consists of two stages:
  • First, we analyze the convergence behavior of the conceptual BCOS method
    $x_{t+1}=(1-\alpha_{t}\lambda)x_{t}-\widetilde{\gamma}_{t}\odot d_{t},\quad\text{where}\quad\widetilde{\gamma}_{t}=\frac{\alpha_{t}}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}.\qquad$ (29)
    It is called “conceptual” because we cannot compute $\mathbf{E}_{t}[d_{t}^{2}]$ exactly in practice. But its analysis is essential for understanding and analyzing the practical method. Our analysis relies on an aiming condition, which originates from the classical stochastic approximation literature RobbinsMonro51 ; Blum1954multidim ; chung1954 ; Sakrison1966 . Under this assumption, we establish almost sure convergence of $x_{t}$ to a target point $x_{*}$ and the $\mathcal{O}(1/t)$ convergence rate of $\mathbf{E}[\|x_{t}-x_{*}\|^{2}]$.
  • To analyze the practical BCOS method (28), we first bound the difference between the expected steps under the practical method and the conceptual method, i.e.,
    $\bigl{|}\mathbf{E}_{t}[\gamma_{t}\odot d_{t}]-\mathbf{E}_{t}[\widetilde{\gamma}_{t}\odot d_{t}]\bigr{|}\leq\sigma_{t}\bigl{|}\mathbf{E}_{t}[\widetilde{\gamma}\odot d_{t}]\bigr{|}+\mathcal{O}(\epsilon),$ (30)
    where $\leq$ denotes coordinate-wise inequalities between two vectors and $\sigma_{t}$ is a scalar that depends on the bias and variance of $v_{t}$ as an online estimator of $\mathbf{E}_{t}[d_{t}^{2}]$. We prove almost sure convergence of $x_{t}$ to a neighborhood of $x_{*}$, where the radius of the neighborhood is determined by the magnitude of $\sigma_{t}$ and $\mathcal{O}(\epsilon)$.
The rest of this section is organized as follows. In Section 5.1, we introduce the aiming condition and discuss its connection with convexity. We present the convergence analysis for the conceptual BCOS method in Section 5.2 and then the analysis for the practical BCOS method in Section 5.3. Our analysis is very general and applies to the whole BCOS family of methods, including RMSprop, Adam, AdamW, and variants of spectral BCOS. To obtain specific results for each method, we only need to estimate the bias and variance of the second-moment estimator $v_{t}$, which are presented in Section 5.4.
5.1 Aiming condition
First consider the stochastic approximation method (7). A typical set of assumptions in the classical stochastic approximation literature (specifically Sakrison1966 ) includes:
  • The random variables $\{\xi_{t}\}_{t\geq 0}$ are independent and $g(x,\xi)$ has bounded variance;
  • The step sizes $\{\alpha_{t}\}_{t\geq 0}$ are nonnegative and satisfy $\sum_{t=0}^{\infty}\alpha_{t}=\infty$ and $\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty$;
  • Suppose $\mathbf{E}_{\xi}[g(x_{*},\xi)]=0$ and there exists $L,\mu>0$ such that for all $x\in\mathbf{R}^{n}$,
    $L\|x-x_{*}\|^{2}~{}\geq~{}\bigl{\langle}x-x_{*},\,\mathbf{E}_{\xi}[g(x,\xi)]\bigr{\rangle}~{}\geq~{}\mu\|x-x_{*}\|^{2}.$ (31)
Under these conditions, the sequence $\{x_{t}\}$ converges to $x_{*}$ in the mean-square sense, i.e., $\lim_{t\to\infty}\mathbf{E}[\|x_{t}-x_{*}\|^{2}]=0$ Sakrison1966 . In particular, the last condition (31) ensures that the expected direction $\mathbf{E}_{\xi}[g(x,\xi)]$ always points towards the target $x_{*}$, implying that the distance from $x_{*}$ can be reduced along this direction (with appropriate stepsize). For this reason, we call it an aiming condition. Similar results can be established under weaker aiming conditions than (31); for example, without the upper bound on the inner product. But they are more subtle and require more sophisticated analysis (e.g., RobbinsMonro51, ; Wolfowitz1952, ; Blum1954prob-one, ; Blum1954multidim, ; Dvoretzky1956, ; DermanSacks1959, ; Wasan69book, ).
Following the same idea, we rewrite the conceptual BCOS method (29) as
$x_{t+1}=x_{t}-\alpha_{t}\lambda x_{t}-\widetilde{\gamma}_{t}\odot d_{t}=x_{t}-\alpha_{t}\left(\frac{d_{t}}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{t}\right)$
and impose an aiming condition on the expected update direction
$\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{t}\right]=\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{t}.$
Assumption A (Aiming condition).
There exists $x_{*}\in\mathbf{R}^{n}$ such that
$\left\langle x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{t}\right\rangle\geq\lambda\|x_{t}-x_{*}\|^{2}$ (32)
holds for all $t\geq 0$ almost surely.
We make the following remarks regarding Assumption A.
  • The condition (32) in general depends on the past trajectory $\{x_{0},d_{0},x_{1},d_{1},\ldots,x_{t}\}$. If $\mathbf{E}_{t}[d_{t}]=\mathbf{E}[d_{t}|x_{t}]$, then we can drop the subscript $t$ and obtain a trajectory-independent condition like (31). This is indeed the case when $d_{t}=g(x_{t},\xi_{t})$ and $\{\xi_{0},\ldots,\xi_{t}\}$ are mutually independent. We need (32) to cover more general cases, e.g., when $d_{t}$ is the stochastic momentum, which clearly depends on the past trajectory.
  • Instead of a generic constant $\mu$ as in (31), we have $\lambda$ on the right-hand side of (32) due to the use of decoupled weight decay, interpreted as $L_{2}$-regularization in (27).
  • We do not need an upper bound on the inner product as in (31) due to the normalization effect of dividing $d_{t}$ by $\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}$ in BCOS. A similar one-sided aiming condition was used in Blum1954multidim to establish almost sure convergence of the first multi-dimensional extension of the Robbins-Monro method.
Let $\rho_{t}\in[0,1]^{n}$ be the vector of coordinate-wise SiF defined in (16). Then we have
$\frac{\mathbf{E}_{t}\left[d_{t}\right]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}=\sqrt{\frac{\mathbf{E}_{t}[d_{t}]^{2}}{\mathbf{E}_{t}[d_{t}^{2}]}}\,\textup{sign}(\mathbf{E}_{t}[d_{t}])=\sqrt{\rho_{t}}\odot\textup{sign}(\mathbf{E}_{t}[d_{t}]).$
Correspondingly, the aiming condition (32) can be written as
$\bigl{\langle}x_{t}-x_{*},\,\sqrt{\rho_{t}}\odot\textup{sign}(\mathbf{E}_{t}[d_{t}])+\lambda x_{t}\bigr{\rangle}\geq\lambda\|x_{t}-x_{*}\|^{2}.$ (33)
Intuitively, the inner product in the aiming condition is predominantly determined by the coordinates that are far from optimal, in other words, the coordinates with relatively large $|x_{t,k}-x_{*,k}|$ and having a large SiF $\rho_{t,k}$ (effectively high signal-to-noise ratio).
5.1.1 Connection with convexity
In the context of stochastic optimization, i.e., minimizing $F(x)=\mathbf{E}_{\xi}[f(x,\xi)]$ with $g(x,\xi)=\nabla f(x,\xi)$, the classical aiming condition (31) is implied by $F$ being $\mu$-strongly convex and $L$-smooth (e.g., Nesterov04book, ). But in general it is a much weaker condition. Our aiming condition in Assumption A does not relate directly to smoothness and is not always implied by (strong) convexity either. However, it does have overlapping characteristics with convexity.
For ease of comparison, we consider the general block structure described in Section 1.3 with $d_{t}=\nabla f(x_{t},\xi_{t})$ and set $\lambda=0$. In this case, we have $\mathbf{E}[d_{t}]=\nabla F(x_{t})$ and
$\sum_{k=1}^{m}\left\langle x_{k}-x_{\ast,k},~{}\frac{\nabla_{k}F(x)}{\sqrt{\mathbf{E}\bigl{[}\left\|\nabla_{k}f(x,\xi)\right\|^{2}\bigr{]}}}\right\rangle\geq 0,\qquad\forall\,x,$ (34)
where $\nabla_{k}F(x)\in\mathbf{R}^{n_{k}}$ denotes the sub-vector of $\nabla F(x)$ corresponding to the $k$th block. In the specific case of a full-dimensional block ($m=1$), the denominator $\sqrt{\mathbf{E}\bigl{[}\|\nabla f(x,\xi)\|^{2}\bigr{]}}$ can be omitted without affecting the inequality and the aiming condition simplifies to:
$\left\langle x-x_{\ast},\nabla F(x)\right\rangle\geq 0,\qquad\forall\,x.$ (35)
This is a direct consequence of convexity, which is characterized by (e.g., Nesterov04book, ):
$\left\langle x-y,\nabla F(x)-\nabla F(y)\right\rangle\geq 0,\qquad\forall\,x,y.$
To see the implication, simply substituting $y=x_{\ast}$ and $\nabla F(x_{\ast})=0$ into the above inequality gives the full-dimensional aiming condition.
However, other than the full-dimensional block case, the aiming condition (34) may exhibit significant departure from convexity. As an example, consider the case of single coordinate blocks with uniform SiF, i.e., $\rho_{t,k}=\rho_{t,k^{\prime}}$ for all $k,k^{\prime}\in\{1,\ldots,n\}$. In light of the equivalent formulation using SiF in (33), the corresponding aiming condition becomes
$\left\langle x-x_{\ast},\,\textup{sign}(\nabla F(x))\right\rangle\geq 0,\qquad\forall\,x.$ (36)
Comparing this condition with convexity, it is easy to construct examples that satisfy one of them while failing the other. We provide some simple examples in Appendix B.1.
5.2 Analysis of conceptual BCOS
In this section, we provide convergence analysis for the conceptual BCOS method in (29). Our first result is about its one-step contraction property.
Lemma 5.1.
Suppose Assumption A holds, $\alpha_{t}\geq 0$ and $\alpha_{t}\lambda<1$ for all $t\geq 0$. Then the sequence $\{x_{t}\}$ generated by (29) satisfies, for all $t\geq 0$,
$\mathbf{E}_{t}\bigl{[}\left\|x_{t+1}-x_{\ast}\right\|^{2}\bigr{]}\leq(1-\alpha_{t}\lambda)^{2}\left\|x_{t}-x_{\ast}\right\|^{2}+\alpha_{t}^{2}B_{*},$ (37)
where
$B_{\ast}=n+\lambda^{2}\left\|x_{\ast}\right\|^{2}+2\lambda\left\|x_{\ast}\right\|_{1}.$ (38)
Proof.
First, we notice that the aiming condition (32) is equivalent to
$\biggl{\langle}x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{*}\biggr{\rangle}\geq 0.$ (39)
To see this, we simply add and subtract $\lambda x_{t}$ inside the inner product:
$\biggl{\langle}x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{*}\biggr{\rangle}$ $=\biggl{\langle}x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{t}-\lambda x_{t}+\lambda x_{*}\biggr{\rangle}$
$=\biggl{\langle}x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{t}\biggr{\rangle}-\lambda\|x_{t}-x_{*}\|^{2}.$
Given $x_{t+1}=x_{t}-\widetilde{\gamma}_{t}\odot d_{t}-\alpha_{t}x_{t}$, we have
$\mathbf{E}_{t}\bigl{[}\|x_{t+1}-x_{\ast}\|^{2}\bigr{]}$ $=\mathbf{E}_{t}\!\left[\bigl{\|}x_{t}-\widetilde{\gamma}_{t}\odot d_{t}-\alpha_{t}\lambda x_{t}-x_{\ast}\bigr{\|}^{2}\right]$
$=\mathbf{E}_{t}\!\left[\bigl{\|}(1-\alpha_{t}\lambda)x_{t}-\widetilde{\gamma}_{t}\odot d_{t}-(1-\alpha_{t}\lambda)x_{\ast}-\alpha_{t}\lambda x_{\ast}\bigr{\|}^{2}\right]$
$=\mathbf{E}_{t}\!\left[\left\|(1-\alpha_{t}\lambda)(x_{t}-x_{*})-\bigl{(}\widetilde{\gamma}_{t}\odot d_{t}+\alpha_{t}\lambda x_{\ast}\bigr{)}\right\|^{2}\right]$
$=\mathbf{E}_{t}\!\left[\bigl{\|}(1-\alpha_{t}\lambda)(x_{t}-x_{\ast})\bigr{\|}^{2}\right]-2\mathbf{E}_{t}\!\left[\bigl{\langle}(1-\alpha_{t}\lambda)(x_{t}-x_{\ast}),\,\widetilde{\gamma}_{t}\odot d_{t}+\alpha_{t}\lambda x_{\ast}\bigr{\rangle}\right]$
$\quad+\mathbf{E}_{t}\!\left[\bigl{\|}\widetilde{\gamma}_{t}\odot d_{t}+\alpha_{t}\lambda x_{\ast}\bigr{\|}^{2}\right].$ (40)
We examine the three terms in the last equation one by one. Using the definition of $\mathbf{E}_{t}[\cdot]$, we can remove the conditional expectation from the first term, i.e.,
$\mathbf{E}_{t}\!\left[\bigl{\|}(1-\alpha_{t}\lambda)(x_{t}-x_{\ast})\bigr{\|}^{2}\right]=\bigl{\|}(1-\alpha_{t}\lambda)(x_{t}-x_{\ast})\bigr{\|}^{2}=(1-\alpha_{t}\lambda)^{2}\bigl{\|}x_{t}-x_{\ast}\bigr{\|}^{2}.$
For the second term, we use $\widetilde{\gamma}_{t}=\alpha_{t}/\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}$ to obtain
$\mathbf{E}_{t}\!\left[\bigl{\langle}(1-\alpha_{t}\lambda)(x_{t}-x_{\ast}),\,\widetilde{\gamma}_{t}\odot d_{t}+\alpha_{t}\lambda x_{\ast}\bigr{\rangle}\right]$ $=\bigl{\langle}(1-\alpha_{t}\lambda)(x_{t}-x_{\ast}),\,\widetilde{\gamma}_{t}\odot\mathbf{E}_{t}[d_{t}]+\alpha_{t}\lambda x_{\ast}\bigr{\rangle}$
$=\alpha_{t}(1-\alpha_{t}\lambda)\biggl{\langle}x_{t}-x_{\ast},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{\ast}\biggl{\rangle}$
$\geq 0,$
where the last inequality is due to (39). For the third term in (40), we have
$\mathbf{E}_{t}\!\left[\bigl{\|}\widetilde{\gamma}_{t}\odot d_{t}+\alpha_{t}\lambda x_{\ast}\bigr{\|}^{2}\right]$ $=\alpha_{t}^{2}\,\mathbf{E}_{t}\!\left[\biggl{\|}\frac{d_{t}}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{*}\biggr{\|}^{2}\right]$
$=\alpha_{t}^{2}\,\mathbf{E}_{t}\!\left[\sum_{k=1}^{n}\left(\frac{d_{t,k}^{2}}{\mathbf{E}_{t}[d_{t,k}^{2}]}+\lambda^{2}x_{*,k}^{2}+2\lambda x_{*,k}\frac{d_{t,k}}{\sqrt{\mathbf{E}_{t}[d_{t,k}^{2}]}}\right)\right]$
$=\alpha_{t}^{2}\,\left[\sum_{k=1}^{n}\left(\frac{\mathbf{E}_{t}[d_{t,k}^{2}]}{\mathbf{E}_{t}[d_{t,k}^{2}]}+\lambda^{2}x_{*,k}^{2}+2\lambda x_{*,k}\frac{\mathbf{E}[d_{t,k}]}{\sqrt{\mathbf{E}_{t}[d_{t,k}^{2}]}}\right)\right]$
$=\alpha_{t}^{2}\sum_{k=1}^{n}\left(1+\lambda^{2}x_{*,k}^{2}+2\lambda x_{*,k}\sqrt{\rho_{t,k}}\right)$
$=\alpha_{t}^{2}\bigl{(}n+\lambda^{2}\|x_{*}\|^{2}+2\lambda\langle x_{*},\sqrt{\rho_{t}}\rangle\bigr{)}$
$\leq\alpha_{t}^{2}\bigl{(}n+\lambda^{2}\|x_{*}\|^{2}+2\lambda\|x_{*}\|_{1}\|\sqrt{\rho_{t}}\|_{\infty}\bigr{)}$
$\leq\alpha_{t}^{2}\bigl{(}n+\lambda^{2}\|x_{*}\|^{2}+2\lambda\|x_{*}\|_{1}\bigr{)},$
where the last two inequalities are due to Hölder’s inequality $\langle x_{*},\sqrt{\rho_{t}}\rangle\leq\|x_{*}\|_{1}\|\sqrt{\rho_{t}}\|_{\infty}$ and the fact that $\|\sqrt{\rho_{t}}\|_{\infty}\leq 1$, respectively. Plugging the above equality and inequalities into (40) yields the desired result. ∎
As a direct consequence of Lemma 5.1, there exists sufficiently small $\alpha_{t}>0$ such that
$\mathbf{E}_{t}\bigl{[}\left\|x_{t+1}\!-\!x_{\ast}\right\|^{2}\bigr{]}\leq\|x_{t}\!-\!x_{*}\|^{2},\qquad\forall\,t\geq 0.$
In fact, we can prove much stronger results, including almost sure convergence of $x_{t}$ to $x_{*}$ and characterizing the rate of convergence of $\mathbf{E}[\|x_{t}-x_{*}\|^{2}]$. These are presented in the following subsections.
5.2.1 Almost sure convergence
The following theorem establishes almost sure convergence of the conceptual BCOS algorithm under a classical condition on the sequence $\{\alpha_{t}\}$.
Theorem 5.1 (Almost sure convergence).
Suppose the stepsize schedule $\{\alpha_{t}\}_{t\geq 0}$ satisfies
$\alpha_{t}\geq 0,\quad\alpha_{t}\lambda\leq 1,\quad\forall\,t\geq 0,\qquad\sum_{t=0}^{\infty}\alpha_{t}=\infty,\qquad\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty.$ (41)
The under Assumption A, the sequence $\{x_{t}\}$ generated by (29) satisfies $\left\|x_{t}-x_{\ast}\right\|\to 0$ a.s.
The proof of Theorem 5.1 is based on the following almost supermartingale lemma of Robbins and Siegmund ROBBINS1971 .
Lemma 5.2 (“Almost supermartingale”, (ROBBINS1971, , Theorem 1)).
) Let $(\Omega,\mathcal{F},P)$ be a probability space, and $\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\ldots$ be a sequence of sub-$\sigma$-algebras of $\mathcal{F}$. For each $t\geq 0$, let $X_{t}$, $a_{t}$, $b_{t}$ and $c_{t}$ be non-negative $\mathcal{F}_{t}$-measurable random variables such that
$\mathbf{E}[X_{t+1}|\mathcal{F}_{t}]\leq X_{t}(1+a_{t})+b_{t}-c_{t}.$ (42)
If $\sum_{t=0}^{\infty}a_{t}<\infty$ and $\sum_{t=0}^{\infty}b_{t}<\infty$, then $\lim_{t\to\infty}X_{t}$ exists and is finite, and $\sum_{t=0}^{\infty}c_{t}<\infty$ almost surely (a.s.).
Proof of Theorem 5.1.
Define $X_{t}:=\left\|x_{t}-x_{\ast}\right\|^{2}$ and $\mathcal{F}_{t}$ to be the $\sigma$-algebra generated by $X_{0},\cdots,X_{t}$. Lemma 5.1 implies the following recursion, in the form of (42),
$\mathbf{E}[X_{t+1}|\mathcal{F}_{t}]$ $\leq(1-\alpha_{t}\lambda)^{2}X_{t}+\alpha_{t}^{2}B_{\ast}$
$=(1+\alpha_{t}^{2}\lambda^{2})X_{t}+\alpha_{t}^{2}B_{\ast}-2\alpha_{t}\lambda X_{t}$
$=\left(1+a_{t}\right)X_{t}+b_{t}-c_{t},$
where $a_{t}=\alpha_{t}^{2}\lambda^{2}$, $b_{t}=\alpha_{t}^{2}B_{\ast}$ and $c_{t}=2\alpha_{t}\lambda X_{t}$. Notice that $X_{t}$, $a_{t}$, $b_{t}$ and $c_{t}$ are all non-negative, and the square-summable assumption of $\alpha_{t}$ guarantees that
$\sum_{t=0}^{\infty}a_{t}=\sum_{t=0}^{\infty}\alpha_{t}^{2}\lambda^{2}<\infty,\qquad\sum_{t=0}^{\infty}b_{t}=\sum_{t=0}^{\infty}\alpha_{t}^{2}B_{\ast}<\infty.$
Therefore, all the assumptions in Lemma 5.2 are satisfied. We may now apply the lemma to conclude that $X_{t}=\left\|x_{t}-x_{\ast}\right\|^{2}\to X$ for some $X<\infty$ and
[…]
This is compatible with the assumption $\sum_{t=0}^{\infty}\alpha_{t}=\infty$ only if $\left\|x_{t}-x_{*}\right\|^{2}\to 0$ a.s. ∎
5.2.2 Rates of convergence
First, with a constant stepsize schedule, we have linear convergence to a neighborhood of $x_{*}$ as stated by the following corollary.
Corollary 5.1.
Consider the conceptual BCOS method (29) and fix the stepsize schedule $\alpha_{t}=\alpha$ where $\alpha$ satisfies $\alpha\lambda<1$. Then under Assumption A, we have
$\mathbf{E}\bigl{[}\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}\leq(1-\alpha\lambda)^{2t}\,\mathbf{E}\bigl{[}\left\|x_{0}-x_{\ast}\right\|^{2}\bigr{]}+\frac{\alpha^{2}B_{\ast}}{1-(1-\alpha\lambda)^{2}},$
where $B_{\ast}$ is defined in (38).
Proof.
Recursive application of Lemma 5.1 with $\alpha_{t}=\alpha$ yields
$\mathbf{E}\bigl{[}\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}$ $\leq(1-\alpha\lambda)^{2t}\,\mathbf{E}\bigl{[}\left\|x_{0}-x_{\ast}\right\|^{2}\bigr{]}+\sum_{s=0}^{t-1}(1-\alpha\lambda)^{2(t-s-1)}\alpha^{2}B_{\ast}$
$=(1-\alpha\lambda)^{2t}\,\mathbf{E}\left[\left\|x_{0}-x_{\ast}\right\|^{2}\right]+\sum_{s=0}^{t-1}(1-\alpha\lambda)^{2s}\alpha^{2}B_{\ast}$
$\leq(1-\alpha\lambda)^{2t}\,\mathbf{E}\left[\left\|x_{0}-x_{\ast}\right\|^{2}\right]+\frac{\alpha^{2}B_{\ast}}{1-(1-\alpha\lambda)^{2}},$
which is the desired result. ∎
With a diminishing stepsize schedule $\alpha_{t}$, we can show that $\mathbf{E}[\|x_{t}-x_{*}\|^{2}]\to 0$ at a sublinear rate, as given in the following theorem.
Theorem 5.2.
Consider the algorithm (29) with stepsize schedule $\alpha_{t}=\frac{\alpha}{t+1}$. Suppose the constant $\alpha$ satisfies $11/2<\alpha\lambda<$ 1 and Assumption A holds. Then we have for all $t\geq 0$,
$\mathbf{E}[\left\|x_{t}-x_{\ast}\right\|^{2}]\leq\frac{\alpha^{2}\left(\lambda^{2}\mathbf{E}\left[\left\|x_{0}-x_{\ast}\right\|^{2}\right]+(1+\pi^{2}/6)B_{\ast}\right)}{2\alpha\lambda-1}\frac{1}{t+1}+\mathcal{O}\left(\frac{1}{(t+1)^{2}}+\frac{1}{(t+1)^{2\alpha\lambda}}\right),$
where $B_{\ast}$ is defined in (38).
Our proof of Theorem 5.2 is based on Lemma 5.1 and a classical result of Chung chung1954 .
Lemma 5.3 (Chung’s lemma, (chung1954, , Lemma 1)).
Suppose that $\{X_{t}\}$ is a sequence of real numbers such that for all $t\geq 1$,
$X_{t+1}\leq\left(1-\frac{a}{t}\right)X_{t}+\frac{b}{t^{p+1}},$ (43)
where $a>p>0$ and $b>0$. Then
$X_{t}\leq\frac{b}{a-p}\frac{1}{t^{p}}+\mathcal{O}\left(\frac{1}{t^{p+1}}+\frac{1}{t^{a}}\right).$
In order to apply Chung’s lemma, we need to first derive an upper bound on $\mathbf{E}[\|x_{t}-x_{*}\|^{2}]$ for any $t\geq 0$, which will be incorporated into the constant $b$ in the lemma. To this end, we start with the inequality (37) and use the law of total expectation to get
$\mathbf{E}\bigl{[}\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}$ $=\mathbf{E}\!\left[\mathbf{E}_{t-1}[\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}\right]$
$\leq\mathbf{E}\!\left[(1-\alpha_{t-1}\lambda)^{2}\|x_{t-1}-x_{*}\|^{2}+\alpha_{t-1}^{2}B_{\ast}\right]$
$\leq\mathbf{E}\!\left[\|x_{t-1}-x_{*}\|^{2}\right]+\alpha_{t-1}^{2}B_{\ast}$
$\leq\mathbf{E}\bigl{[}\|x_{0}-x_{*}\|^{2}\bigr{]}+B_{\ast}\sum_{s=0}^{t-1}\alpha_{s}^{2},$ (44)
where the last term is bounded if $\sum_{t=1}^{\infty}\alpha_{t}^{2}<\infty$.
Proof of Theorem 5.2.
Taking total expectation on both sides of (37), we obtain
$\mathbf{E}\bigl{[}\left\|x_{t+1}-x_{\ast}\right\|^{2}\bigr{]}$ $\leq(1-\alpha_{t}\lambda)^{2}\mathbf{E}\bigl{[}\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}+\alpha_{t}^{2}B_{\ast}$
$=(1-2\alpha_{t}\lambda)\mathbf{E}\bigl{[}\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}+\alpha_{t}^{2}\left(\lambda^{2}\mathbf{E}\bigl{[}\|x_{t}-x_{*}\|^{2}\bigr{]}+B_{\ast}\right).$ (45)
We bound $\mathbf{E}[\|x_{t}-x_{*}\|^{2}]$ by substituting $\alpha_{t}=\frac{\alpha}{t+1}$ into (44), which leads to
$\mathbf{E}\bigl{[}\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}$ $\leq\mathbf{E}\bigl{[}\|x_{0}-x_{*}\|^{2}\bigr{]}+B_{\ast}\sum_{s=0}^{\infty}\frac{\alpha^{2}}{(s+1)^{2}}$
$=\mathbf{E}\bigl{[}\|x_{0}-x_{*}\|^{2}\bigr{]}+\frac{\pi^{2}}{6}\alpha^{2}B_{\ast}.$
Applying the above bound to the last term of (45) and setting $\alpha_{t}=\frac{\alpha}{t+1}$, we get
$\mathbf{E}\bigl{[}\left\|x_{t+1}-x_{\ast}\right\|^{2}\bigr{]}$ $\leq\left(1-\frac{2\alpha\lambda}{t+1}\right)\mathbf{E}\bigl{[}\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}+\frac{\alpha^{2}\!\left(\lambda^{2}\mathbf{E}\bigl{[}\|x_{0}-x_{*}\|^{2}\bigr{]}+(\pi^{2}/6)\alpha^{2}\lambda^{2}B_{\ast}+B_{\ast}\right)}{(t+1)^{2}}$
$\leq\left(1-\frac{2\alpha\lambda}{t+1}\right)\mathbf{E}\bigl{[}\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}+\frac{\alpha^{2}\!\left(\lambda^{2}\mathbf{E}\bigl{[}\|x_{0}-x_{*}\|^{2}\bigr{]}+(1+\pi^{2}/6)B_{\ast}\right)}{(t+1)^{2}},$
where in the second inequality we used the assumption $\alpha\lambda<1$.
Let $X_{t+1}=\mathbf{E}[\|x_{t}-x_{*}\|^{2}]$, $a=2\alpha\lambda$, $p=1$ and $b=\alpha^{2}\left(\lambda^{2}\mathbf{E}[\|x_{0}-x_{*}\|^{2}]+(1+\pi^{2}/6)B_{\ast}\right)$. Then the above inequality can be written as
$X_{t+2}\leq\left(1-\frac{a}{t+1}\right)X_{t+1}+\frac{b}{(t+1)^{p+1}},$
which is in the form of (43) except with the index shift $t\to t+1$. Applying Lemma 5.3 gives the desired result. ∎
Theorem 5.2 requires $\alpha\lambda>1/2$ to obtain the $\mathcal{O}(1/t)$ rate of convergence. We can remove this condition by working with the stepsize schedule $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ with $p\in(\frac{1}{2},1)$. In this case we have an asymptotic $\mathcal{O}(1/t^{p})$ rate, as stated in the following theorem.
Theorem 5.3.
Consider the algorithm (29) with stepsize schedule $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ where $p\in\left(\frac{1}{2},1\right)$ and the constant $\alpha$ satisfies $\alpha\lambda<1$. Then under Assumption A, we have
$\limsup_{t\to\infty}\,(t+1)^{p}\,\mathbf{E}\bigl{[}\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}\leq\frac{\alpha\left(\lambda^{2}\mathbf{E}\bigl{[}\left\|x_{0}-x_{\ast}\right\|^{2}\bigr{]}+(1+K_{p})B_{\ast}\right)}{2\lambda},$ (46)
where $K_{p}=\sum_{t=1}^{\infty}\frac{1}{(t+1)^{2p}}$ is finite for $p\in(\frac{1}{2},1)$.
The proof is based on another lemma of Chung chung1954 , restated below.
Lemma 5.4 ((chung1954, , Lemma 4)).
Suppose that $\{X_{t}\}$ is a sequence of real numbers such that for all $t\geq 1$,
$X_{t+1}\leq\left(1-\frac{a_{t}}{t^{p}}\right)X_{t}+\frac{b}{t^{q}},$ (47)
where $00<p<1,p<q,a_{t}\geq a>0,b>$ 0. Then
$\limsup_{t\to\infty}\,t^{q-p}X_{t}\leq\frac{b}{a}.$
Proof of Theorem 5.3.
Substituting $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ into (44), we obtain
$\mathbf{E}\bigl{[}\|x_{t}-x_{*}\|^{2}\bigr{]}\leq\mathbf{E}\bigl{[}\|x_{0}-x_{*}\|^{2}\bigr{]}+B_{\ast}\sum_{s=0}^{\infty}\frac{\alpha^{2}}{(t+1)^{2p}}=\mathbf{E}\bigl{[}\|x_{0}-x_{*}\|^{2}\bigr{]}+K_{p}\alpha^{2}B_{\ast}.$
Taking total expectation of both sides of (37) and plugging in $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$, we get
$\mathbf{E}\bigl{[}\left\|x_{t+1}-x_{\ast}\right\|^{2}\bigr{]}$ $\leq(1-\alpha_{t}\lambda)^{2}\mathbf{E}\bigl{[}\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}+\alpha_{t}^{2}B_{\ast}$
$=\left(1-2\alpha_{t}\lambda\right)\mathbf{E}[\left\|x_{t}-x_{\ast}\right\|^{2}]+\alpha_{t}^{2}\left(B_{\ast}+\lambda^{2}\mathbf{E}\bigl{[}\left\|x_{t}-x_{\ast}\right\|^{2}\bigr{]}\right)$
$\leq\left(1-\frac{2\alpha\lambda}{(t+1)^{p}}\right)\mathbf{E}[\left\|x_{t}-x_{\ast}\right\|^{2}]+\frac{\alpha^{2}\left(\lambda^{2}\mathbf{E}\bigl{[}\left\|x_{0}-x_{\ast}\right\|^{2}\bigr{]}+(1+K_{p})B_{\ast}\right)}{(t+1)^{2p}}.$
The desired result follows by applying Lemma 5.4 with the definitions $X_{t+1}=\mathbf{E}[\left\|x_{t}-x_{\ast}\right\|^{2}]$, $a_{t}=a=2\alpha\lambda$, $b=\alpha^{2}\left(\lambda^{2}\mathbf{E}[\left\|x_{0}-x_{\ast}\right\|^{2}]+(1+K_{p})B_{\ast}\right)$ and $q=2p$. ∎
5.2.3 Analysis for not using decoupled weight decay
Here we briefly explain the convergence analysis of BCOS in the case of $\lambda=0$, or in the case of $\lambda>0$ but without using decoupled weight decay. In other words, we consider
$x_{t+1}=x_{t}-\widetilde{\gamma}_{t}\odot d_{t}=x_{t}-\alpha_{t}\frac{d_{t}}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}.$ (48)
The aiming condition (32) becomes
$\biggl{\langle}x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\biggr{\rangle}\geq 0.$ (49)
In order to establish similar convergence guarantees as before, we need slightly stronger aiming conditions, as discussed below.
  1. 1.
    First, with $\lambda=0$, we have the following inequality in place of (37) (see Lemma 5.1):
    $\mathbf{E}_{t}\bigl{[}\|x_{t+1}-x_{*}\|^{2}\bigr{]}\leq\|x_{t}-x_{*}\|^{2}-\alpha_{t}\biggl{\langle}x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\biggr{\rangle}+\alpha_{t}^{2}n.$
    To have contraction with sufficiently small $\alpha_{t}$, we need the aiming condition (49) to hold with strict inequality for any $x_{t}\neq x_{*}$.
  2. 2.
    Almost sure convergence as in Theorem 5.4 requires a slightly stronger condition
    $\biggl{\langle}x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\biggr{\rangle}\geq\phi(\|x_{t}-x_{*}\|),$
    where $\phi$ is any strictly positive and continuous function.
  3. 3.
    To obtain sublinear rates of convergence as in Theorems 5.2 and 5.3, we need an even stronger aiming condition, specifically,
    $\biggl{\langle}x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\biggr{\rangle}\geq\mu\|x_{t}-x_{*}\|^{2},$ (50)
    with some $\mu>0$ for all $t\geq 0$.
Finally, we comment on the case of adding the regularization $(\lambda/2)\|x\|^{2}$ with $\lambda>0$ but without using decoupled weight-decay. Specifically, we compute $g_{t}=\nabla f(x_{t},\xi_{t})+\lambda x_{t}$ with $\lambda>0$ and apply the BCOS method (48) with $d_{t}$ being $g_{t}$ or its EMA. It is worth emphasizing that, in this case, the aiming condition (50) does not automatically hold with $\mu=\lambda$. This has to do with the coordinate-wise scaling of $d_{t}$ by $\sqrt{\mathbf{E}_{t}[d_{t}]}$. To see this, we consider the loss function $F(x)$ defined in (27), i.e.,
$F(x):=\mathbf{E}_{\xi}[f(x,\xi)]+\frac{\lambda}{2}\|x\|^{2},$
and assume that $\mathbf{E}_{\xi}[f(x,\xi)]$ is convex. We contrast the two cases discussed in Section 5.1.1.
  • With full coordinate block, it is straightforward to show that under the convex assumption, the aiming condition (35) is strengthened to
    $\langle x-x_{*},\,\nabla F(x)\rangle\geq\lambda\|x-x_{*}\|^{2}.$
  • However, the above argument does not extend to the case with coordinate-wise scaling. As an example, consider the special case of having uniform SiF across all coordinates. In this case, the corresponding aiming condition (36) cannot be strengthened to
    $\langle x-x_{*},\,\textup{sign}(\nabla F(x))\rangle\geq\lambda\|x-x_{*}\|^{2},$
    because the $\textup{sign}(\cdot)$ function is indifferent or insensitive to the curvature brought by the regularization $(\lambda/2)\|x\|^{2}$. In general, with coordinate-wise scaling and nonunifom SiF, the condition (50) may hold with some positive $\mu$ that is much smaller than $\lambda$.
The above discussion gives a plausible explanation to the empirical observation that decoupled weight decay works better for methods with coordinate-wise scaling (e.g., see Figure 3).
5.3 Analysis for practical BCOS
In this section, we provide convergence analysis for the practical BCOS method
$x_{t+1}=(1-\alpha_{t}\lambda)x_{t}-\alpha_{t}\frac{d_{t}}{\sqrt{v_{t}+\epsilon}},$ (51)
where $v_{t}$ is an online estimator of the conditional second moment $\mathbf{E}_{t}[d_{t}^{2}]$. As explained at the beginning of Section 5, our analysis is based on bounding the difference between the expected steps under the practical method and the conceptual method, in the form of (30). To proceed, we first decompose their difference into two terms, i.e.,
$\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}[d_{t}^{2}]}}-\mathbf{E}_{t}\biggl{[}\frac{d_{t}}{\sqrt{v_{t}\!+\!\epsilon}}\biggr{]}\right|\leq\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}[d_{t}^{2}]}}-\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]\!+\!\epsilon}}\right|+\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]\!+\!\epsilon}}-\mathbf{E}_{t}\biggl{[}\frac{d_{t}}{\sqrt{v_{t}\!+\!\epsilon}}\biggr{]}\right|,$ (52)
and then bound the two error terms on the right-hand side separately. Intuitively, the magnitudes of these two error terms are determined by the quality of the estimator $v_{t}$. We make the following assumption on the bias of $v_{t}$.
Assumption B (Bias of second-moment estimator).
There exists $\tau\in(0,1)$ and $\epsilon>0$ such that for all $t\geq 0$,
$\left|\mathbf{E}_{t}[v_{t}]-\mathbf{E}_{t}[d_{t}^{2}]\right|\leq\tau\mathbf{E}_{t}[d_{t}^{2}]+\epsilon.$ (53)
Here $\epsilon$ is an additive error bound that leaves room for $\mathbf{E}_{t}[v_{t}]$ when $\mathbf{E}_{t}[d_{t}^{2}]$ is very small. There is no loss of generality by restricting $\tau<1$ in the sense that we can always find an estimator that satisfies the assumption with sufficiently large $\epsilon$. As an extreme case, by setting $\epsilon\geq\|\mathbf{E}_{t}[d_{t}^{2}]\|_{\infty}$ for all $t$, the trivial estimator $v_{t}=0$ satisfies (53) for any $\tau\in(0,1)$.
5.3.1 Bounding the difference between conceptual and practical BCOS updates
The following lemma derives an upper bound on the first term on the right-hand side of (52). The proof is given in Appendix B.2.
Lemma 5.5.
Under Assumption B, it holds that:
$\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}-\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]+\epsilon}}\right|\leq\frac{\tau+\mathcal{O}(\tau^{2})}{2}\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|+\mathcal{O}(\epsilon).$ (54)
To bound the second term on the right-hand side of (52), we need the following lemma, whose proof is given in Appendix B.3.
Lemma 5.6.
Let $Y,Z$ be two random variables and $Z>0$ almost surely, then
$\mathbf{E}\left[\frac{Y}{\sqrt{Z}}\right]=$ $\frac{\mathbf{E}[Y]}{\sqrt{\mathbf{E}[Z]}}\left(1-\frac{\textup{Cov}(Y,Z)}{2\mathbf{E}[Y]\mathbf{E}[Z]}+\frac{3\textup{Var}(Z)}{8\mathbf{E}[Z]^{2}}+\mathcal{O}\left(\frac{\textup{Var}(Z)}{\mathbf{E}[Z]^{2}}\right)\right)$ (55)
where $\mathcal{O}\left(\frac{\textup{Var}(Z)}{\mathbf{E}[Z]^{2}}\right)$ includes higher-order statistical terms
$\frac{\mathbf{E}\bigl{[}(Y-\mathbf{E}[Y])(Z-\mathbf{E}[Z])^{p-1}\bigr{]}}{\mathbf{E}[Y]\mathbf{E}[Z]^{p-1}}\qquad\text{and}\qquad\frac{\mathbf{E}\bigl{[}(Z-\mathbf{E}[Z])^{p}\bigr{]}}{\mathbf{E}[Z]^{p}},$
for $p\geq 3$.
Applying Lemma 5.6 with $Y=d_{t}$ and $Z=v_{t}+\epsilon$, we get the coordinate-wise inequality
$\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]\!+\!\epsilon}}-\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}\!+\!\epsilon}}\right]\right|\leq\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]\!+\!\epsilon}}\right|\odot\left|\frac{-\textup{Cov}_{t}(d_{t},v_{t}\!+\!\epsilon)}{2\mathbf{E}_{t}[d_{t}]\mathbf{E}_{t}[v_{t}\!+\!\epsilon]}+\mathcal{O}\!\left(\frac{\textup{Var}_{t}(v_{t}\!+\!\epsilon)}{\mathbf{E}_{t}[v_{t}\!+\!\epsilon]^{2}}\right)\right|.$ (56)
In order to gain more intuition, we express the above bound using the conditional signal-to-noise ratio (SNR) of $d_{t}$ and $v_{t}+\epsilon$, defined as
$\mathrm{SNR}_{t}(d_{t})=\frac{\mathbf{E}_{t}[d_{t}]^{2}}{\textup{Var}_{t}(d_{t})},\qquad\mathrm{SNR}_{t}(v_{t}+\epsilon)=\frac{\mathbf{E}_{t}[v_{t}+\epsilon]^{2}}{\textup{Var}_{t}(v_{t}+\epsilon)}=\frac{\mathbf{E}_{t}[v_{t}+\epsilon]^{2}}{\textup{Var}_{t}(v_{t})}.$
With the definition of SNRs, the term containing $\textup{Cov}_{t}(d_{t},v_{t}+\epsilon)$ in (56) can be expressed as
$\frac{\textup{Cov}_{t}(d_{t},v_{t}+\epsilon)}{\mathbf{E}_{t}[d_{t}]\mathbf{E}_{t}[v_{t}+\epsilon]}$ $=\frac{\textup{Cov}_{t}(d_{t},v_{t})}{\mathbf{E}_{t}[d_{t}]\mathbf{E}_{t}[v_{t}+\epsilon]}$
$=\frac{\textup{Cov}_{t}(d_{t},v_{t})}{\sqrt{\textup{Var}_{t}(d_{t})\textup{Var}_{t}(v_{t})}}\sqrt{\frac{\textup{Var}_{t}(d_{t})\textup{Var}_{t}(v_{t})}{\mathbf{E}_{t}[d_{t}]^{2}\mathbf{E}_{t}[v_{t}+\epsilon]^{2}}}$
$=\textup{Corr}_{t}(d_{t},v_{t})\frac{1}{\sqrt{\mathrm{SNR}_{t}(d_{t})\mathrm{SNR}_{t}(v_{t}+\epsilon)}},$
where we use the definition of correlation of two random variables: $\textup{Corr}(Y,Z)=\frac{\textup{Cov}(Y,Z)}{\sqrt{\textup{Var}(Y)\textup{Var}(Z)}}$. Substituting the expressions using SNRs into (56), we obtain
$\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]\!+\!\epsilon}}-\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}\!+\!\epsilon}}\right]\right|\leq\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]\!+\!\epsilon}}\right|\odot\left|\frac{-(1/2)\textup{Corr}_{t}(d_{t},v_{t})}{\sqrt{\mathrm{SNR}_{t}(d_{t})\mathrm{SNR}_{t}(v_{t}\!+\!\epsilon)}}+\mathcal{O}\!\left(\frac{1}{\mathrm{SNR}_{t}(v_{t}\!+\!\epsilon)}\right)\right|.$
We are ready to bound the difference between the expected updates of the practical BCOS and the conceptual BCOS methods. Specifically, bounding the first term on the right-hand side of (52) using Lemma 5.5 and the second term using the last inequality above, we obtain
$\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}-\mathbf{E}_{t}\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]\right|$ $\leq\frac{\tau+\mathcal{O}(\tau^{2})}{2}\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|+\mathcal{O}(\epsilon)$
$\quad+\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{v_{t}+\epsilon}}\right|\odot\left|\frac{-(1/2)\textup{Corr}_{t}(d_{t},v_{t})}{\sqrt{\mathrm{SNR}_{t}(d_{t})\mathrm{SNR}_{t}(v_{t}+\epsilon)}}+\mathcal{O}\!\left(\frac{1}{\mathrm{SNR}_{t}(v_{t}+\epsilon)}\right)\right|.$
We can further merge the two terms on the right-hand side by bounding $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{v_{t}+\epsilon}}\right|$ in terms of $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|$. To this end, we use Lemma 5.5 to obtain
$\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]+\epsilon}}\right|\leq\left(1+\frac{\tau+\mathcal{O}(\tau^{2})}{2}\right)\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|+\mathcal{O}(\epsilon).$
Combining the two inequalities above, we arrive at the following result.
Lemma 5.7.
Under Assumptions B, we have for all $t\geq 0$,
$\left|\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]-\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|$ $\leq\sigma_{t}\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|+\mathcal{O}(\epsilon),$
where
$\sigma_{t}=\frac{\tau\!+\!\mathcal{O}(\tau^{2})}{2}+\left(1+\frac{\tau\!+\!\mathcal{O}(\tau^{2})}{2}\right)\left\|\frac{-(1/2)\textup{Corr}_{t}(d_{t},v_{t}\!)}{\sqrt{\mathrm{SNR}_{t}(d_{t})\mathrm{SNR}_{t}(v_{t}\!+\!\epsilon)}}+\mathcal{O}\!\left(\frac{1}{\mathrm{SNR}_{t}(v_{t}\!+\!\epsilon)}\right)\right\|_{\infty}\!\!.$ (57)
5.3.2 Almost sure convergence of practical BCOS
The following lemma is a variant of the aiming condition with the practical BCOS update.
Lemma 5.8.
Under Assumptions B, it holds that
$\left\langle x_{t}-x_{\ast},\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]\right\rangle\geq$ $\left\langle x_{t}-x_{\ast},\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\rangle-\left\|x_{t}-x_{\ast}\right\|_{1}\bigl{(}\sigma_{t}+\mathcal{O}(\epsilon)\bigr{)},$ (58)
where $\sigma_{t}$ is given by (57).
Proof.
Adding and subtracting the term $\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}$ from the inner product, we obtain:
$\left\langle x_{t}-x_{\ast},\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]\right\rangle$
[…]
where the first inequality is due to Hölder’s inequality and the last inequality is due to Lemma 5.7. In addition, we notice that
$\biggl{\|}\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\biggr{\|}_{\infty}=\left\|\sqrt{\rho_{t}}\right\|_{\infty}\leq 1,$ (59)
which concludes the proof. ∎
Theorem 5.4 (Almost sure convergence of BCOSW).
Consider the practical BCOSW method (28). Suppose that Assumptions A and B hold, $d_{t}$ is bounded almost surely, and $\{\alpha_{t}\}$ satisfies (41). Then there exists $\delta>0$ such that we have with probability one,
$\limsup_{t\to\infty}\|x_{t}-x_{*}\|^{2}\leq\delta^{2}.$ (60)
More specifically, $\delta$ is the smallest constant such that $\|x-x_{*}\|\geq\delta$ implies
$2\|x-x_{*}\|_{1}\bigl{(}\sigma_{t}+\mathcal{O}(\epsilon)\bigr{)}\leq\lambda\|x-x_{*}\|^{2},\qquad\forall\,t\geq 0,$ (61)
where $\sigma_{t}$ is given by (57).
The proof of Theorem 5.4 is based on a classical result of stochastic approximation due to Dvoretzky DermanSacks1959 . Here we present an extension that best fits our purpose.
Theorem 5.5 (An extension of Dvoretzky’s Theorem).
Let $(\Omega=\{\omega\},\mathcal{F},P)$ be a probability space. Let $\{x_{t}\}$ and $\{y_{t}\}$ be sequences of random variables such that, for all $t\geq 0$,
$x_{t+1}(\omega)=T_{t}\bigl{(}x_{0}(\omega),\ldots,x_{t}(\omega)\bigr{)}+y_{t}(\omega),$ (62)
where the transformation $T_{t}$ satisfy, for any $x_{0},\ldots,x_{t}\in\mathbf{R}^{n}$,
$\bigl{\|}T_{t}(x_{0},\ldots,x_{t})-x_{*}\|^{2}\leq\max\bigl{\{}a_{t},\,(1+b_{t})\|x_{t}-x_{*}\|^{2}-c_{t}+h_{t}\bigr{\}},$ (63)
and the sequences $\{a_{t}\}$, $\{b_{t}\}$, $\{c_{t}\}$ and $\{h_{t}\}$ are non-negative and satisfy
$\lim_{t\to\infty}a_{t}=a_{\infty},\qquad\sum_{t=0}^{\infty}b_{t}<\infty,\qquad\sum_{t=0}^{\infty}c_{t}=\infty,\qquad\sum_{t=1}^{\infty}h_{t}<\infty.$ (64)
In addition, suppose the following conditions hold with probability one:
$\mathbf{E}[\|x_{0}\|^{2}]<\infty,\qquad\sum_{t=0}^{\infty}\mathbf{E}[\|y_{t}\|^{2}]<\infty,\qquad\mathbf{E}\bigl{[}y_{t}|x_{0},\ldots,x_{t}\bigr{]}=0\quad\forall\,t\geq 0.$ (65)
Then we have with probability one,
$\limsup_{t\to\infty}\|x_{t}-x_{*}\|^{2}\leq a_{\infty}.$
Remark.
There are many extensions of Dvoretzky’s original results Dvoretzky1956 . Theorem 5.5 is a minor variation of Venter (Venter1966, , Theorem 1). More concretely,
  • Theorem 1 of Venter Venter1966 has the sequence $\{a_{t}\}$ being a constant sequence, i.e., $a_{t}=a_{\infty}$ for all $t\geq 0$. The extension to a non-constant sequence $\{a_{t}\}$ is outlined in the original work of Dvoretzky Dvoretzky1956 and admits a simple proof due to Derman and Sacks DermanSacks1959 .
  • Theorem 1 of Venter Venter1966 does not include the sequence $\{h_{t}\}$. The extension with $\sum_{t=0}^{\infty}h_{t}<\infty$ is straightforward based on a simple argument of Dvoretzky Dvoretzky1956 .
  • More generally, the sequences $\{a_{t}\}$, $\{b_{t}\}$, $\{c_{t}\}$, $\{h_{t}\}$ can be non-negative measurable functions of $x_{0},\ldots,x_{t}$, and the conclusion of Theorem 5.5 holds if $a_{\infty}$ is an upper bound on $\limsup_{t\to\infty}a_{t}(x_{0},\ldots,x_{t})$ uniformly for all sequences $\{x_{t}\}$ DermanSacks1959 ; ROBBINS1971 .
Proof of Theorem 5.5.
We first write the practical BCOSW algorithm in the form of (62):
$x_{t+1}$ $=(1-\alpha_{t}\lambda)x_{t}-\alpha_{t}\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}$
$=(1-\alpha_{t}\lambda)x_{t}-\alpha_{t}\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]+\alpha_{t}\left(\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]-\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right).$
Thus we have $x_{t+1}=T_{t}(x_{0},\ldots,x_{t})+y_{t}$ with
$T_{t}(x_{0},\ldots,x_{t})$ $=(1-\alpha_{t}\lambda)x_{t}-\alpha_{t}\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right],$
$y_{t}$ $=\alpha_{t}\left(\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]-\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right).$
Apparently $\mathbf{E}_{t}[y_{t}]=\mathbf{E}[y_{t}|x_{0},\ldots,x_{t}]=0$. We also have $\sum_{t=0}^{\infty}\mathbf{E}[\|y_{t}\|^{2}]<\infty$ due to the assumptions that $\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty$ and $d_{t}$ is bounded almost surely. Therefore the conditions in (65) are all satisfied.
The squared distance between $T_{t}(x_{0},\ldots,x_{t})$ and $x_{*}$ admits the following expansion:
$\bigl{\|}T_{t}(x_{0},\ldots,x_{t})-x_{*}\bigr{\|}^{2}$ $=\left\|(1-\alpha_{t}\lambda)x_{t}-\alpha_{t}\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]-x_{*}\right\|^{2}$
$=\left\|(1-\alpha_{t}\lambda)(x_{t}-x_{*})-\alpha_{t}\left(\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]+\lambda x_{*}\right)\right\|^{2}$
$=(1-\alpha_{t}\lambda)^{2}\|x_{t}-x_{*}\|^{2}-2\alpha_{t}(1-\alpha_{t}\lambda)\left\langle x_{t}-x_{*},\,\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]+\lambda x_{*}\right\rangle$
$\quad+\alpha_{t}^{2}\left\|\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]+\lambda x_{*}\right\|^{2}.$
Using Lemma 5.8 and then the aiming condition (39), we obtain
$\left\langle x_{t}-x_{\ast},\,\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]\!+\!\lambda x_{*}\right\rangle$ $\geq\left\langle x_{t}-x_{\ast},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\!\lambda x_{*}\right\rangle-\|x_{t}-x_{*}\|_{1}\left(\sigma_{t}+\mathcal{O}(\epsilon)\right)$
$\geq-\|x_{t}-x_{*}\|_{1}\left(\sigma_{t}+\mathcal{O}(\epsilon)\right).$
By the assumption that $d_{t}$ is bounded almost surely, there exists a constant $B$ such that
$\left\|\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]+\lambda x_{*}\right\|^{2}\leq B,\qquad\forall\,t\geq 0.$
Together with $10<1-\alpha_{t}\lambda<$ 1, we conclude that
$\bigl{\|}T_{t}(x_{0},\ldots,x_{t})-x_{*}\bigr{\|}^{2}$ $\leq(1-\alpha_{t}\lambda)^{2}\left\|x_{t}-x_{\ast}\right\|^{2}+2\alpha_{t}(1-\alpha_{t}\lambda)\|x_{t}-x_{*}\|_{1}(\sigma_{t}+\mathcal{O}(\epsilon))+\alpha_{t}^{2}B$
$\leq(1-\alpha_{t}\lambda)^{2}\left\|x_{t}-x_{\ast}\right\|^{2}+2\alpha_{t}\|x_{t}-x_{*}\|_{1}(\sigma_{t}+\mathcal{O}(\epsilon))+\alpha_{t}^{2}B$
$=(1+\alpha_{t}^{2}\lambda^{2})\left\|x_{t}-x_{\ast}\right\|^{2}-2\alpha_{t}\lambda\|x_{t}-x_{*}\|^{2}+2\alpha_{t}\|x_{t}-x_{*}\|_{1}(\sigma_{t}+\mathcal{O}(\epsilon))$
$\quad~{}+\alpha_{t}^{2}B$
$=(1+\alpha_{t}^{2}\lambda^{2})\left\|x_{t}-x_{\ast}\right\|^{2}+\alpha_{t}\left(2\|x_{t}-x_{*}\|_{1}(\sigma_{t}+\mathcal{O}(\epsilon))-\lambda\|x_{t}-x_{*}\|^{2}\right)$
$\quad~{}-\alpha_{t}\lambda\|x_{t}-x_{*}\|^{2}+\alpha_{t}^{2}B.$ (66)
We observe that there exist $\delta>0$ such that for all $t\geq 0$,
$\|x_{t}-x_{*}\|\geq\delta\quad\Longrightarrow\quad 2\|x_{t}-x_{*}\|_{1}(\sigma_{t}+\mathcal{O}(\epsilon))-\lambda\|x_{t}-x_{*}\|^{2}\leq 0.$
Therefore, when $\|x_{t}-x_{*}\|\geq\delta$, we deduce from (66) that
$\left\|T_{t}(x_{0},\ldots,x_{t})-x_{\ast}\right\|^{2}\leq(1+\alpha_{t}^{2}\lambda^{2})\|x_{t}-x_{*}\|^{2}-\alpha_{t}\lambda\|x_{t}-x_{*}\|^{2}+\alpha_{t}^{2}B.$
Otherwise, when $\|x_{t}-x_{*}\|\leq\delta$, we have
$\left\|T_{t}(x_{0},\ldots,x_{t})-x_{\ast}\right\|^{2}\leq(1-\alpha_{t}\lambda)^{2}\delta^{2}+2\alpha_{t}(1-\alpha_{t}\lambda)\sqrt{n}\delta(\sigma_{t}+\mathcal{O}(\epsilon))+\alpha_{t}^{2}B,$
where we used $\|x_{t}-x_{*}\|_{1}\leq\sqrt{n}\|x_{t}-x_{*}\|\leq\sqrt{n}\delta$. With the definition
$a_{t}=(1-\alpha_{t}\lambda)^{2}\delta^{2}+2\alpha_{t}(1-\alpha_{t}\lambda)\sqrt{n}\delta(\sigma_{t}+\mathcal{O}(\epsilon))+\alpha_{t}^{2}B,$ (67)
we can combine the above two cases as
$\left\|T_{t}(x_{0},\ldots,x_{t})-x_{\ast}\right\|^{2}\leq\max\bigl{\{}a_{t},\,(1+\alpha_{t}^{2}\lambda^{2})\|x_{t}-x_{*}\|^{2}-\alpha_{t}\lambda\|x_{t}-x_{*}\|^{2}+\alpha_{t}^{2}B\bigr{\}}.$
With the additional definitions of
$b_{t}=\alpha_{t}^{2}\lambda^{2},\qquad c_{t}=\alpha_{t}\lambda\|x_{t}-x_{*}\|^{2},\qquad h_{t}=\alpha_{t}^{2}B,$
we arrive at the key inequality (63).
In order to apply Theorem 5.5, we are left to check the conditions in (64). Using the assumption $\alpha_{t}\to 0$, the definition of $a_{t}$ in (67) implies that $\lim_{t\to\infty}a_{t}=\delta^{2}$. The conditions on $\{b_{t}\}$ and $\{h_{t}\}$ are satisfied due to the assumption $\sum_{t=0}^{\infty}\alpha_{t}^{2}\leq\infty$. For $\{c_{t}\}$, if $\sum_{t=0}^{\infty}c_{t}<\infty$, then we must have $\|x_{t}-x_{\star}\|^{2}\to 0$ almost surely because of the assumption $\sum_{t=0}^{\infty}\alpha_{t}=\infty$, and the conclusion of the theorem holds trivially. Otherwise, $\sum_{t=0}^{\infty}c_{t}=\infty$ allows all the conditions in (64) to hold, so we can invoke Theorem 5.5 to finish the proof. ∎
Following the last remark on Dvoretzky’s theorem on page 5.3.2, the radius of the neighborhood $\delta$ can be refined as follows. Let $\delta_{t}$ be the smallest constant such that $\|x_{t}-x_{*}\|\geq\delta_{t}$ implies
$2\|x_{t}-x_{*}\|_{1}\bigl{(}c_{t}+\mathcal{O}(\epsilon)\bigr{)}\leq\lambda\|x_{t}-x_{*}\|^{2},$
where $\sigma_{t}$ is given by (57). Then we can set $\delta=\limsup_{t\to\infty}\delta_{t}$ in (60).
5.4 Bias and variance trade-off
Theorem 5.4 states that the sequence $\{x_{t}\}$ generated by BCOSW converges almost surely to a neighborhood of $x_{*}$, and the radius of the neighborhood is determined by the magnitude of $\sigma_{t}$ and $\epsilon$. Specifically, smaller $\sigma_{t}$ and $\epsilon$ ensure a smaller neighborhood of convergence. According to (57), the magnitude of $\sigma_{t}$ depends on both the bias and variance of the estimator $v_{t}$. In this section, we examine the delicacy of this bias-variance trade-off and illustrate its effects through several examples.
For convenience, we repeat the definition of $\sigma_{t}$ here for quick reference:
$\sigma_{t}=\frac{\tau\!+\!\mathcal{O}(\tau^{2})}{2}+\left(1+\frac{\tau\!+\!\mathcal{O}(\tau^{2})}{2}\right)\left\|\frac{-(1/2)\textup{Corr}_{t}(d_{t},v_{t})}{\sqrt{\mathrm{SNR}_{t}(d_{t})\mathrm{SNR}_{t}(v_{t}\!+\!\epsilon)}}+\mathcal{O}\!\left(\frac{1}{\mathrm{SNR}_{t}(v_{t}\!+\!\epsilon)}\right)\right\|_{\infty}\!\!.$ (68)
Apparently, in order to make $\sigma_{t}$ small, it is desirable for the random variable $v_{t}$ (as an estimator of $\mathbf{E}_{t}[d_{t}^{2}]$) to possess the following properties.
  • Low bias. This means that Assumption B holds with small $\tau$ and $\epsilon$.
  • Low variance, which naturally leads to high SNR. Notice that using a larger $\epsilon$ increases $\mathrm{SNR}_{t}(v_{t}+\epsilon)=\mathbf{E}_{t}[v_{t}+\epsilon]^{2}/\textup{Var}_{t}(v_{t})$, leading to a smaller $\sigma_{t}$. However, it also increases the $\mathcal{O}(\epsilon)$ term in (58), which affects the radius of neighborhood through (61).
In practice, it is hard to construct estimators with low bias and low variance simultaneously. Below we give several examples to illustrate the trade-off in algorithm design.
  1. 1.
    Using $d_{t}^{2}$ as the estimator. The simplest estimator of $\mathbf{E}_{t}[d_{t}^{2}]$ is $d_{t}^{2}$ itself. Setting $v_{t}=d_{t}^{2}$ exhibiting low bias and high variance, specifically,
    $\mathrm{Bias}=\left|\mathbf{E}_{t}[v_{t}]-\mathbf{E}_{t}[d_{t}^{2}]\right|=0,\qquad\mathrm{Variance}=\textup{Var}_{t}(v_{t})=\textup{Var}_{t}(d_{t}^{2}).$
    In this case, Assumption B is satisfied with $\tau=0$ and $\epsilon=0$. Then we have
    $x_{t+1}~{}=~{}x_{t}-\alpha_{t}\frac{d_{t}}{\sqrt{d_{t}^{2}+\epsilon}}~{}=~{}x_{t}-\alpha_{t}\textup{sign}(d_{t}),$
    which corresponds to the sign-gradient method (21) or the sign-momentum method (22), depending on the choice of $d_{t}$. Among them, sign-momentum is preferred because $\textup{Var}_{t}(m_{t}^{2})$ is usually much smaller than $\textup{Var}_{t}(g_{t}^{2})$.
  2. 2.
    Using a constant estimator across coordinates. Setting $v_{t}=c\mathbf{1}_{n}$ for some constant $c>0$ exhibits high bias and low variance, specifically,
    $\mathrm{Bias}=\left|\mathbf{E}_{t}[v_{t}]-\mathbf{E}_{t}[d_{t}^{2}]\right|=\left|c\mathbf{1}_{n}-\mathbf{E}_{t}[d_{t}^{2}]\right|,\qquad\mathrm{Variance}=\textup{Var}_{t}(v_{t})=0.$
    In this case, we need large $\tau$ and $\epsilon$ for Assumption B to hold. Correspondingly
    $x_{t+1}=x_{t}-\alpha_{t}\frac{d_{t}}{\sqrt{c\mathbf{1}_{n}+\epsilon}}=x_{t}-\alpha_{t}^{\prime}d_{t},$
    where $\alpha_{t}^{\prime}=\alpha_{t}/\sqrt{c+\epsilon}$ is a scalar stepsize. This is equivalent to the classical SGD with or without momentum, depending on $d_{t}=m_{t}$ or $d_{t}=g_{t}$ respectively.
  3. 3.
    Using EMA of $d_{t}^{2}$ as the estimator. EMA is widely adopted as an online estimator to approximate the mean of a random variable. For example, $v_{t}=\beta_{1}v_{t-1}+(1-\beta)d_{t}^{2}$ is used in BCOS-g (Algorithm 1) and BCOS-m (Algorithm 2) to approximate $\mathbf{E}_{t}[d_{t}^{2}]$, with $d_{t}=g_{t}$ and $d_{t}=m_{t}$ respectively. The bias can be expressed as
    $\mathrm{Bias}$ $=\left|\mathbf{E}_{t}[v_{t}]-\mathbf{E}_{t}[d_{t}^{2}]\right|=\left|\mathbf{E}_{t}\biggl{[}\sum_{k=1}^{t}(1-\beta)\beta^{t-k}d_{k}^{2}\biggr{]}-\mathbf{E}_{t}[d_{t}^{2}]\right|.$
    It is possible to derive a sensible bound on the bias with a smoothness assumption on the loss function, but we do not delve into it here. On the other hand, we have a simple expression for its variance (derivation given in Appendix B.4):
    $\textup{Var}_{t}(v_{t})$ $=(1-\beta)^{2}\textup{Var}_{t}(d_{t}^{2}),$
    which is much smaller than $\textup{Var}_{t}(d_{t}^{2})$ obtained by simply using $d_{t}^{2}$ as the estimator.
  4. 4.
    The Adam estimator. Adam uses stochastic momentum as the search direction, i.e., $d_{t}=m_{t}=\beta_{1}d_{t-1}+(1-\beta_{1})g_{t}$. However, it uses the EMA of $g_{t}^{2}$, rather than the EMA of $m_{t}^{2}$ (as in the previous example), to estimate $\mathbf{E}_{t}[m_{t}^{2}]$. This causes a mismatch from the perspective of BCOS, and it is compensated by using a larger smoothing factor $\beta_{2}$ in the estimator $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2}$. Its bias can be expressed as
    $\mathrm{Bias}=\left|\mathbf{E}_{t}\biggl{[}\sum_{k=1}^{t}(1-\beta_{2})\beta_{2}^{t-k}g_{k}^{2}\biggr{]}-\mathbf{E}_{t}\left[\biggl{(}\sum_{k=1}^{t}(1-\beta_{1})\beta_{1}^{t-k}g_{k}\biggr{)}^{\!2}\right]\right|.$
    Again, deriving a sensible upper bound on the bias can be quite involved and we do not pursue it here. As for the variance, we get (see Appendix B.4)
    $\textup{Var}_{t}(v_{t})$ $=(1-\beta_{2})^{2}\textup{Var}_{t}\left(g_{t}^{2}\right).$
    If we choose $\beta_{2}=1-(1-\beta_{1})^{2}$, then $\textup{Var}_{t}(v_{t})=(1-\beta_{1})^{4}\textup{Var}_{t}(g_{t}^{2})$. Therefore, the Adam estimator exhibits a very low variance.
  5. 5.
    The conditional estimator of BCOS-c. The BCOS-c method (Algorithm 3) also uses $m_{t}$ as the search direction. Instead of relying on EMA, it uses a conditional estimator for $\mathbf{E}_{t}[m_{t}^{2}]$, i.e., $v_{t}=(1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$, where $\beta$ is the smoothing factor for computing $m_{t}$. We have simple expressions for both the bias and the variance:
    $\mathrm{Bias_{t}}(v_{t})$ $=2\beta(1-\beta)\left|m_{t-1}\odot\bigl{(}m_{t-1}-\mathbf{E}_{t}[g_{t}]\bigr{)}\right|,$
    $\textup{Var}_{t}(v_{t})$ $=(1-\beta)^{4}\,\textup{Var}_{t}(g_{t}^{2}).$
    Notice that $|m_{t-1}-\mathbf{E}_{t}[g_{t}]|$ is the bias of using $m_{t-1}$ to estimate $\mathbf{E}_{t}[d_{t}]$, and the bias of $v_{t}$ can be much smaller than that. Meanwhile, the variance of BCOS-c is the same as Adam when setting $\beta_{2}$ in Adam as $\beta_{2}=1-(1-\beta)^{2}$.
In addition to the bias-variance trade-off illustrated above, the expression of $\sigma_{t}$ in (68) also reveals the following tips for algorithm design:
  • High SNR of $d_{t}$. This is consistent with the empirical observation that using the momentum $m_{t}$ as search direction is often better than using the stochastic gradient $g_{t}$, because $m_{t}$ has much smaller variance, thus higher SNR, than $g_{t}$.
  • Positive correlation between $d_{t}$ and $v_{t}$. The negative sign before $\textup{Corr}_{t}(d_{t},v_{t})$ implies that a positive correlation can lead to a smaller $\sigma_{t}$. Intuitively, a negative correlation between $d_{t}$ and $v_{t}$ may cause large fluctuations of the ratio $d_{t}/\sqrt{v_{t}+\epsilon}$.
In summary, our theory captures several essential trade-offs that match intuitions and empirical observations.
6 Conclusion
BCOS is a family of stochastic approximation methods that exploit the flexibility of using different block-coordinate stepsizes. Rather than using sophisticated ideas from optimization, such as preconditioning, it is based on the simple idea of minimizing the distance from the next iterate to an optimal point. While the optimal stepsizes are not computable, we make several simplifications and focus on constructing efficient statistical estimators (for the second moment of the search direction) in determining the stepsizes. In particular, by leveraging a simple conditional estimator, we derive variants of BCOS that achieve competitive performance against AdamW but require less memory and fewer hyperparameters.
Our convergence analysis builds upon several classical results from the stochastic approximation literature. In particular, we extend the classical aiming condition towards a target point to account for coordinate-wise stepsizes and decoupled weight-decay (interpreted as $L_{2}$-regularization). For the conceptual BCOS method, which assumes exact second moment, we establish almost-sure convergence and $\mathcal{O}(1/t)$ rate of convergence. For practical BCOS methods with various second-moment estimators, we establish almost-sure convergence to a neighborhood of the target point, where the radius of the neighborhood is determined by the bias and variance of the second-moment estimator. This framework covers convergence analysis for a broad family of methods, including the sign stochastic gradient/momentum method, RMSProp and Adam(W). In addition, it provides novel insights that can guide the development of new algorithms.
Acknowledgments
We thank Zeyuan Allen-Zhu for helping us with running experiments on GPT2, especially with rotary embedding and distributed data parallelism in PyTorch. We are grateful to Lisa Jin for her codebase that made our lives much easier on experimenting with ResNet and vision Transformers. We also thank Yann Olivier for careful reading of an early version of the draft and providing feedback. Last but not least, we extend our gratitude to Damek Davis, Anil Damle, Adrian Lewis, James Renegar and Katya Scheinberg for their comments and insights that helped us clarify and refine the theoretical underpinnings of the paper.
Appendix A PyTorch implementation of BCOS
The following code is also available at https://github.com/facebookresearch/bcos.
1import torch
2from torch.optim import Optimizer
3
4class BCOS(Optimizer):
5 def __init__(self, params, lr, beta=0.9, eps=1e-6,
6 weight_decay=0.1, mode='c', decouple_wd=True):
7
8 defaults = dict(lr=lr, beta=beta, eps=eps, wd=weight_decay)
9 super().__init__(params, defaults)
10
11 if mode not in ['g', 'm', 'c']:
12 raise ValueError(f"BCOS mode {mode} not supported")
13 self.mode = mode
14 self.decouple_wd = decouple_wd # True for BCOSW
15
16 def step(self, closure = None):
17
18 for group in self.param_groups:
19 lr = group["lr"]
20 beta = group["beta"]
21 eps = group["eps"]
22 wd = group["wd"]
23
24 for p in group["params"]:
25 if not p.requires_grad:
26 continue
27
28 state = self.state[p]
29 g = p.grad
30
31 # initialize optimizer states for specific modes
32 if self.mode in ['m', 'c'] and 'm' not in state:
33 state['m'] = g.detach().clone()
34 if self.mode in ['g', 'm'] and 'v' not in state:
35 state['v'] = g.detach().square()
36
37 # decoupled weight decay or absorb in gradient
38 if self.decouple_wd: # p := (1 - lr * wd) * p
39 p.data.mul_(1 - lr * wd)
40 else: # g := g + wd * p
41 g.data.add_(p.data, alpha = wd)
42
43 if self.mode in ['m', 'c']:
44 m = state['m']
45 if self.mode == 'c': # conditional estimator
46 beta_v = 1 - (1 - beta)**2
47 g2 = g.detach().square()
48 v = beta_v * m.square() + (1 - beta_v) * g2
49 # update momentum
50 m.mul_(beta).add_(g.detach(), alpha=1 - beta)
51 d = m
52 else:
53 d = g.detach()
54
55 if self.mode in ['g', 'm']: # EMA estimator
56 v = state['v']
57 v.mul_(beta).add_(d.square(), alpha=1 - beta)
58
59 # BCOS update: p := p - lr * (d / (sqrt(v) + eps))
60 p.data.add_(d.div(v.sqrt() + eps), alpha= - lr)
Appendix B Appendix
B.1 Aiming versus convexity
To illustrate the fundamental differences between the coordinate-wise aiming condition (36) and the standard convexity condition (35), we provide the following two counterexamples, each satisfying one condition while failing the other:
  • Aiming but not convex: Let $f(x):=\log(x)$ with the optimal solution $x_{\ast}=0$. On the domain of $\mathbf{R}_{+}$, the gradient is $f^{\prime}(x)=\frac{1}{x}$, and thus $\textup{sign}(f^{\prime}(x))=1$ for all $x>0$. Consequently, for any $x\in\mathbf{R}_{+}$, we have
    $\langle x-x_{\ast},\textup{sign}(\nabla f(x))\rangle=x\geq 0,$
    satisfying the aiming condition (36). However, $\log(x)$ is a concave function, thus failing the convex inequality (35).
  • Convex but not aiming: Consider the quadratic function $f:\mathbf{R}^{2}\to\mathbf{R},f(x)=\frac{1}{2}x^{T}Ax$. Choose the coefficient matrix as
    $A=\begin{pmatrix}1\\ -2\end{pmatrix}\begin{pmatrix}1\\ -2\end{pmatrix}^{T}=\begin{bmatrix}1&-2\\ -2&4\end{bmatrix}\succeq 0.$
    Since $A$ is positive semidefinite, the function $f$ is convex and attains its minimum at $x_{\ast}=\mathbf{0}$. The gradient of $f$ is
    $\nabla f(x)=Ax=\begin{bmatrix}x_{1}-2x_{2}\\ -2x_{1}+4x_{2}.\end{bmatrix}.$
    Evaluating the aiming condition (36) at $x=(1.5,1)^{T}$, we get
    $\langle x-x_{\ast},\textup{sign}(\nabla f(x))\rangle=1.5\times\textup{sign}(-0.5)+1\times\textup{sign}(1)=1.5\times(-1)+1\times 1=-0.5\leq 0.$
    Thus, the aiming condition (36) does not hold at this point even though $f$ is convex.
B.2 Proof of Lemma 5.5
Proof.
The proof leverages the Taylor expansion of $\phi(y):=1/\sqrt{y}$. Specifically, we have
$\phi(y+\delta)=\sum_{p=0}^{\infty}\frac{\phi^{(p)}(y)}{p!}\delta^{p},$
where
$\phi^{(p)}=(-1)^{p}\left(\frac{1}{2}\cdot\frac{3}{2}\cdots\left(p-\frac{1}{2}\right)\right)y^{-\frac{2p+1}{2}}=(-1)^{p}\frac{(2p)!}{4^{p}p!}y^{-\frac{2p+1}{2}}.$
Taylor expansion (element-wise) at $y=\mathbf{E}_{t}[d_{t}^{2}]$ with $\delta=\mathbf{E}_{t}[v_{t}]+\epsilon-\mathbf{E}_{t}[d_{t}^{2}]$ yields the following approximation:
$\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}-\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]+\epsilon}}\right|$ $=\left|\mathbf{E}_{t}[d_{t}](\phi(y)-\phi(y+\delta))\right|=\left|\mathbf{E}_{t}[d_{t}]\left(\sum_{p=1}^{\infty}\frac{\phi^{(p)}(y)}{p!}\delta^{p}\right)\right|$
$=\left|\mathbf{E}_{t}[d_{t}]\left(\sum_{p=1}^{\infty}(-1)^{p}\frac{(2p)!}{4^{p}(p!)^{2}}\mathbf{E}_{t}[d_{t}^{2}]^{-\frac{2p+1}{2}}(\mathbf{E}_{t}[v_{t}]+\epsilon-\mathbf{E}_{t}[d_{t}^{2}])^{p}\right)\right|$
$\leq\left|\mathbf{E}_{t}[d_{t}]\left(\sum_{p=1}^{\infty}\frac{(2p)!}{4^{p}(p!)^{2}}\mathbf{E}_{t}[d_{t}^{2}]^{-\frac{2p+1}{2}}(\tau\mathbf{E}_{t}[d_{t}^{2}]+2\epsilon)^{p}\right)\right|$ (69)
$=\left|\mathbf{E}_{t}[d_{t}]\left(\sum_{p=1}^{\infty}\frac{(2p)!}{4^{p}(p!)^{2}}\mathbf{E}_{t}[d_{t}^{2}]^{-\frac{2p+1}{2}}\tau^{p}\mathbf{E}_{t}[d_{t}^{2}]^{p}+\mathcal{O}(\epsilon)\right)\right|$
$=\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\left(\sum_{p=1}^{\infty}\frac{(2p)!\tau^{p}}{4^{p}(p!)^{2}}+\mathcal{O}(\epsilon)\right)\right|$
$=\frac{\tau+\mathcal{O}(\tau^{2})}{2}\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|+\mathcal{O}(\epsilon)$
where the inequality (69) is a consequence of Assumption B. ∎
B.3 Proof of Lemma 5.6
We first derive a useful approximation for a general smooth function.
Lemma B.1.
Let $\phi:\mathbf{R}^{n}\to\mathbf{R}$ be a smooth ($C^{\infty}$) function and $X\in\mathbf{R}^{n}$ a random variable. Then we have
$\mathbf{E}[\phi(X)]=\phi(\mathbf{E}[X])+\frac{1}{2}\bigl{\langle}\nabla^{2}\phi(\mathbf{E}[X]),\textup{Cov}(X)\bigr{\rangle}+\sum_{|\bm{p}|=3}^{\infty}\!\frac{D^{(\bm{p})}\phi(\mathbf{E}[X])}{\bm{p}!}\mathbf{E}\bigl{[}(X\!-\!\mathbf{E}[X])^{(\bm{p})}\bigr{]},$ (70)
[…]
$D^{(\bm{p})}\phi$ $=\frac{\partial^{|p|}\phi}{\partial X^{p}}=\frac{\partial^{p_{1}+\cdots+p_{n}}\phi}{\partial X_{1}^{p_{1}}\cdots\partial X_{n}^{p_{n}}},$
$(X-\mathbf{E}[X])^{(\bm{p})}$ $=(X_{1}-\mathbf{E}[X_{1}])^{p_{1}}\cdots(X_{n}-\mathbf{E}[X_{n}])^{p_{n}}.$
Proof.
Let $\delta:=X-\mathbf{E}[X]$ and consider the Taylor expansion of $\phi$ at $\mathbf{E}[X]$:
$\phi(X)$ $=\phi(\mathbf{E}[X])+\nabla\phi(\mathbf{E}[X])^{T}\delta+\frac{1}{2}\delta^{T}\nabla^{2}\phi(\mathbf{E}[X])\delta+\sum_{|\bm{p}|=3}^{\infty}\frac{D^{(\bm{p})}\phi(\mathbf{E}[X])}{\bm{p}!}\delta^{(\bm{p})}.$
Taking expectation with respect to $X$, we have $\mathbf{E}[\delta]=\mathbf{E}\bigl{[}X-\mathbf{E}[X]\bigr{]}=0$. Therefore,
$\mathbf{E}\bigl{[}\nabla\phi(\mathbf{E}[X])^{T}\delta\bigr{]}=\nabla\phi(\mathbf{E}[X])^{T}\mathbf{E}\bigl{[}\delta\bigr{]}=0,$
which, together with the equation $\textup{Cov}(X)=\mathbf{E}[\delta\delta^{T}]$, yields the desired result. ∎
Proof of Lemma 5.6.
We apply Lemma B.1 with $X:=(Y,Z)$ and $\phi(x)=\phi(y,z):=\frac{y}{\sqrt{z}}$. First, the gradient and Hessian of $g$ can be calculated as
[…]
For general $p$-th partial derivative, we derive the following result for any $q\in\{0,\ldots,p\}$:
$\frac{\partial^{p}\phi}{\partial y^{q}\partial z^{p-q}}=\frac{\partial^{p-q}}{\partial z^{p-q}}\left(\frac{\partial^{q}\phi}{\partial y^{q}}\right)=\begin{cases}\displaystyle 0\quad&\text{if }q\geq 2,\\[8.61108pt] \displaystyle\frac{\partial^{p-1}}{\partial z^{p-1}}\frac{1}{\sqrt{z}}=(-1)^{p-1}\frac{(2p-2)!}{4^{p-1}(p-1)!}z^{-\frac{2p-1}{2}}\quad&\text{if }q=1,\\[8.61108pt] \displaystyle y\cdot\frac{\partial^{p}}{\partial z^{p}}\frac{1}{\sqrt{z}}=(-1)^{p}\frac{(2p)!}{4^{p}p!}yz^{-\frac{2p+1}{2}}\quad&\text{if }q=0.\\ \end{cases}$
Substitute the Hessian and $p$-th order partial derivative into (70), we get
$\mathbf{E}\left[\frac{Y}{\sqrt{Z}}\right]$ $=\frac{\mathbf{E}[Y]}{\sqrt{\mathbf{E}[Z]}}-\mathbf{E}\left[\frac{(Y-\mathbf{E}[Y])(Z-\mathbf{E}[Z])}{2\mathbf{E}[Z]^{3/2}}\right]+\mathbf{E}\left[\frac{3\mathbf{E}[Y](Z-\mathbf{E}[Z])^{2}}{8\mathbf{E}[Z]^{5/2}}\right]$
$\quad+\sum_{p=3}^{\infty}\frac{1}{p!}\left(\frac{p(2p-2)!}{4^{p-1}(p-1)!}\mathbf{E}\left[\frac{(Y-\mathbf{E}[Y])(Z-\mathbf{E}[Z])^{p-1}}{\mathbf{E}[Z]^{\frac{2p-1}{2}}}\right]+(-1)^{p}\frac{(2p)!}{4^{p}p!}\mathbf{E}\left[\frac{\mathbf{E}[Y](Z-\mathbf{E}[Z])^{p}}{\mathbf{E}[Z]^{\frac{2p+1}{2}}}\right]\right)$
$=\frac{\mathbf{E}[Y]}{\sqrt{\mathbf{E}[Z]}}-\frac{\textup{Cov}(Y,Z)}{2\mathbf{E}[Z]^{3/2}}+\frac{3\mathbf{E}[Y]\textup{Var}(Z)}{8\mathbf{E}[Z]^{5/2}}$
$\quad+\mathcal{O}\left(\frac{\mathbf{E}\bigl{[}(Y-\mathbf{E}[Y])(Z-\mathbf{E}[Z])^{p-1}\bigr{]}}{\mathbf{E}[Z]^{p-1/2}}\right)+\mathcal{O}\left(\frac{\mathbf{E}[Y]\mathbf{E}\bigl{[}(Z-\mathbf{E}[Z])^{p}\bigr{]}}{\mathbf{E}[Z]^{p+1/2}}\right)$
$=\frac{\mathbf{E}[Y]}{\sqrt{\mathbf{E}[Z]}}\Biggl{(}1-\frac{\textup{Cov}(Y,Z)}{2\mathbf{E}[Y]\mathbf{E}[Z]}+\frac{3\textup{Var}(Z)}{8\mathbf{E}[Z]^{2}}$
$\quad+\mathcal{O}\left(\frac{\mathbf{E}\bigl{[}(Y-\mathbf{E}[Y])(Z-\mathbf{E}[Z])^{p-1}\bigr{]}}{\mathbf{E}[Y]\mathbf{E}[Z]^{p-1}}\right)+\mathcal{O}\left(\frac{\mathbf{E}\bigl{[}(Z-\mathbf{E}[Z])^{p}\bigr{]}}{\mathbf{E}[Z]^{p}}\right)\Biggr{)},$
which is the desired result. ∎
B.4 Bias and variance calculation
Here we derive the bias and variance expressions for some estimators listed in Section 5.4.
  • Using EMA of $d_{t}^{2}$ as estimator. In this case, we have $v_{t}=\beta_{1}v_{t-1}+(1-\beta)d_{t}^{2}$ and
    $\textup{Var}_{t}(v_{t})$ $=\mathbf{E}_{t}\!\left[\left(\sum_{k=1}^{t}(1-\beta)\beta^{t-k}d_{k}^{2}-\sum_{k=1}^{t}(1-\beta)\beta^{t-k}\mathbf{E}_{t}[d_{k}^{2}]\right)^{2}\right]$
    $=\mathbf{E}_{t}\!\left[\left(\sum_{k=1}^{t-1}(1-\beta)\beta^{t-k}d_{k}^{2}+(1-\beta)d_{t}^{2}-\sum_{k=1}^{t-1}(1-\beta)\beta^{t-k}d_{k}^{2}-(1-\beta)\mathbf{E}_{t}[d_{k}^{2}]\right)^{2}\right]$
    $=(1-\beta)^{2}\mathbf{E}_{t}\!\left[\left(d_{t}^{2}-\mathbf{E}_{t}[d_{t}^{2}]\right)^{2}\right]$
    $=(1-\beta)^{2}\textup{Var}_{t}\left(d_{t}^{2}\right).$
  • The Adam estimator. In this case, we have $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2}$ and
    $\textup{Var}_{t}(v_{t})$ $=\mathbf{E}_{t}\left[\left(\sum_{k=1}^{t}(1-\beta_{2})\beta_{2}^{t-k}g_{k}^{2}-\sum_{k=1}^{t}(1-\beta_{2})\beta_{2}^{t-k}\mathbf{E}_{t}\left[g_{k}^{2}\right]\right)^{2}\right]$
    $=\mathbf{E}_{t}\left[\left(\sum_{k=1}^{t-1}(1-\beta_{2})\beta_{2}^{t-k}g_{k}^{2}+(1-\beta_{2})g_{t}^{2}-\sum_{k=1}^{t-1}(1-\beta_{2})\beta_{2}^{t-k}g_{k}^{2}-(1-\beta_{2})\mathbf{E}_{t}[g_{t}^{2}]\right)^{2}\right]$
    $=(1-\beta_{2})^{2}\mathbf{E}_{t}\left[\left(g_{t}^{2}-\mathbf{E}_{t}[g_{t}^{2}]\right)^{2}\right]$
    $=(1-\beta_{2})^{2}\textup{Var}_{t}\left(g_{t}^{2}\right).$
  • The conditional estimator of BCOS-c. In this case, we have $d_{t}=m_{t}=\beta m_{t-1}+(1-\beta)g_{t}$ and $v_{t}=(1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$. Therefore,
    $\mathrm{Bias}$ $=|\mathbf{E}_{t}[v_{t}]-\mathbf{E}_{t}[d_{t}^{2}]|$
    $=\left|\mathbf{E}_{t}\left[(1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}\right]-\mathbf{E}_{t}\left[\left(\beta m_{t-1}+(1-\beta)g_{t}\right)^{2}\right]\right|$
    $=\left|(2\beta-\beta^{2})m_{t-1}^{2}+(1-\beta)^{2}\mathbf{E}_{t}\left[g_{t}^{2}\right]-\beta^{2}m^{2}_{t-1}-2\beta(1-\beta)m_{t-1}\mathbf{E}_{t}\left[g_{t}\right]-(1-\beta)\mathbf{E}_{t}\left[g_{t}^{2}\right]\right|$
    $=\left|(2\beta-2\beta^{2})m_{t-1}^{2}-2\beta(1-\beta)m_{t-1}\mathbf{E}_{t}\left[g_{t}\right]\right|$
    $=2\beta(1-\beta)\left|m_{t-1}\left(m_{t-1}-\mathbf{E}_{t}\left[g_{t}\right]\right)\right|,$
    and
    $\textup{Var}_{t}(v_{t})$ $=\mathbf{E}_{t}\left[\left((1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}-(1-(1-\beta)^{2})m_{t-1}^{2}-(1-\beta)^{2}\mathbf{E}_{t}[g_{t}^{2}]\right)^{2}\right]$
    $=(1-\beta)^{4}\mathbf{E}_{t}\left[\left(g_{t}^{2}-\mathbf{E}_{t}[g_{t}^{2}]\right)^{2}\right]$
    $=(1-\beta)^{4}\textup{Var}_{t}\left(g_{t}^{2}\right).$

### v2

@@
 Abstract
-We consider stochastic approximation with block-coordinate stepsizes and propose adaptive stepsize rules that aim to minimize the expected distance from the next iterate to an optimal point.
+We consider stochastic approximation with block-coordinate stepsizes and propose adaptive stepsize rules that aim to minimize the expected distance from the next iterate to an (unknown) target point.
 These stepsize rules employ online estimates of the second moment of the search direction along each block coordinate.
-The popular Adam algorithm can be interpreted as a particular heuristic for such estimation.
-By leveraging a simple conditional estimator, we derive a new method that obtains comparable performance as Adam but requires less memory and fewer hyper-parameters.
-We prove that this family of methods converges almost surely to a small neighborhood of the optimal point, and the radius of the neighborhood depends on the bias and variance of the second-moment estimator.
+The popular Adam algorithm can be interpreted as a variant with a specific estimator.
+By leveraging a simple conditional estimator, we derive a new method that obtains competitive performance against Adam but requires less memory and fewer hyper-parameters.
+We prove that this family of methods converges almost surely to a small neighborhood of the target point, and the radius of the neighborhood depends on the bias and variance of the second-moment estimator.
 Our analysis relies on a simple aiming condition that assumes neither convexity nor smoothness, thus has broad applicability.
+Keywords: stochastic approximation, block coordinate methods, conditional estimator, aiming condition, almost sure convergence.
 1 Introduction
@@
 where $x\in\mathbf{R}^{n}$ is the decision variable, $\xi$ is a random variable, and $f$ is the loss function.
-In the context of statistical machine learning (e.g., hastie2009elements, ), $x$ represents the parameters of a prediction model, $\xi$ represents randomly sampled data from some (unknown) probability distribution, and $f(x,\xi)$ is the error in making predictions about $\xi$ using the model parametrized by $x$.
-We assume that $F$ is bounded below and there exists an optimal solution $x_{*}$.
+In the context of statistical machine learning (e.g., hastie2009elements), $x$ represents the parameters of a prediction model, $\xi$ represents randomly sampled data from some (unknown) probability distribution, and $f(x,\xi)$ is the error in making predictions about $\xi$ using the model parametrized by $x$.
+We assume that $F$ is bounded below.
 Suppose that for any pair $x$ and $\xi$, we can evaluate the gradient of $f$ with respect to $x$, denoted as $\nabla f(x,\xi)$.
-Starting with an initial point $x_{0}\in\mathbf{R}^{n}$, the classical stochastic approximation method RobbinsMonro51 generates a sequence $\{x_{1},x_{2},\ldots\}$ with the update rule
+Starting with an initial point $x_{0}\in\mathbf{R}^{n}$, the classical stochastic approximation method (RobbinsMonro51) generates a sequence $\{x_{1},x_{2},\ldots\}$ with the update rule
 $x_{t+1}=x_{t}-\alpha_{t}\nabla f(x_{t},\xi_{t}),$ (2)
 where $\alpha_{t}$ is the stepsize (often called the learning rate in the machine learning literature).
-The convergence properties of this method are well studied in the stochastic approximation literature (e.g., RobbinsMonro51, ; Wasan69book, ; KushnerYin03book, ).
+The convergence properties of this method are well studied in the stochastic approximation literature (e.g., RobbinsMonro51; Wasan69book; KushnerYin03book).
 In particular, the stepsizes $\alpha_{t}$ should be nonnegative and unsummable, i.e., $\sum_{t=0}^{\infty}\alpha_{t}=\infty$.
-Stronger guarantees such as almost sure convergence also require $\alpha_{t}$ to be square-summable, i.e., $\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty$ (e.g., RobbinsMonro51, ; Blum1954multidim, ; chung1954, ; ROBBINS1971, ).
+Stronger guarantees such as almost sure convergence also require $\alpha_{t}$ to be square-summable, i.e., $\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty$ (e.g., RobbinsMonro51; Blum1954multidim; chung1954; ROBBINS1971).
 Despite the rich literature on their convergence theory, stochastic approximation methods in practice often require heuristics and trial and error in choosing the stepsize sequence $\{\alpha_{t}\}$.
-Adaptive rules that can adjust stepsizes on the fly have been developed in both the optimization literature (e.g., Kesten58, ; MirzoakhmedovUryasev83, ; RuszczynskiSyski83stats, ; RuszczynskiSyski84IFAC, ; DelyonJuditsky93, ) and by the machine learning community (e.g., Jacobs88DBD, ; Sutton92IDBD, ; Schraudolph1999, ; MahmoodSutton12TuningFree, ).
-More recently, adaptive algorithms that use coordinate-wise stepsizes have become very popular following the seminal works of AdaGrad DuchiHazanSinger2011adagrad and Adam KingmaBa2014adam , and they demonstrate significantly better performance in training large-scale deep neural networks.
+Adaptive rules that can adjust stepsizes on the fly have been developed in both the optimization literature (e.g., Kesten58; MirzoakhmedovUryasev83; RuszczynskiSyski83stats; RuszczynskiSyski84IFAC; DelyonJuditsky93) and by the machine learning community (e.g., Jacobs88DBD; Sutton92IDBD; Schraudolph1999; MahmoodSutton12TuningFree).
+More recently, adaptive algorithms that use coordinate-wise stepsizes have become very popular following the seminal works of AdaGrad (DuchiHazanSinger2011adagrad) and Adam (KingmaBa2014adam), and they demonstrate significantly better performance in training large-scale deep neural networks.
 In this paper, we propose a family of stochastic approximation methods with block-coordinate stepsizes and present a general framework for their convergence analysis.
@@
 where $\beta_{1}\in[0,1)$ is a smoothing factor. This is often called stochastic momentum.
-The Adam algorithm KingmaBa2014adam uses the stochastic momentum in (4) as its search direction and sets the coordinate-wise stepsizes as
+The Adam algorithm (KingmaBa2014adam) uses the stochastic momentum in (4) as its search direction and sets the coordinate-wise stepsizes as
 $s_{t,i}=\frac{\alpha_{t}}{\sqrt{v_{t,i}}+\epsilon},\qquad i=1,\ldots,n,$ (5)
@@
 For Adam to perform well, the second smoothing factor $\beta_{2}$ is often set to be larger than $\beta_{1}$. For example, a pair of commonly used choices is $\beta_{1}=0.9$ and $\beta_{2}=0.99$.
-Adam KingmaBa2014adam and its variant AdamW Loshchilov2019AdamW (which implements decoupled weight decay) have been very successful in training large-scale deep neural networks.
-However, theoretical understanding of their superior performance is still incomplete despite many recent efforts (e.g., reddi2019convergence, ; bock2018improvement, ; balles2018dissecting, ; zou2019sufficient, ; defossez2020simple, ; Luo2022neurips, ; kunstner2024heavy, ).
-On the other hand, there are also many works that propose new variants or alternatives to Adam and AdamW, either starting from fundamental optimization principles (e.g., yao2021adahessian, ; gomes2024adafisher, ; hwang2024fadam, ; lin2024can, ; jordan2024muon, ; pethick2025LMOs, ) or based on empirical algorithmic search (e.g., chen2023symbolic, ; zhang2024adam, ).
+Adam (KingmaBa2014adam) and its variant AdamW (Loshchilov2019AdamW), which implements decoupled weight decay, have been very successful in training large-scale deep neural networks.
+However, theoretical understanding of their superior performance is still incomplete despite many recent efforts (e.g., bock2018improvement; balles2018dissecting; reddi2019convergence; zou2019sufficient; defossez2020simple; Luo2022neurips; kunstner2024heavy).
+On the other hand, there are also many works that propose new variants or alternatives to Adam and AdamW, either starting from fundamental optimization principles (e.g., yao2021adahessian; gomes2024adafisher; hwang2024fadam; jordan2024muon; lin2024can; pethick2025LMOs) or based on empirical algorithmic search (e.g., chen2023symbolic; zhang2024adam).
 But all have limited success.
@@
 More specifically:
-  • In Section 2, we derive BCOS by minimizing the expected distance from the next iterate to an optimal point.
+  • In Section 2, we derive BCOS by minimizing the expected distance from the next iterate to a target point.
 While the optimal stepsizes cannot be computed exactly, we make several simplifications and approximate the second moments of the search direction with simple EMA estimators.
   • In Section 3, we instantiate BCOS with specific search directions.
-In particular, we show that RMSProp rmsprop2012lecture and Adam KingmaBa2014adam can be interpreted as special cases of BCOS.
-With the stochastic momentum as search direction, we leverage a simple conditional estimator to derive a variant that requires fewer optimizer states and hyper-parameters than Adam.
-Integrating with decoupled weight decay Loshchilov2019AdamW gives the BCOSW variants that enjoy comparable performance as AdamW.
+In particular, we show that RMSProp (rmsprop2012lecture) and Adam (KingmaBa2014adam) can be interpreted as special cases of BCOS.
+With the stochastic momentum as search direction, we leverage a simple conditional estimator to derive a variant that requires less memory and fewer hyper-parameters than Adam.
+Integrating with decoupled weight decay (Loshchilov2019AdamW) gives the BCOSW variants that enjoy competitive performance as AdamW.
   • In Section 4, we present numerical experiments to compare several variants of BCOS(W) with Adam(W) on several deep learning tasks.
@@
 Generality
-The general form of stochastic approximation (e.g., RobbinsMonro51, ; Wasan69book, ; KushnerYin03book, ) concerns solving a nonlinear system of equations $G(x)=0$ where the map $G:\mathbf{R}^{n}\to\mathbf{R}^{n}$ only admits noisy evaluations in the form of $g(\cdot,\xi)$, where $\xi$ is a random variable, under the assumption that $G(x)=\mathbf{E}_{\xi}[g(x,\xi)]$.
+The general form of stochastic approximation (e.g., RobbinsMonro51; Wasan69book; KushnerYin03book) concerns solving a nonlinear system of equations $G(x)=0$ where the map $G:\mathbf{R}^{n}\to\mathbf{R}^{n}$ only admits noisy evaluations in the form of $g(\cdot,\xi)$, where $\xi$ is a random variable, under the assumption that $G(x)=\mathbf{E}_{\xi}[g(x,\xi)]$.
 In this case, stochastic approximation takes the form
@@
 Notice that $\gamma_{t,k}$ is always a scalar and $\gamma_{t}$ is a vector in $\mathbf{R}^{m}$ instead of $\mathbf{R}^{n}$ (unless $m=n$).
-[…]
-The $\textup{sign}(\cdot)$ function is defined as $\textup{sign}(\alpha)=1$ if $\alpha>0$, $-1$ if $\alpha<0$ and 00 if $\alpha=0$.
+Throughout this paper, $\langle\cdot,\cdot\rangle$ denotes the standard inner product in the Euclidean space $\mathbf{R}^{n}$ and $\|\cdot\|$ denotes the induced Euclidean norm.
+The $\textup{sign}(\cdot)$ function is defined as $\textup{sign}(\alpha)=1$ if $\alpha>0$, 1-1 if $\alpha<0$ and 0 if $\alpha=0$; for vectors, it applies element-wise.
 2 Derivation of BCOS
@@
 2.1 Block-coordinate optimal stepsizes
-We consider the distance of the next iterate $x_{t+1}$ to an optimal point $x_{*}$. Specifically, in terms of the squared Euclidean distance,
+We consider the distance of the next iterate $x_{t+1}$ to an (unknown) target point $x_{*}$. Specifically, in terms of the squared Euclidean distance,
 $\|x_{t+1}-x_{*}\|^{2}$ $=\|x_{t}-s_{t}\odot d_{t}-x_{*}\|^{2}$
@@
 2.2 Simplification and approximation
-We first try to avoid the direct reliance on $x_{*}$. To this end, rewrite the numerator in (10) as
+We first try to avoid the direct reliance on $x_{*}$, which is unknown to us. To this end, rewrite the numerator in (10) as
 $\bigl\langle x_{t,k}-x_{*,k},\,\mathbf{E}_{t}[d_{t,k}]\bigr\rangle=\|x_{t,k}-x_{*,k}\|\|\mathbf{E}_{t}[d_{t,k}]\|\cos\theta_{t,k},$
@@
 In particular, while the optimal stepsizes $\widehat{\gamma}_{t,k}$ can be positive or negative, it is very hard to even correctly estimate the sign of the inner product $\langle x_{t,k}-x_{*,k},\,\mathbf{E}_{t}[d_{t,k}]\rangle$.
-We take the pragmatic approach of restricting $\alpha_{t,k}>0$, effectively being optimistic that the expected direction $-\mathbf{E}_{t}[d_{t,k}]$ is always pointing towards $x_{*,k}$ (i.e., assuming $\cos\theta_{t,k}>0$).
+We take the pragmatic approach of restricting $\alpha_{t,k}>0$, effectively being optimistic that the expected direction $\mathbf{E}_{t}[d_{t,k}]$ is always pointing towards $x_{*,k}$ (i.e., assuming $\cos\theta_{t,k}>0$).
 A further simplification is to use a common stepsize schedule $\alpha_{t}$ across all blocks, i.e., $\alpha_{t,k}=\alpha_{t}$ for all $k=1,\ldots,m$.
-This can be a reasonable choice for training deep neural networks, where the model parameters are initialized randomly coordinate-wise such that $\mathbf{E}[\|x_{0,k}\|]$ is constant for each coordinate $k$ HeZhangRenSun2015 ; RobertsYaidaHanin2022 ; DinanYaidaZhang2023 .
+This can be a reasonable choice for training deep neural networks, where the model parameters are initialized randomly coordinate-wise such that $\mathbf{E}[\|x_{0,k}\|]$ is constant for each coordinate $k$ (HeZhangRenSun2015; RobertsYaidaHanin2022; DinanYaidaZhang2023).
 This brings us to
@@
 Therefore, we expect $\alpha_{t}$ to decrease as $\|x_{t}-x_{*}\|$ gradually shrinks.
-A simple strategy is to use a monotonic stepsize schedule on $\alpha_{t}$, such as the popular cosine decay (loshchilov2016sgdr, ) or linear decay (defazio2024optimallineardecaylearning, ).
+A simple strategy is to use a monotonic stepsize schedule on $\alpha_{t}$, such as the popular cosine decay (loshchilov2016sgdr) or linear decay (defazio2024optimallineardecaylearning).
 Next, we need to replace the conditional expectations $\mathbf{E}_{t}[d_{t,k}]$ and $\mathbf{E}_{t}[\|d_{t,k}\|^{2}]$ in (13) with computable approximations.
@@
 $\rho_{t,k}=\frac{\|\mathbf{E}_{t}[d_{t,k}]\|^{2}}{\mathbf{E}_{t}[\|d_{t,k}\|^{2}]}=\frac{\|\mathbf{E}_{t}[d_{t,k}]\|^{2}}{\|\mathbf{E}_{t}[d_{t,k}]\|^{2}+\textup{Var}_{t}(d_{t,k})}.$ (16)
-Apparently we have $\rho_{t,k}\in[0,1]$, and the two boundary values of 00 and $1$ correspond to the cases of zero signal and full signal (zero noise) respectively.
+Apparently we have $\rho_{t,k}\in[0,1]$, and the two boundary values of 0 and $1$ correspond to the cases of zero signal and full signal (zero noise) respectively.
 With the definition of SiF, we can decompose the stepsizes in (12) as
@@
 are invariant of scaling of the search directions $d_{t,k}$, which makes the tuning of $\alpha^{\prime}_{t}$ much easier in practice.
-This invariance property is also shared by Adam KingmaBa2014adam , as seen from (5) and (6).
+This invariance property is also shared by Adam (KingmaBa2014adam), as seen from (5) and (6).
 The similarity of Adam and the simplified BCOS in (17) is apparent, and we will explain their connection in detail in the next section.
@@
 We call it BCOS-g to signify the use of gradient as search direction.
-We immediately recognize that BCOS-g is exactly the RMSprop algorithm rmsprop2012lecture , which is one of the first effective algorithms for training deep neural networks.
+We immediately recognize that BCOS-g is exactly the RMSprop algorithm (rmsprop2012lecture), which is one of the first effective algorithms for training deep neural networks.
 Our BCOS framework gives a novel interpretation of RMSprop and its effectiveness.
@@
 $x_{t+1}=x_{t}-\alpha_{t}\,\textup{sign}(g_{t}),$ (21)
-which also received significant attention in the literature Polyak1973pseudo ; bernstein2018signsgd ; safaryan2021signdescent ; jiang2024efficient .
+which also received significant attention in the literature (Polyak1973pseudo; bernstein2018signsgd; safaryan2021signdescent; jiang2024efficient).
 BCOS-m
-Using the stochastic momentum as search direction has a long history in stochastic approximation (e.g., GupalBazhenov72, ; Polyak77comparison, ; RuszczynskiSyski83stats, ).
+Using the stochastic momentum as search direction has a long history in stochastic approximation (e.g., GupalBazhenov72; Polyak77comparison; RuszczynskiSyski83stats).
 It has become the default option for modern deep learning due to its superior performance compared with using plain stochastic gradients.
@@
 This avoids the initial bias caused by setting $v_{-1}=0$.
-An alternative way for bias correction is to initialize $m_{-1}$ and $v_{-1}$ as the zero vector, and then multiply $m_{t}$ and $v_{t}$ by $1/(1-\beta^{t+1})$, which quickly converges to $1$, as done in Adam KingmaBa2014adam .
+An alternative way for bias correction is to initialize $m_{-1}$ and $v_{-1}$ as the zero vector, and then multiply $m_{t}$ and $v_{t}$ by $1/(1-\beta^{t+1})$, which quickly converges to $1$, as done in Adam (KingmaBa2014adam).
 Our numerical experiments show that these two choices do not have significant differences in practice.
 3.2 BCOS with conditional estimator
-Recall that the optimal stepsizes $\widehat{\gamma}_{t,k}$ in (10) and their simplifications $\widetilde{\gamma}_{t,k}$ in (13) are all based on conditional expectation.
-In Section 3.1, we used coordinate-wise EMA of $d_{t}^{2}$ to approximate the conditional expectation $\mathbf{E}_{t}[d_{t}^{2}]$, i.e., $v_{t}$ as estimator of $\mathbf{E}_{t}[d_{t}^{2}]$ in BCOS-g and of $\mathbf{E}_{t}[m_{t}^{2}]$ in BCOS-m, respectively.
-In this section, we show that with $m_{t}$ as the search direction, we can exploit its update form to derive effective conditional estimators that can avoid using EMA.
-Algorithm 3 BCOS-c
-input: $x_{0}$, $\{\alpha_{t}\}$, $\beta\in[0,1)$, $\epsilon>0$
-$m_{-1}=\nabla f(x_{0},\xi_{0})$
-for $t=0,1,2,\ldots$ do
-     $g_{t}=\nabla f(x_{t},\xi_{t})$
-     $m_{t}=\beta m_{t-1}+(1-\beta)g_{t}$
-     $v_{t}=\bigl{(}1-(1-\beta)^{2}\bigr{)}m^{2}_{t-1}+(1-\beta)^{2}g^{2}_{t}$
-     $x_{t+1}=x_{t}-\alpha_{t}\frac{m_{t}}{\sqrt{v_{t}}+\epsilon}$
-Algorithm 4 BCOSW-c
-input: $x_{0}$, $\{\alpha_{t}\}$, $\beta\in[0,1)$, $\lambda\geq 0$, $\epsilon>0$
-$m_{-1}=\nabla f(x_{0},\xi_{0})$
-for $t=0,1,2,\ldots$ do
-     $g_{t}=\nabla f(x_{t},\xi_{t})$
-     $m_{t}=\beta m_{t-1}+(1-\beta)g_{t}$
-     $v_{t}=\bigl{(}1-(1-\beta)^{2}\bigr{)}m^{2}_{t-1}+(1-\beta)^{2}g^{2}_{t}$
-     $x_{t+1}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\alpha_{t}\lambda)}x_{t}-\alpha_{t}\frac{m_{t}}{\sqrt{v_{t}}+\epsilon}$
+Recall that the derivations of optimal stepsizes $\widehat{\gamma}_{t,k}$ and $\widetilde{\gamma}_{t,k}$ in Section 2 are all based on conditional expectation $\mathbf{E}_{t}[\cdot]$.
+In Section 3.1, we used coordinate-wise EMA to approximate the conditional expectations, specifically, $v_{t}$ as an estimator of $\mathbf{E}_{t}[d_{t}^{2}]$ in BCOS-g and an estimator of $\mathbf{E}_{t}[m_{t}^{2}]$ in BCOS-m.
+In this section, we show that with $m_{t}$ as the search direction, we can exploit its update form to derive more effective conditional estimators.
 We first repeat the definition of momentum with a smoothing factor $\beta\in[0,1)$, i.e.,
 $m_{t}=\beta m_{t-1}+(1-\beta)g_{t}.$ (23)
-To derive an estimator of $\mathbf{E}_{t}[m_{t}^{2}]$, we expand the square and take the expectation of each term:
+To derive an estimator of $\mathbf{E}_{t}[m_{t}^{2}]$, we expand the square and take the conditional expectation of each term:
 $\mathbf{E}_{t}\bigl[m_{t}^{2}\bigr]$ $=\mathbf{E}_{t}\bigl[(\beta m_{t-1}+(1-\beta)g_{t})^{2}\bigr]$
 $=\beta^{2}\mathbf{E}_{t}\bigl[m_{t-1}^{2}\bigr]+2\beta(1-\beta)\mathbf{E}_{t}[m_{t-1}\odot g_{t}]+(1-\beta)^{2}\mathbf{E}_{t}\bigl[g_{t}^{2}\bigr]$
-$=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}\odot\mathbf{E}_{t}[g_{t}]+(1-\beta)^{2}\mathbf{E}_{t}\bigl{[}g_{t}^{2}\bigr{]}.$ (24)
+$=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}\odot\mathbf{E}_{t}[g_{t}]+(1-\beta)^{2}\mathbf{E}_{t}\bigl[g_{t}^{2}\bigr],$ (24)
 where we used the fact $\mathbf{E}_{t}[m_{t-1}^{2}]=m_{t-1}^{2}$ and $\mathbf{E}_{t}[m_{t-1}]=m_{t-1}$ thanks to the definition of $\mathbf{E}_{t}[\cdot]$ in (8).
@@
 $v_{t}=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}\odot m_{t}+(1-\beta)^{2}g_{t}^{2}.$ (25)
-Our numerical experiments confirmed that this indeed is a good estimator which obtains performance comparable to Adam.
-While effective in practice, the estimator in (25) looks complicated and does not carry much intuition.
-We are motivated to derive another one that is much simpler and as effective.
-The key is to approximate $\mathbf{E}[g_{t}]$ in (24) with $m_{t-1}$ instead of $m_{t}$, which results in
+The estimator in (25) does not need to store $v_{t-1}$, thus requiring less memory (fewer optimizer states). Moreover, this estimator eliminates $\beta_{2}$ as a second hyper-parameter.
+Algorithm 3 BCOS-c
+input: $x_{0}$, $\{\alpha_{t}\}$, $\beta\in[0,1)$, $\epsilon>0$,
+$m_{-1}=\nabla f(x_{0},\xi_{0})$
+for $t=0,1,2,\ldots$ do
+  $g_{t}=\nabla f(x_{t},\xi_{t})$
+  $m_{t}=\beta m_{t-1}+(1-\beta)g_{t}$
+  $v_{t}=\beta^{2}m^{2}_{t-1}+2\beta(1-\beta)m_{t-1}\odot m_{t}+(1-\beta)^{2}g^{2}_{t}$
+  ($\Big$(Simple alternative: $v_{t}=\bigl(1-(1-\beta)^{2}\bigr)m^{2}_{t-1}+(1-\beta)^{2}g^{2}_{t}$ $\Big)$
+  $x_{t+1}=x_{t}-\alpha_{t}\frac{m_{t}}{\sqrt{v_{t}+\epsilon}}$
+Replacing $v_{t}$ in BCOS-m with (25) leads to Algorithm 3. We call it BCOS-c to signify the conditional estimator.
+The estimator in (25) can be further simplified if we approximate $\mathbf{E}_{t}[g_{t}]$ in (24) with $m_{t-1}$ instead of $m_{t}$, which results in
 $v_{t}$ $=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$
 $=\left(1-(1-\beta)^{2}\right)m^{2}_{t-1}+(1-\beta)^{2}g^{2}_{t}.$ (26)
-This estimator is much simpler than (25). Moreover, it resembles the standard EMA estimator in Adam, shown in (6), with an effective smoothing factor
-$\beta^{\prime}=1-(1-\beta)^{2},$
+This estimator resembles the standard EMA estimator in Adam, shown in (6), with an effective smoothing factor
+$\beta^{\prime}=1-(1-\beta)^{2},$ (27)
 but with $v_{t-1}$ replaced by $m^{2}_{t-1}$.
-As a result, the estimator in (26) does not need to store $v_{t-1}$, thus requiring fewer optimizer states (less memory).
 This also explains that the second smoothing factor in Adam, $\beta_{2}$, corresponding to $\beta^{\prime}$ here, should be much larger or closer to 1 than $\beta$.
 Specifically, $\beta=0.9$ roughly corresponds to $\beta^{\prime}=0.99$.
-The estimator in (26) eliminates $\beta_{2}$ as a second hyper-parameter.
-Finally, replacing $v_{t}$ in BCOS-m with the one in (26) produces Algorithm 3.
-We call it BCOS-c to signify the conditional estimator.
-It has fewer optimizer states and fewer hyper-parameters to tune than BCOS-m and Adam.
-We note the slight abuse of notation in using the symbol $v_{t}$ for both the EMA estimator in BCOS-g and -m and the conditional estimator here.
-In general, $v_{t}$ denotes an online estimator of $\mathbf{E}_{t}[d_{t}^{2}]$.
-The corresponding choice of $d_{t}$ and type of estimator should be clear from the context.
-Figure 1: Comparing AdamW and BCOSW-c with different momentum parameters.
+We include the estimator (26) as a simple alternative of (25) in Algorithm 3.
+Empirically we found that it behaves almost identically as (25) on the experiments we conduct in Section 4.
+Concurrently, zhang2025adams propose a similar algorithm with
+$v_{t}=\beta^{\prime}m_{t-1}^{2}+(1-\beta^{\prime})g_{t}^{2},$
+and show that tuning $\beta^{\prime}$ as a hyper-parameter can obtain slightly better performance than using the value in (27).
+Algorithm 4 BCOSW{-g, -m, -c}
+input: $x_{0}$, $\{\alpha_{t}\}$, $\beta\in[0,1)$, $\epsilon>0$, $\lambda\geq 0$
+for $t=0,1,2,\ldots$ do
+  compute $d_{t}$ ($g_{t}$ or $m_{t}$) and $v_{t}$ as in Algorithm 1, 2 or 3 (-g, -m or -c, respectively)
+  $x_{t+1}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1-\alpha_{t}\lambda)}x_{t}-\alpha_{t}\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}$
 3.3 BCOS with decoupled weight decay
@@
 We can apply the BCOS family of algorithms to incorporate the regularization term by simply replacing $g_{t}=\nabla f(x_{t},\xi_{t})$ with $g_{t}=\nabla f(x_{t},\xi_{t})+\lambda x_{t}$.
-A more effective way in practice is to use decoupled weight decay as proposed in the AdamW algorithm Loshchilov2019AdamW . Specifically, we apply weight decay separately in the BCOS update:
+A more effective way in practice is to use decoupled weight decay as proposed in the AdamW algorithm (Loshchilov2019AdamW). Specifically, we apply weight decay separately in the BCOS update:
 $x_{t+1}=x_{t}-\gamma_{t}\odot d_{t}-\alpha_{t}\lambda x_{t}=(1-\alpha_{t}\lambda)x_{t}-\gamma_{t}\odot d_{t}.$
 We call the resulting method BCOSW following the naming convention of AdamW.
-Algorithm 4 shows BCOSW-c, which uses the conditional estimator.
-Other variants, namely BCOSW-g and BCOSW-m, can be obtained similarly.
-A PyTorch implementation of all variants of BCOS and BCOSW is given in Appendix A.
-The code is also available at https://github.com/facebookresearch/bcos.
+Algorithm 4 shows the BCOSW framework, which can be instantiated with specific $d_{t}$ and $v_{t}$ as in BCOS-g, -m or -c. A PyTorch implementation of all variants is available at
+https://github.com/facebookresearch/bcos.
 4 Numerical experiments
-Figure 2: Comparing AdamW and three variants of BCOSW. Left: the first 10k iterations; Right: all 100k iterations.
-In this section, we present preliminary experiments to compare BCOS with Adam, specifically their variants with decoupled weight decay. Among the BCOSW family, we focus on BCOSW-c (Algorithm 4).
-Our first set of experiments is conducted on training a small GPT2 model with 124 million parameters radford2019language on the OpenWebText dataset Gokaslan2019OpenWeb .
+In this section, we present preliminary experiments to compare BCOS with Adam, specifically their variants with decoupled weight decay: AdamW and BCOSW. Among the BCOSW family, we focus on BCOSW-c.
+Our first set of experiments is conducted on training a small GPT2 model with 124 million parameters (radford2019language) on the OpenWebText dataset (Gokaslan2019OpenWeb).
 We use global batch size 512 and run all experiments for 100k iterations.
 We use linear warmup on the stepsize schedule $\{\alpha_{t}\}$ to reach $\alpha_{\text{max}}$ in the first 2k iterations, and then use cosine decay to toward the final stepsize $\alpha_{\text{min}}=0.01\alpha_{\text{max}}$.
-The default hyper-parameters are chosen (based on a coarse sweep) as: peak stepsize $\alpha_{\text{max}}=0.002$, $\epsilon=10^{-6}$ and weight decay $\lambda=0.1$.
+The default hyper-parameters are chosen (based on a coarse sweep) as: peak stepsize $\alpha_{\text{max}}=0.002$, $\epsilon=10^{-12}$ and weight decay $\lambda=0.1$.
+Figure 1: Comparing AdamW and BCOSW-c with different momentum parameters.
 Figure 1 (left) shows the test loss of AdamW on the GPT2 task with different combinations of $\beta_{1}$ and $\beta_{2}$.
@@
 In contrast, Figure 1 (right) shows that BCOSW-c obtains the same final loss but with very smooth loss curves.
+Figure 2: Comparing AdamW and three variants of BCOSW. Left: the first 10k iterations; Right: all 100k iterations.
 Figure 3: Left: Adam(W) with $\beta_{1}=0.9$ and $\beta_{2}=0.99$. Right: BCOS(W)-c with $\beta=0.9$.
@@
 Figure 3 illustrates the difference between algorithms with and without decoupled weight decay.
+For Adam and BCOS-c, the weight decay is accounted for in the gradient with $g_{t}=\nabla f(x_{t},\xi_{t})+\lambda x_{t}$.
 BCOS-c converges to much higher loss than BCOSW-c, and different values of $\lambda$ (weight decay) make a dramatic difference for BCOS-c but cause little change to BCOSW-c.
@@
 BCOSW has slightly better performance for smaller values of $\alpha_{\text{max}}$, but becomes more sensitive for larger values of $\alpha_{\text{max}}$.
-[…]
+Finally, in Figure 5, we compare different algorithms for training ResNet-20 (he2016resnet) on the CIFAR10 dataset (krizhevsky2009learning), and also training the Vision Transformer (ViT) (touvron21vit) on the ImageNet dataset (imagenet2009).
+For the ResNet task, we tried both cosine decay (drop by factor 100) and step decay (drop by 10 at epochs 80, 120, 150).
+The hyper-parameters chosen are: $\beta=0.9$ for SGD and BCOSW-c, and $\beta_{1,2}=(0.9,\,0.99)$ for AdamW.
+We observe that the best-performing stepsize schedules are quite different for different methods.
+This prompts the need to tune hyper-parameters for BCOSW for different tasks even though it shares similar tuned hyper-parameters as AdamW on the GPT2 task.
 For training ViT on ImageNet, although the best tuned stepsize schedules are similar between AdamW and BCOSW, their training and test curves look quite different.
 Figure 5 (right) shows that the test precision curves for BCOSW-c rise slowly but reach slightly higher precision at the end.
-These preliminary experiments demonstrate that BCOSW-c can obtain competitive performance compared with the state-of-the-art method AdamW, but with fewer optimizer states and fewer hyper-parameters to tune.
-More comprehensive empirical studies are needed to fully understand its potential in training large deep learning models.
+These preliminary experiments demonstrate that BCOSW-c can obtain competitive performance compared with the state-of-the-art method AdamW, but require less memory and fewer hyper-parameters to tune.
+More comprehensive empirical studies are needed to fully understand its potential in training larger deep learning models.
 5 Convergence analysis
-In this section, we present convergence analysis for BCOS with a general search direction and decoupled weight decay (see Section 3.3).
-To simplify presentation, we stick with the setting of single-coordinate blocks.
-Specifically, we analyze the algorithm
-$x_{t+1}=(1-\alpha_{t}\lambda)x_{t}-\gamma_{t}\odot d_{t},\quad\text{where}\quad\gamma_{t}=\frac{\alpha_{t}}{\sqrt{v_{t}+\epsilon}}.$ (28)
-Here, stepsize $\gamma_{t}$ takes a slightly different form from the one used throughout Section 3, where $\epsilon$ is outside the square-root in the denominator.
-The resulting difference between the two forms is negligible for corresponding choices of $\epsilon$, with the one in (28) being squared (e.g., being $10^{-12}$ if the one in Section 3 is $10^{-6}$).
-We choose the form in (28) for ease of presenting the analysis.
-Recall that $v_{t}\in\mathbf{R}_{+}^{n}$ is an online estimator of $\mathbf{E}_{t}[d_{t}^{2}]$. Our analysis consists of two stages:
+In this section, we present convergence analysis for BCOS(W). To simplify presentation, we stick with the setting of single-coordinate blocks. Specifically, we analyze the algorithm
+$x_{t+1}=(1-\alpha_{t}\lambda)x_{t}-\gamma_{t}\odot d_{t},\quad\text{where}\quad\gamma_{t}=\frac{\alpha_{t}}{\sqrt{v_{t}+\epsilon}},$ (29)
+where $d_{t}$ is a general search direction and $v_{t}\in\mathbf{R}_{+}^{n}$ is an online estimator of $\mathbf{E}_{t}[d_{t}^{2}]$. Our analysis consists of two stages:
   • First, we analyze the convergence behavior of the conceptual BCOS method
@@
 But its analysis is essential for understanding and analyzing the practical method.
-Our analysis relies on an aiming condition, which originates from the classical stochastic approximation literature RobbinsMonro51 ; Blum1954multidim ; chung1954 ; Sakrison1966 .
+Our analysis relies on an aiming condition, which originates from the classical stochastic approximation literature (RobbinsMonro51; Blum1954multidim; chung1954; Sakrison1966).
 Under this assumption, we establish almost sure convergence of $x_{t}$ to a target point $x_{*}$ and the $\mathcal{O}(1/t)$ convergence rate of $\mathbf{E}[\|x_{t}-x_{*}\|^{2}]$.
@@
 The rest of this section is organized as follows.
-In Section 5.1, we introduce the aiming condition and discuss its connection with convexity.
+In Section 5.1, we introduce the aiming condition and give some empirical evidence.
 We present the convergence analysis for the conceptual BCOS method in Section 5.2 and then the analysis for the practical BCOS method in Section 5.3.
-Our analysis is very general and applies to the whole BCOS family of methods, including RMSprop, Adam, AdamW, and variants of spectral BCOS.
+Our analysis is very general and applies to the whole BCOS family of methods, including the sign-gradient method, RMSprop and Adam(W).
 To obtain specific results for each method, we only need to estimate the bias and variance of the second-moment estimator $v_{t}$, which are presented in Section 5.4.
 5.1 Aiming condition
-First consider the stochastic approximation method (7). A typical set of assumptions in the classical stochastic approximation literature (specifically Sakrison1966 ) includes:
+First consider the stochastic approximation method (7). A typical set of assumptions in the classical stochastic approximation literature includes:
   • The random variables $\{\xi_{t}\}_{t\geq 0}$ are independent and $g(x,\xi)$ has bounded variance;
@@
     $L\|x-x_{*}\|^{2}~\geq~\bigl\langle x-x_{*},\,\mathbf{E}_{\xi}[g(x,\xi)]\bigr\rangle~\geq~\mu\|x-x_{*}\|^{2}.$ (32)
-Under these conditions, the sequence $\{x_{t}\}$ converges to $x_{*}$ in the mean-square sense, i.e., $\lim_{t\to\infty}\mathbf{E}[\|x_{t}-x_{*}\|^{2}]=0$ Sakrison1966 .
-In particular, the last condition (31) ensures that the expected direction $\mathbf{E}_{\xi}[g(x,\xi)]$ always points towards the target $x_{*}$, implying that the distance from $x_{*}$ can be reduced along this direction (with appropriate stepsize).
+Under these conditions, the sequence $\{x_{t}\}$ converges to $x_{*}$ in the mean-square sense, i.e., $\lim_{t\to\infty}\mathbf{E}[\|x_{t}-x_{*}\|^{2}]=0$ (e.g., Sakrison1966).
+In particular, the last condition (32) ensures that the direction $\mathbf{E}_{\xi}[g(x,\xi)]$ always points towards the target $x_{*}$, implying that the distance from $x_{*}$ can be reduced along this direction (with appropriate stepsize).
 For this reason, we call it an aiming condition.
-Similar results can be established under weaker aiming conditions than (31); for example, without the upper bound on the inner product.
-But they are more subtle and require more sophisticated analysis (e.g., RobbinsMonro51, ; Wolfowitz1952, ; Blum1954prob-one, ; Blum1954multidim, ; Dvoretzky1956, ; DermanSacks1959, ; Wasan69book, ).
-Following the same idea, we rewrite the conceptual BCOS method (29) as
+In the context of stochastic optimization, i.e., minimizing $F(x)=\mathbf{E}_{\xi}[f(x,\xi)]$ with $g(x,\xi)=\nabla f(x,\xi)$ and $\mathbf{E}_{\xi}[g(x,\xi)]=\nabla F(x)$, the aiming condition (32) is implied by $F$ being $\mu$-strongly convex and $L$-smooth (e.g., Nesterov04book).
+But in general it is a much weaker condition.
+Moreover, the upper bound on the inner product can often be removed by leveraging more sophisticated analysis (e.g., RobbinsMonro51; Wolfowitz1952) or careful choice of the search direction $g(x,\xi)$ (e.g.
+Blum1954multidim; Dvoretzky1956).
+This is the case of our aiming condition, which we present next.
+In order to allow reduction of the expected distance from $x_{t+1}$ to $x_{*}$, we rewrite the conceptual BCOS method (30) as
 $x_{t+1}=x_{t}-\alpha_{t}\lambda x_{t}-\widetilde{\gamma}_{t}\odot d_{t}=x_{t}-\alpha_{t}\left(\frac{d_{t}}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{t}\right)$
@@
 $\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{t}\right]=\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{t}.$
-Assumption A (Aiming condition).
+Assumption 1 (Aiming condition)
 There exists $x_{*}\in\mathbf{R}^{n}$ such that
@@
 holds for all $t\geq 0$ almost surely.
-We make the following remarks regarding Assumption A.
+We make the following remarks regarding Assumption 1.
   • The condition (33) in general depends on the past trajectory $\{x_{0},d_{0},x_{1},d_{1},\ldots,x_{t}\}$.
@@
   • We do not need an upper bound on the inner product as in (32) due to the normalization effect of dividing $d_{t}$ by $\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}$ in BCOS.
-A similar one-sided aiming condition was used in Blum1954multidim to establish almost sure convergence of the first multi-dimensional extension of the Robbins-Monro method.
+A similar one-sided aiming condition was used in (Blum1954multidim) to establish almost sure convergence of the first multi-dimensional extension of the Robbins-Monro method.
 Let $\rho_{t}\in[0,1]^{n}$ be the vector of coordinate-wise SiF defined in (16). Then we have
@@
 $\bigl\langle x_{t}-x_{*},\,\sqrt{\rho_{t}}\odot\textup{sign}(\mathbf{E}_{t}[d_{t}])+\lambda x_{t}\bigr\rangle\geq\lambda\|x_{t}-x_{*}\|^{2}.$ (34)
-Intuitively, the inner product in the aiming condition is predominantly determined by the coordinates that are far from optimal, in other words, the coordinates with relatively large $|x_{t,k}-x_{*,k}|$ and having a large SiF $\rho_{t,k}$ (effectively high signal-to-noise ratio).
-5.1.1 Connection with convexity
-In the context of stochastic optimization, i.e., minimizing $F(x)=\mathbf{E}_{\xi}[f(x,\xi)]$ with $g(x,\xi)=\nabla f(x,\xi)$, the classical aiming condition (31) is implied by $F$ being $\mu$-strongly convex and $L$-smooth (e.g., Nesterov04book, ).
-But in general it is a much weaker condition.
-Our aiming condition in Assumption A does not relate directly to smoothness and is not always implied by (strong) convexity either.
-However, it does have overlapping characteristics with convexity.
-For ease of comparison, we consider the general block structure described in Section 1.3 with $d_{t}=\nabla f(x_{t},\xi_{t})$ and set $\lambda=0$.
-In this case, we have $\mathbf{E}[d_{t}]=\nabla F(x_{t})$ and
-$\sum_{k=1}^{m}\left\langle x_{k}-x_{\ast,k},~{}\frac{\nabla_{k}F(x)}{\sqrt{\mathbf{E}\bigl{[}\left\|\nabla_{k}f(x,\xi)\right\|^{2}\bigr{]}}}\right\rangle\geq 0,\qquad\forall\,x,$ (34)
-where $\nabla_{k}F(x)\in\mathbf{R}^{n_{k}}$ denotes the sub-vector of $\nabla F(x)$ corresponding to the $k$th block.
-In the specific case of a full-dimensional block ($m=1$), the denominator $\sqrt{\mathbf{E}\bigl{[}\|\nabla f(x,\xi)\|^{2}\bigr{]}}$ can be omitted without affecting the inequality and the aiming condition simplifies to:
-$\left\langle x-x_{\ast},\nabla F(x)\right\rangle\geq 0,\qquad\forall\,x.$ (35)
-This is a direct consequence of convexity, which is characterized by (e.g., Nesterov04book, ):
-$\left\langle x-y,\nabla F(x)-\nabla F(y)\right\rangle\geq 0,\qquad\forall\,x,y.$
-To see the implication, simply substituting $y=x_{\ast}$ and $\nabla F(x_{\ast})=0$ into the above inequality gives the full-dimensional aiming condition.
-However, other than the full-dimensional block case, the aiming condition (34) may exhibit significant departure from convexity.
-As an example, consider the case of single coordinate blocks with uniform SiF, i.e., $\rho_{t,k}=\rho_{t,k^{\prime}}$ for all $k,k^{\prime}\in\{1,\ldots,n\}$.
-In light of the equivalent formulation using SiF in (33), the corresponding aiming condition becomes
-$\left\langle x-x_{\ast},\,\textup{sign}(\nabla F(x))\right\rangle\geq 0,\qquad\forall\,x.$ (36)
-Comparing this condition with convexity, it is easy to construct examples that satisfy one of them while failing the other. We provide some simple examples in Appendix B.1.
+Intuitively, the inner product in the aiming condition is predominantly determined by the coordinates that are far from target, in other words, the coordinates with relatively large $|x_{t,k}-x_{*,k}|$ and having a large SiF $\rho_{t,k}$ (effectively high signal-to-noise ratio).
+In general we cannot check the aiming condition analytically, since it involves conditional expectations and an unknown target point $x_{*}$.
+However, we can check empirically if the inner product $\langle x_{t}-x_{*},d_{t}/\sqrt{v_{t}}+\lambda x_{t}\rangle$ is positive while running any particular algorithm.
+In order to do so, we approximate $x_{*}$ with the last iterate after running the algorithm for many iterations.
+Figure 6 and 7 depict the empirical aiming condition while running SGD, SGD with momentum, BCOS-c and BCOSW-c on GPT2-124M with hyper-parameters as chosen in Section 4.
+Specifically, we record the cosines between $x_{t}-x_{\ast}$ and various directions including stochastic gradients $g_{t}$, momentum $m_{t}$, $m_{t}/\sqrt{v_{t}}$ and $m_{t}/\sqrt{v_{t}}+\lambda x_{t}$ along the training trajectory.
+We notice that the cosines of the angles between $x_{t}-x_{*}$ and the search directions are positive (correct aiming) most of the time, especially in the later stage of training.
+Moreover, the actual update directions in each algorithm exhibit the highest cosine values with $x_{t}-x_{\ast}$ among all these directions.
+The above results give empirical evidence that the aiming condition can be much weaker than convexity and often holds in training deep learning models.
+We explain the connection between aiming and convexity in more detail in Appendix A.1.
+Figure 6: Empirical aiming conditions on GPT2. Left: SGD. Right: SGD with momentum.
+Figure 7: Empirical aiming conditions on GPT2. Left: BCOS-c. Right: BCOSW-c.
 5.2 Analysis of conceptual BCOS
 In this section, we provide convergence analysis for the conceptual BCOS method in (30). Our first result is about its one-step contraction property.
-Lemma 5.1.
-Suppose Assumption A holds, $\alpha_{t}\geq 0$ and $\alpha_{t}\lambda<1$ for all $t\geq 0$. Then the sequence $\{x_{t}\}$ generated by (29) satisfies, for all $t\geq 0$,
+Lemma 1
+Suppose Assumption 1 holds, $\alpha_{t}\geq 0$ and $\alpha_{t}\lambda<1$ for all $t\geq 0$. Then the sequence $\{x_{t}\}$ generated by (30) satisfies, for all $t\geq 0$,
 $\mathbf{E}_{t}\bigl[\left\|x_{t+1}-x_{\ast}\right\|^{2}\bigr]\leq(1-\alpha_{t}\lambda)^{2}\left\|x_{t}-x_{\ast}\right\|^{2}+\alpha_{t}^{2}B_{*},$ (35)
@@
 $B_{\ast}=n+\lambda^{2}\left\|x_{\ast}\right\|^{2}+2\lambda\left\|x_{\ast}\right\|_{1}.$ (36)
-Proof.
-First, we notice that the aiming condition (32) is equivalent to
+Proof First, we notice that the aiming condition (33) is equivalent to
 $\biggl\langle x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\lambda x_{*}\biggr\rangle\geq 0.$ (37)
@@
 Plugging the above equality and inequalities into (38) yields the desired result.
-∎
-As a direct consequence of Lemma 5.1, there exists sufficiently small $\alpha_{t}>0$ such that
+As a direct consequence of Lemma 1, there exists sufficiently small $\alpha_{t}>0$ such that
 $\mathbf{E}_{t}\bigl[\left\|x_{t+1}\!-\!x_{\ast}\right\|^{2}\bigr]\leq\|x_{t}\!-\!x_{*}\|^{2},\qquad\forall\,t\geq 0.$
@@
 5.2.1 Almost sure convergence
-The following theorem establishes almost sure convergence of the conceptual BCOS algorithm under a classical condition on the sequence $\{\alpha_{t}\}$.
-Theorem 5.1 (Almost sure convergence).
+The following theorem establishes almost sure (a.s.) convergence of the conceptual BCOS algorithm under a classical condition on the sequence $\{\alpha_{t}\}$.
+Theorem 2 (Almost sure convergence)
 Suppose the stepsize schedule $\{\alpha_{t}\}_{t\geq 0}$ satisfies
 $\alpha_{t}\geq 0,\quad\alpha_{t}\lambda\leq 1,\quad\forall\,t\geq 0,\qquad\sum_{t=0}^{\infty}\alpha_{t}=\infty,\qquad\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty.$ (39)
-The under Assumption A, the sequence $\{x_{t}\}$ generated by (29) satisfies $\left\|x_{t}-x_{\ast}\right\|\to 0$ a.s.
-The proof of Theorem 5.1 is based on the following almost supermartingale lemma of Robbins and Siegmund ROBBINS1971 .
-Lemma 5.2 (“Almost supermartingale”, (ROBBINS1971, , Theorem 1)).
-) Let $(\Omega,\mathcal{F},P)$ be a probability space, and $\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\ldots$ be a sequence of sub-$\sigma$-algebras of $\mathcal{F}$.
+The under Assumption 1, the sequence $\{x_{t}\}$ generated by (30) satisfies $\left\|x_{t}-x_{\ast}\right\|\to 0$ a.s.
+The proof of Theorem 2 is based on the following almost supermartingale lemma of ROBBINS1971.
+Lemma 3 (ROBBINS1971)
+Let $(\Omega,\mathcal{F},P)$ be a probability space, and $\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\ldots$ be a sequence of sub-$\sigma$-algebras of $\mathcal{F}$.
 For each $t\geq 0$, let $X_{t}$, $a_{t}$, $b_{t}$ and $c_{t}$ be non-negative $\mathcal{F}_{t}$-measurable random variables such that
@@
 If $\sum_{t=0}^{\infty}a_{t}<\infty$ and $\sum_{t=0}^{\infty}b_{t}<\infty$, then $\lim_{t\to\infty}X_{t}$ exists and is finite, and $\sum_{t=0}^{\infty}c_{t}<\infty$ almost surely (a.s.).
-Proof of Theorem 5.1.
-Define $X_{t}:=\left\|x_{t}-x_{\ast}\right\|^{2}$ and $\mathcal{F}_{t}$ to be the $\sigma$-algebra generated by $X_{0},\cdots,X_{t}$. Lemma 5.1 implies the following recursion, in the form of (42),
+Proof [Proof of Theorem 2] Define $X_{t}:=\left\|x_{t}-x_{\ast}\right\|^{2}$ and $\mathcal{F}_{t}$ to be the $\sigma$-algebra generated by $X_{0},\cdots,X_{t}$.
+Lemma 1 implies the following recursion, in the form of (40),
 $\mathbf{E}[X_{t+1}|\mathcal{F}_{t}]$ $\leq(1-\alpha_{t}\lambda)^{2}X_{t}+\alpha_{t}^{2}B_{\ast}$
@@
 $\sum_{t=0}^{\infty}a_{t}=\sum_{t=0}^{\infty}\alpha_{t}^{2}\lambda^{2}<\infty,\qquad\sum_{t=0}^{\infty}b_{t}=\sum_{t=0}^{\infty}\alpha_{t}^{2}B_{\ast}<\infty.$
-Therefore, all the assumptions in Lemma 5.2 are satisfied. We may now apply the lemma to conclude that $X_{t}=\left\|x_{t}-x_{\ast}\right\|^{2}\to X$ for some $X<\infty$ and
-[…]
-This is compatible with the assumption $\sum_{t=0}^{\infty}\alpha_{t}=\infty$ only if $\left\|x_{t}-x_{*}\right\|^{2}\to 0$ a.s. ∎
+Therefore, all the assumptions in Lemma 3 are satisfied. We may now apply the lemma to conclude that $X_{t}=\left\|x_{t}-x_{\ast}\right\|^{2}\to X$ for some $X<\infty$ and
+$\sum_{t=0}^{\infty}c_{t}=\sum_{t=0}^{\infty}2\alpha_{t}\lambda\left\|x_{t}-x_{\ast}\right\|^{2}<\infty\quad\text{ a.s.}$
+This is compatible with the assumption $\sum_{t=0}^{\infty}\alpha_{t}=\infty$ only if $\left\|x_{t}-x_{*}\right\|^{2}\to 0$ a.s.
 5.2.2 Rates of convergence
 First, with a constant stepsize schedule, we have linear convergence to a neighborhood of $x_{*}$ as stated by the following corollary.
-Corollary 5.1.
-Consider the conceptual BCOS method (29) and fix the stepsize schedule $\alpha_{t}=\alpha$ where $\alpha$ satisfies $\alpha\lambda<1$. Then under Assumption A, we have
+Corollary 4
+Consider the conceptual BCOS method (30) with a constant step size schedule $\alpha_{t}=\alpha$ where $\alpha$ satisfies $\alpha\lambda<1$. Then under Assumption 1, we have
 $\mathbf{E}\bigl[\left\|x_{t}-x_{\ast}\right\|^{2}\bigr]\leq(1-\alpha\lambda)^{2t}\,\mathbf{E}\bigl[\left\|x_{0}-x_{\ast}\right\|^{2}\bigr]+\frac{\alpha^{2}B_{\ast}}{1-(1-\alpha\lambda)^{2}},$
 where $B_{\ast}$ is defined in (36).
-Proof.
-Recursive application of Lemma 5.1 with $\alpha_{t}=\alpha$ yields
+Proof Recursive application of Lemma 1 with $\alpha_{t}=\alpha$ and taking total expectation yields
 $\mathbf{E}\bigl[\left\|x_{t}-x_{\ast}\right\|^{2}\bigr]$ $\leq(1-\alpha\lambda)^{2t}\,\mathbf{E}\bigl[\left\|x_{0}-x_{\ast}\right\|^{2}\bigr]+\sum_{s=0}^{t-1}(1-\alpha\lambda)^{2(t-s-1)}\alpha^{2}B_{\ast}$
@@
 $\leq(1-\alpha\lambda)^{2t}\,\mathbf{E}\left[\left\|x_{0}-x_{\ast}\right\|^{2}\right]+\frac{\alpha^{2}B_{\ast}}{1-(1-\alpha\lambda)^{2}},$
-which is the desired result. ∎
+which is the desired result.
 With a diminishing stepsize schedule $\alpha_{t}$, we can show that $\mathbf{E}[\|x_{t}-x_{*}\|^{2}]\to 0$ at a sublinear rate, as given in the following theorem.
-Theorem 5.2.
+Theorem 5
 Consider the algorithm (30) with stepsize schedule $\alpha_{t}=\frac{\alpha}{t+1}$.
-Suppose the constant $\alpha$ satisfies $11/2<\alpha\lambda<$ 1 and Assumption A holds.
+Suppose the constant $\alpha$ satisfies $1/2<\alpha\lambda<1$ and Assumption 1 holds.
 Then we have for all $t\geq 0$,
@@
 where $B_{\ast}$ is defined in (36).
-Our proof of Theorem 5.2 is based on Lemma 5.1 and a classical result of Chung chung1954 .
-Lemma 5.3 (Chung’s lemma, (chung1954, , Lemma 1)).
+Our proof of Theorem 5 is based on Lemma 1 and a classical result of chung1954.
+Lemma 6 (chung1954)
 Suppose that $\{X_{t}\}$ is a sequence of real numbers such that for all $t\geq 1$,
@@
 where the last term is bounded if $\sum_{t=1}^{\infty}\alpha_{t}^{2}<\infty$.
-Proof of Theorem 5.2.
-Taking total expectation on both sides of (37), we obtain
+Proof [Proof of Theorem 5] Taking total expectation on both sides of (35), we obtain
 $\mathbf{E}\bigl[\left\|x_{t+1}-x_{\ast}\right\|^{2}\bigr]$ $\leq(1-\alpha_{t}\lambda)^{2}\mathbf{E}\bigl[\left\|x_{t}-x_{\ast}\right\|^{2}\bigr]+\alpha_{t}^{2}B_{\ast}$
@@
 $X_{t+2}\leq\left(1-\frac{a}{t+1}\right)X_{t+1}+\frac{b}{(t+1)^{p+1}},$
-which is in the form of (43) except with the index shift $t\to t+1$. Applying Lemma 5.3 gives the desired result. ∎
-Theorem 5.2 requires $\alpha\lambda>1/2$ to obtain the $\mathcal{O}(1/t)$ rate of convergence.
+which is in the form of (41) except with the index shift $t\to t+1$. Applying Lemma 6 gives the desired result.
+Theorem 5 requires $\alpha\lambda>1/2$ to obtain the $\mathcal{O}(1/t)$ rate of convergence.
 We can remove this condition by working with the stepsize schedule $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ with $p\in(\frac{1}{2},1)$.
 In this case we have an asymptotic $\mathcal{O}(1/t^{p})$ rate, as stated in the following theorem.
-Theorem 5.3.
+Theorem 7
 Consider the algorithm (30) with stepsize schedule $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ where $p\in\left(\frac{1}{2},1\right)$ and the constant $\alpha$ satisfies $\alpha\lambda<1$.
-Then under Assumption A, we have
+Then under Assumption 1, we have
 $\limsup_{t\to\infty}\,(t+1)^{p}\,\mathbf{E}\bigl[\left\|x_{t}-x_{\ast}\right\|^{2}\bigr]\leq\frac{\alpha\left(\lambda^{2}\mathbf{E}\bigl[\left\|x_{0}-x_{\ast}\right\|^{2}\bigr]+(1+K_{p})B_{\ast}\right)}{2\lambda},$ (44)
 where $K_{p}=\sum_{t=1}^{\infty}\frac{1}{(t+1)^{2p}}$ is finite for $p\in(\frac{1}{2},1)$.
-The proof is based on another lemma of Chung chung1954 , restated below.
-Lemma 5.4 ((chung1954, , Lemma 4)).
+The proof is based on another lemma of chung1954, restated below.
+Lemma 8 (chung1954)
 Suppose that $\{X_{t}\}$ is a sequence of real numbers such that for all $t\geq 1$,
 $X_{t+1}\leq\left(1-\frac{a_{t}}{t^{p}}\right)X_{t}+\frac{b}{t^{q}},$ (45)
-where $00<p<1,p<q,a_{t}\geq a>0,b>$ 0. Then
+where $0<p<1,p<q,a_{t}\geq a>0,b>0$. Then
 $\limsup_{t\to\infty}\,t^{q-p}X_{t}\leq\frac{b}{a}.$
-Proof of Theorem 5.3.
-Substituting $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ into (44), we obtain
+Proof [Proof of Theorem 7] Substituting $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ into (42), we obtain
 $\mathbf{E}\bigl[\|x_{t}-x_{*}\|^{2}\bigr]\leq\mathbf{E}\bigl[\|x_{0}-x_{*}\|^{2}\bigr]+B_{\ast}\sum_{s=0}^{\infty}\frac{\alpha^{2}}{(t+1)^{2p}}=\mathbf{E}\bigl[\|x_{0}-x_{*}\|^{2}\bigr]+K_{p}\alpha^{2}B_{\ast}.$
@@
 $\leq\left(1-\frac{2\alpha\lambda}{(t+1)^{p}}\right)\mathbf{E}[\left\|x_{t}-x_{\ast}\right\|^{2}]+\frac{\alpha^{2}\left(\lambda^{2}\mathbf{E}\bigl[\left\|x_{0}-x_{\ast}\right\|^{2}\bigr]+(1+K_{p})B_{\ast}\right)}{(t+1)^{2p}}.$
-The desired result follows by applying Lemma 5.4 with the definitions $X_{t+1}=\mathbf{E}[\left\|x_{t}-x_{\ast}\right\|^{2}]$, $a_{t}=a=2\alpha\lambda$, $b=\alpha^{2}\left(\lambda^{2}\mathbf{E}[\left\|x_{0}-x_{\ast}\right\|^{2}]+(1+K_{p})B_{\ast}\right)$ and $q=2p$.
-∎
-5.2.3 Analysis for not using decoupled weight decay
-Here we briefly explain the convergence analysis of BCOS in the case of $\lambda=0$, or in the case of $\lambda>0$ but without using decoupled weight decay. In other words, we consider
+The desired result follows by applying Lemma 8 with the definitions $X_{t+1}=\mathbf{E}[\left\|x_{t}-x_{\ast}\right\|^{2}]$, $a_{t}=a=2\alpha\lambda$, $b=\alpha^{2}\left(\lambda^{2}\mathbf{E}[\left\|x_{0}-x_{\ast}\right\|^{2}]+(1+K_{p})B_{\ast}\right)$ and $q=2p$.
+5.2.3 Convergence analysis with $\lambda=0$
+Here we briefly explain the convergence analysis of BCOS in the case of $\lambda=0$. In other words, we consider
 $x_{t+1}=x_{t}-\widetilde{\gamma}_{t}\odot d_{t}=x_{t}-\alpha_{t}\frac{d_{t}}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}.$ (46)
@@
   1. 1.
-    First, with $\lambda=0$, we have the following inequality in place of (37) (see Lemma 5.1):
+    First, with $\lambda=0$, we have the following inequality in place of (35) (see Lemma 1):
     $\mathbf{E}_{t}\bigl[\|x_{t+1}-x_{*}\|^{2}\bigr]\leq\|x_{t}-x_{*}\|^{2}-\alpha_{t}\biggl\langle x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\biggr\rangle+\alpha_{t}^{2}n.$
@@
   2. 2.
-    Almost sure convergence as in Theorem 5.4 requires a slightly stronger condition
+    Almost sure convergence as in Theorem 13 requires a slightly stronger condition
     $\biggl\langle x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\biggr\rangle\geq\phi(\|x_{t}-x_{*}\|),$
@@
   3. 3.
-    To obtain sublinear rates of convergence as in Theorems 5.2 and 5.3, we need an even stronger aiming condition, specifically,
+    To obtain sublinear rates of convergence as in Theorems 5 and 7, we need an even stronger aiming condition, specifically,
     $\biggl\langle x_{t}-x_{*},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\biggr\rangle\geq\mu\|x_{t}-x_{*}\|^{2},$ (48)
     with some $\mu>0$ for all $t\geq 0$.
-Finally, we comment on the case of adding the regularization $(\lambda/2)\|x\|^{2}$ with $\lambda>0$ but without using decoupled weight-decay.
-Specifically, we compute $g_{t}=\nabla f(x_{t},\xi_{t})+\lambda x_{t}$ with $\lambda>0$ and apply the BCOS method (48) with $d_{t}$ being $g_{t}$ or its EMA.
-It is worth emphasizing that, in this case, the aiming condition (50) does not automatically hold with $\mu=\lambda$.
-This has to do with the coordinate-wise scaling of $d_{t}$ by $\sqrt{\mathbf{E}_{t}[d_{t}]}$.
-To see this, we consider the loss function $F(x)$ defined in (27), i.e.,
-$F(x):=\mathbf{E}_{\xi}[f(x,\xi)]+\frac{\lambda}{2}\|x\|^{2},$
-and assume that $\mathbf{E}_{\xi}[f(x,\xi)]$ is convex. We contrast the two cases discussed in Section 5.1.1.
-  • With full coordinate block, it is straightforward to show that under the convex assumption, the aiming condition (35) is strengthened to
-    $\langle x-x_{*},\,\nabla F(x)\rangle\geq\lambda\|x-x_{*}\|^{2}.$
-  • However, the above argument does not extend to the case with coordinate-wise scaling.
-As an example, consider the special case of having uniform SiF across all coordinates.
-In this case, the corresponding aiming condition (36) cannot be strengthened to
-    $\langle x-x_{*},\,\textup{sign}(\nabla F(x))\rangle\geq\lambda\|x-x_{*}\|^{2},$
-    because the $\textup{sign}(\cdot)$ function is indifferent or insensitive to the curvature brought by the regularization $(\lambda/2)\|x\|^{2}$.
-In general, with coordinate-wise scaling and nonunifom SiF, the condition (50) may hold with some positive $\mu$ that is much smaller than $\lambda$.
-The above discussion gives a plausible explanation to the empirical observation that decoupled weight decay works better for methods with coordinate-wise scaling (e.g., see Figure 3).
 5.3 Analysis for practical BCOS
@@
 We make the following assumption on the bias of $v_{t}$.
-Assumption B (Bias of second-moment estimator).
+Assumption 2 (Bias of second-moment estimator)
 There exists $\tau\in(0,1)$ and $\epsilon>0$ such that for all $t\geq 0$,
@@
 As an extreme case, by setting $\epsilon\geq\|\mathbf{E}_{t}[d_{t}^{2}]\|_{\infty}$ for all $t$, the trivial estimator $v_{t}=0$ satisfies (51) for any $\tau\in(0,1)$.
-5.3.1 Bounding the difference between conceptual and practical BCOS updates
-The following lemma derives an upper bound on the first term on the right-hand side of (52). The proof is given in Appendix B.2.
-Lemma 5.5.
-Under Assumption B, it holds that:
+5.3.1 Bounding the difference between conceptual and practical updates
+The following lemma derives an upper bound on the first term on the right-hand side of (50). The proof is given in Appendix A.2.
+Lemma 9
+Under Assumption 2, it holds that:
 $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}-\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]+\epsilon}}\right|\leq\frac{\tau+\mathcal{O}(\tau^{2})}{2}\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|+\mathcal{O}(\epsilon).$ (52)
-To bound the second term on the right-hand side of (52), we need the following lemma, whose proof is given in Appendix B.3.
-Lemma 5.6.
+To bound the second term on the right-hand side of (50), we need the following lemma, whose proof is given in Appendix A.3.
+Lemma 10
 Let $Y,Z$ be two random variables and $Z>0$ almost surely, then
@@
 for $p\geq 3$.
-Applying Lemma 5.6 with $Y=d_{t}$ and $Z=v_{t}+\epsilon$, we get the coordinate-wise inequality
+Applying Lemma 10 with $Y=d_{t}$ and $Z=v_{t}+\epsilon$, we get the coordinate-wise inequality
 $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]\!+\!\epsilon}}-\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}\!+\!\epsilon}}\right]\right|\leq\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]\!+\!\epsilon}}\right|\odot\left|\frac{-\textup{Cov}_{t}(d_{t},v_{t}\!+\!\epsilon)}{2\mathbf{E}_{t}[d_{t}]\mathbf{E}_{t}[v_{t}\!+\!\epsilon]}+\mathcal{O}\!\left(\frac{\textup{Var}_{t}(v_{t}\!+\!\epsilon)}{\mathbf{E}_{t}[v_{t}\!+\!\epsilon]^{2}}\right)\right|.$ (54)
@@
 $\mathrm{SNR}_{t}(d_{t})=\frac{\mathbf{E}_{t}[d_{t}]^{2}}{\textup{Var}_{t}(d_{t})},\qquad\mathrm{SNR}_{t}(v_{t}+\epsilon)=\frac{\mathbf{E}_{t}[v_{t}+\epsilon]^{2}}{\textup{Var}_{t}(v_{t}+\epsilon)}=\frac{\mathbf{E}_{t}[v_{t}+\epsilon]^{2}}{\textup{Var}_{t}(v_{t})}.$
-With the definition of SNRs, the term containing $\textup{Cov}_{t}(d_{t},v_{t}+\epsilon)$ in (56) can be expressed as
+With the definition of SNRs, the covariance term in (54) can be expressed as
 $\frac{\textup{Cov}_{t}(d_{t},v_{t}+\epsilon)}{\mathbf{E}_{t}[d_{t}]\mathbf{E}_{t}[v_{t}+\epsilon]}$ $=\frac{\textup{Cov}_{t}(d_{t},v_{t})}{\mathbf{E}_{t}[d_{t}]\mathbf{E}_{t}[v_{t}+\epsilon]}$
@@
 We are ready to bound the difference between the expected updates of the practical BCOS and the conceptual BCOS methods.
-Specifically, bounding the first term on the right-hand side of (52) using Lemma 5.5 and the second term using the last inequality above, we obtain
+Specifically, bounding the first term on the right-hand side of (50) using Lemma 9 and the second term using the last inequality above, we obtain
 $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}-\mathbf{E}_{t}\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]\right|$ $\leq\frac{\tau+\mathcal{O}(\tau^{2})}{2}\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|+\mathcal{O}(\epsilon)$
-$\quad+\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{v_{t}+\epsilon}}\right|\odot\left|\frac{-(1/2)\textup{Corr}_{t}(d_{t},v_{t})}{\sqrt{\mathrm{SNR}_{t}(d_{t})\mathrm{SNR}_{t}(v_{t}+\epsilon)}}+\mathcal{O}\!\left(\frac{1}{\mathrm{SNR}_{t}(v_{t}+\epsilon)}\right)\right|.$
-We can further merge the two terms on the right-hand side by bounding $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{v_{t}+\epsilon}}\right|$ in terms of $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|$.
-To this end, we use Lemma 5.5 to obtain
+$\quad+\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]+\epsilon}}\right|\odot\left|\frac{-(1/2)\textup{Corr}_{t}(d_{t},v_{t})}{\sqrt{\mathrm{SNR}_{t}(d_{t})\mathrm{SNR}_{t}(v_{t}+\epsilon)}}+\mathcal{O}\!\left(\frac{1}{\mathrm{SNR}_{t}(v_{t}+\epsilon)}\right)\right|.$
+We can further merge the two terms on the right-hand side by bounding $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]+\epsilon}}\right|$ in terms of $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|$.
+To this end, we use Lemma 9 to obtain
 $\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[v_{t}]+\epsilon}}\right|\leq\left(1+\frac{\tau+\mathcal{O}(\tau^{2})}{2}\right)\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|+\mathcal{O}(\epsilon).$
 Combining the two inequalities above, we arrive at the following result.
-Lemma 5.7.
-Under Assumptions B, we have for all $t\geq 0$,
+Lemma 11
+Under Assumptions 2, we have for all $t\geq 0$,
 $\left|\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]-\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|$ $\leq\sigma_{t}\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|+\mathcal{O}(\epsilon),$
@@
 The following lemma is a variant of the aiming condition with the practical BCOS update.
-Lemma 5.8.
-Under Assumptions B, it holds that
+Lemma 12
+Under Assumptions 2, it holds that
 $\left\langle x_{t}-x_{\ast},\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]\right\rangle\geq$ $\left\langle x_{t}-x_{\ast},\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\rangle-\left\|x_{t}-x_{\ast}\right\|_{1}\bigl(\sigma_{t}+\mathcal{O}(\epsilon)\bigr),$ (56)
 where $\sigma_{t}$ is given by (55).
-Proof.
-Adding and subtracting the term $\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}$ from the inner product, we obtain:
+Proof Adding and subtracting the term $\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}$ from the inner product, we obtain:
 $\left\langle x_{t}-x_{\ast},\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]\right\rangle$
-[…]
-[…]
-[…]
-where the first inequality is due to Hölder’s inequality and the last inequality is due to Lemma 5.7. In addition, we notice that
+$=$~ $\left\langle x_{t}-x_{\ast},\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\rangle+\left\langle x_{t}-x_{\ast},\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]-\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\rangle$
+$\geq$~ $\left\langle x_{t}-x_{\ast},\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\rangle-\left\|x_{t}-x_{\ast}\right\|_{1}\cdot\left\|\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]-\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\|_{\infty}$
+$\geq$~ $\left\langle x_{t}-x_{\ast},\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\rangle-\left\|x_{t}-x_{\ast}\right\|_{1}\left(\sigma_{t}\left\|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right\|_{\infty}+\mathcal{O}(\epsilon)\right),$
+where the first inequality is due to Hölder’s inequality and the last inequality is due to Lemma 11. In addition, we notice that
 $\biggl\|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\biggr\|_{\infty}=\left\|\sqrt{\rho_{t}}\right\|_{\infty}\leq 1,$ (57)
-which concludes the proof. ∎
-Theorem 5.4 (Almost sure convergence of BCOSW).
-Consider the practical BCOSW method (28).
-Suppose that Assumptions A and B hold, $d_{t}$ is bounded almost surely, and $\{\alpha_{t}\}$ satisfies (41).
+which concludes the proof.
+Theorem 13 (Almost sure convergence of practical BCOS)
+Consider the practical BCOS method (29).
+Suppose that Assumptions 1 and 2 hold, $d_{t}$ is bounded almost surely, and $\{\alpha_{t}\}$ satisfies (39).
 Then there exists $\delta>0$ such that we have with probability one,
@@
 where $\sigma_{t}$ is given by (55).
-The proof of Theorem 5.4 is based on a classical result of stochastic approximation due to Dvoretzky DermanSacks1959 . Here we present an extension that best fits our purpose.
-Theorem 5.5 (An extension of Dvoretzky’s Theorem).
+The proof of Theorem 13 is based on a classical result of stochastic approximation due to Dvoretzky1956 and its extension by DermanSacks1959. Here we present a version that best fits our purpose.
+Theorem 14 (An extension of Dvoretzky’s Theorem)
 Let $(\Omega=\{\omega\},\mathcal{F},P)$ be a probability space. Let $\{x_{t}\}$ and $\{y_{t}\}$ be sequences of random variables such that, for all $t\geq 0$,
@@
 Remark.
-There are many extensions of Dvoretzky’s original results Dvoretzky1956 . Theorem 5.5 is a minor variation of Venter (Venter1966, , Theorem 1). More concretely,
-  • Theorem 1 of Venter Venter1966 has the sequence $\{a_{t}\}$ being a constant sequence, i.e., $a_{t}=a_{\infty}$ for all $t\geq 0$.
-The extension to a non-constant sequence $\{a_{t}\}$ is outlined in the original work of Dvoretzky Dvoretzky1956 and admits a simple proof due to Derman and Sacks DermanSacks1959 .
-  • Theorem 1 of Venter Venter1966 does not include the sequence $\{h_{t}\}$.
-The extension with $\sum_{t=0}^{\infty}h_{t}<\infty$ is straightforward based on a simple argument of Dvoretzky Dvoretzky1956 .
-  • More generally, the sequences $\{a_{t}\}$, $\{b_{t}\}$, $\{c_{t}\}$, $\{h_{t}\}$ can be non-negative measurable functions of $x_{0},\ldots,x_{t}$, and the conclusion of Theorem 5.5 holds if $a_{\infty}$ is an upper bound on $\limsup_{t\to\infty}a_{t}(x_{0},\ldots,x_{t})$ uniformly for all sequences $\{x_{t}\}$ DermanSacks1959 ; ROBBINS1971 .
-Proof of Theorem 5.5.
-We first write the practical BCOSW algorithm in the form of (62):
+There are many extensions of the original result by Dvoretzky1956. Theorem 14 is a minor variation of Venter1966. More concretely,
+  • Theorem 1 of Venter1966 has the sequence $\{a_{t}\}$ being a constant sequence, i.e., $a_{t}=a_{\infty}$ for all $t\geq 0$.
+The extension to a non-constant sequence $\{a_{t}\}$ is outlined in the original work of Dvoretzky1956 and admits a simple proof due to DermanSacks1959.
+  • Theorem 1 of Venter1966 does not include the sequence $\{h_{t}\}$. The extension with $\sum_{t=0}^{\infty}h_{t}<\infty$ is straightforward based on a simple argument of Dvoretzky1956.
+  • More generally, the sequences $\{a_{t}\}$, $\{b_{t}\}$, $\{c_{t}\}$, $\{h_{t}\}$ can be non-negative measurable functions of $x_{0},\ldots,x_{t}$, and the conclusion of Theorem 14 holds if $a_{\infty}$ is an upper bound on $\limsup_{t\to\infty}a_{t}(x_{0},\ldots,x_{t})$ uniformly for all sequences $\{x_{t}\}$ (DermanSacks1959; ROBBINS1971).
+Proof [Proof of Theorem 14] We first write the practical BCOS method in the form of (60):
 $x_{t+1}$ $=(1-\alpha_{t}\lambda)x_{t}-\alpha_{t}\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}$
@@
 $\quad+\alpha_{t}^{2}\left\|\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]+\lambda x_{*}\right\|^{2}.$
-Using Lemma 5.8 and then the aiming condition (39), we obtain
+Using Lemma 12 and then the aiming condition (37), we obtain
 $\left\langle x_{t}-x_{\ast},\,\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]\!+\!\lambda x_{*}\right\rangle$ $\geq\left\langle x_{t}-x_{\ast},\,\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}+\!\lambda x_{*}\right\rangle-\|x_{t}-x_{*}\|_{1}\left(\sigma_{t}+\mathcal{O}(\epsilon)\right)$
@@
 $\left\|\mathbf{E}_{t}\!\left[\frac{d_{t}}{\sqrt{v_{t}+\epsilon}}\right]+\lambda x_{*}\right\|^{2}\leq B,\qquad\forall\,t\geq 0.$
-Together with $10<1-\alpha_{t}\lambda<$ 1, we conclude that
+Together with $0<1-\alpha_{t}\lambda<1$, we conclude that
 $\bigl\|T_{t}(x_{0},\ldots,x_{t})-x_{*}\bigr\|^{2}$ $\leq(1-\alpha_{t}\lambda)^{2}\left\|x_{t}-x_{\ast}\right\|^{2}+2\alpha_{t}(1-\alpha_{t}\lambda)\|x_{t}-x_{*}\|_{1}(\sigma_{t}+\mathcal{O}(\epsilon))+\alpha_{t}^{2}B$
@@
 we arrive at the key inequality (61).
-In order to apply Theorem 5.5, we are left to check the conditions in (64).
+In order to apply Theorem 14, we are left to check the conditions in (62).
 Using the assumption $\alpha_{t}\to 0$, the definition of $a_{t}$ in (65) implies that $\lim_{t\to\infty}a_{t}=\delta^{2}$.
@@
 For $\{c_{t}\}$, if $\sum_{t=0}^{\infty}c_{t}<\infty$, then we must have $\|x_{t}-x_{\star}\|^{2}\to 0$ almost surely because of the assumption $\sum_{t=0}^{\infty}\alpha_{t}=\infty$, and the conclusion of the theorem holds trivially.
-Otherwise, $\sum_{t=0}^{\infty}c_{t}=\infty$ allows all the conditions in (64) to hold, so we can invoke Theorem 5.5 to finish the proof.
-∎
-Following the last remark on Dvoretzky’s theorem on page 5.3.2, the radius of the neighborhood $\delta$ can be refined as follows.
+Otherwise, $\sum_{t=0}^{\infty}c_{t}=\infty$ allows all the conditions in (62) to hold, so we can invoke Theorem 14 to finish the proof.
+Following the third remark after Theorem 14, the radius of the neighborhood $\delta$ can be refined as follows.
 Let $\delta_{t}$ be the smallest constant such that $\|x_{t}-x_{*}\|\geq\delta_{t}$ implies
-$2\|x_{t}-x_{*}\|_{1}\bigl{(}c_{t}+\mathcal{O}(\epsilon)\bigr{)}\leq\lambda\|x_{t}-x_{*}\|^{2},$
+$2\|x_{t}-x_{*}\|_{1}\bigl(\sigma_{t}+\mathcal{O}(\epsilon)\bigr)\leq\lambda\|x_{t}-x_{*}\|^{2},$
 where $\sigma_{t}$ is given by (55). Then we can set $\delta=\limsup_{t\to\infty}\delta_{t}$ in (58).
 5.4 Bias and variance trade-off
-Theorem 5.4 states that the sequence $\{x_{t}\}$ generated by BCOSW converges almost surely to a neighborhood of $x_{*}$, and the radius of the neighborhood is determined by the magnitude of $\sigma_{t}$ and $\epsilon$.
+Theorem 13 states that the sequence $\{x_{t}\}$ generated by the practical BCOS method converges almost surely to a neighborhood of $x_{*}$, and the radius of the neighborhood is determined by the magnitude of $\sigma_{t}$ and $\epsilon$.
 Specifically, smaller $\sigma_{t}$ and $\epsilon$ ensure a smaller neighborhood of convergence.
@@
 Apparently, in order to make $\sigma_{t}$ small, it is desirable for the random variable $v_{t}$ (as an estimator of $\mathbf{E}_{t}[d_{t}^{2}]$) to possess the following properties.
-  • Low bias. This means that Assumption B holds with small $\tau$ and $\epsilon$.
+  • Low bias. This means that Assumption 2 holds with small $\tau$ and $\epsilon$.
   • Low variance, which naturally leads to high SNR.
@@
     $\mathrm{Bias}=\left|\mathbf{E}_{t}[v_{t}]-\mathbf{E}_{t}[d_{t}^{2}]\right|=0,\qquad\mathrm{Variance}=\textup{Var}_{t}(v_{t})=\textup{Var}_{t}(d_{t}^{2}).$
-    In this case, Assumption B is satisfied with $\tau=0$ and $\epsilon=0$. Then we have
+    In this case, Assumption 2 is satisfied with $\tau=0$ and $\epsilon=0$. Then we have
     $x_{t+1}~=~x_{t}-\alpha_{t}\frac{d_{t}}{\sqrt{d_{t}^{2}+\epsilon}}~=~x_{t}-\alpha_{t}\textup{sign}(d_{t}),$
     which corresponds to the sign-gradient method (21) or the sign-momentum method (22), depending on the choice of $d_{t}$.
-Among them, sign-momentum is preferred because $\textup{Var}_{t}(m_{t}^{2})$ is usually much smaller than $\textup{Var}_{t}(g_{t}^{2})$.
+Between them, sign-momentum is preferred because $\textup{Var}_{t}(m_{t}^{2})$ is usually much smaller than $\textup{Var}_{t}(g_{t}^{2})$.
   2. 2.
@@
     $\mathrm{Bias}=\left|\mathbf{E}_{t}[v_{t}]-\mathbf{E}_{t}[d_{t}^{2}]\right|=\left|c\mathbf{1}_{n}-\mathbf{E}_{t}[d_{t}^{2}]\right|,\qquad\mathrm{Variance}=\textup{Var}_{t}(v_{t})=0.$
-    In this case, we need large $\tau$ and $\epsilon$ for Assumption B to hold. Correspondingly
+    In this case, we need large $\tau$ and $\epsilon$ for Assumption 2 to hold. Correspondingly
     $x_{t+1}=x_{t}-\alpha_{t}\frac{d_{t}}{\sqrt{c\mathbf{1}_{n}+\epsilon}}=x_{t}-\alpha_{t}^{\prime}d_{t},$
@@
     It is possible to derive a sensible bound on the bias with a smoothness assumption on the loss function, but we do not delve into it here.
-On the other hand, we have a simple expression for its variance (derivation given in Appendix B.4):
+On the other hand, we have a simple expression for its variance (derivation given in Appendix A.4):
     $\textup{Var}_{t}(v_{t})$ $=(1-\beta)^{2}\textup{Var}_{t}(d_{t}^{2}),$
@@
     $\mathrm{Bias}=\left|\mathbf{E}_{t}\biggl[\sum_{k=1}^{t}(1-\beta_{2})\beta_{2}^{t-k}g_{k}^{2}\biggr]-\mathbf{E}_{t}\left[\biggl(\sum_{k=1}^{t}(1-\beta_{1})\beta_{1}^{t-k}g_{k}\biggr)^{\!2}\right]\right|.$
-    Again, deriving a sensible upper bound on the bias can be quite involved and we do not pursue it here. As for the variance, we get (see Appendix B.4)
+    Again, deriving a sensible upper bound on the bias can be quite involved and we do not pursue it here. As for the variance, we get (see Appendix A.4)
     $\textup{Var}_{t}(v_{t})$ $=(1-\beta_{2})^{2}\textup{Var}_{t}\left(g_{t}^{2}\right).$
@@
 The BCOS-c method (Algorithm 3) also uses $m_{t}$ as the search direction.
-Instead of relying on EMA, it uses a conditional estimator for $\mathbf{E}_{t}[m_{t}^{2}]$, i.e., $v_{t}=(1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$, where $\beta$ is the smoothing factor for computing $m_{t}$.
-We have simple expressions for both the bias and the variance:
+Instead of relying on EMA, it uses a conditional estimator for $\mathbf{E}_{t}[m_{t}^{2}]$, i.e., $v_{t}=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}m_{t}+(1-\beta)^{2}g_{t}^{2}$, where $\beta$ is the smoothing factor for computing $m_{t}$.
+The bias and variance can be expressed as:
+    $\mathrm{Bias_{t}}(v_{t})$ $=2\beta^{2}(1-\beta)\left|m_{t-1}\odot\bigl(m_{t-1}-\mathbf{E}_{t}[g_{t}]\bigr)\right|,$
+    $\textup{Var}_{t}(v_{t})$ $=(1-\beta)^{4}\left(4\beta^{2}m_{t-1}^{2}\textup{Var}_{t}(g_{t})+\,\textup{Var}_{t}(g_{t}^{2})+4\beta m_{t-1}(\mathbf{E}_{t}[g_{t}^{3}]-\mathbf{E}_{t}[g_{t}]\mathbf{E}_{t}[g_{t}^{2}])\right).$
+    Notice that $|m_{t-1}-\mathbf{E}_{t}[g_{t}]|$ is the bias of using $m_{t-1}$ to estimate $\mathbf{E}_{t}[d_{t}]$, and the bias of $v_{t}$ can be much smaller than that.
+For the simple alternative $v_{t}=(1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$, we have simple expressions for both the bias and the variance:
     $\mathrm{Bias_{t}}(v_{t})$ $=2\beta(1-\beta)\left|m_{t-1}\odot\bigl(m_{t-1}-\mathbf{E}_{t}[g_{t}]\bigr)\right|,$
     $\textup{Var}_{t}(v_{t})$ $=(1-\beta)^{4}\,\textup{Var}_{t}(g_{t}^{2}).$
-    Notice that $|m_{t-1}-\mathbf{E}_{t}[g_{t}]|$ is the bias of using $m_{t-1}$ to estimate $\mathbf{E}_{t}[d_{t}]$, and the bias of $v_{t}$ can be much smaller than that.
-Meanwhile, the variance of BCOS-c is the same as Adam when setting $\beta_{2}$ in Adam as $\beta_{2}=1-(1-\beta)^{2}$.
+    The variance of the simple alternative is the same as Adam when setting $\beta_{2}$ in Adam as $\beta_{2}=1-(1-\beta)^{2}$.
 In addition to the bias-variance trade-off illustrated above, the expression of $\sigma_{t}$ in (66) also reveals the following tips for algorithm design:
@@
 BCOS is a family of stochastic approximation methods that exploit the flexibility of using different block-coordinate stepsizes.
-Rather than using sophisticated ideas from optimization, such as preconditioning, it is based on the simple idea of minimizing the distance from the next iterate to an optimal point.
+Rather than using sophisticated techniques from optimization such as preconditioning, it is derived from the simple idea of minimizing the distance from the next iterate to a target point.
 While the optimal stepsizes are not computable, we make several simplifications and focus on constructing efficient statistical estimators (for the second moment of the search direction) in determining the stepsizes.
@@
 Our convergence analysis builds upon several classical results from the stochastic approximation literature.
-In particular, we extend the classical aiming condition towards a target point to account for coordinate-wise stepsizes and decoupled weight-decay (interpreted as $L_{2}$-regularization).
+In particular, we extend the classical aiming condition towards a target point to account for coordinate-wise stepsizes and decoupled weight-decay.
 For the conceptual BCOS method, which assumes exact second moment, we establish almost-sure convergence and $\mathcal{O}(1/t)$ rate of convergence.
@@
 Last but not least, we extend our gratitude to Damek Davis, Anil Damle, Adrian Lewis, James Renegar and Katya Scheinberg for their comments and insights that helped us clarify and refine the theoretical underpinnings of the paper.
-Appendix A PyTorch implementation of BCOS
-The following code is also available at https://github.com/facebookresearch/bcos.
-1import torch
-2from torch.optim import Optimizer
-3
-4class BCOS(Optimizer):
-5 def __init__(self, params, lr, beta=0.9, eps=1e-6,
-6 weight_decay=0.1, mode='c', decouple_wd=True):
-7
-8 defaults = dict(lr=lr, beta=beta, eps=eps, wd=weight_decay)
-9 super().__init__(params, defaults)
-10
-11 if mode not in ['g', 'm', 'c']:
-12 raise ValueError(f"BCOS mode {mode} not supported")
-13 self.mode = mode
-14 self.decouple_wd = decouple_wd # True for BCOSW
-15
-16 def step(self, closure = None):
-17
-18 for group in self.param_groups:
-19 lr = group["lr"]
-20 beta = group["beta"]
-21 eps = group["eps"]
-22 wd = group["wd"]
-23
-24 for p in group["params"]:
-25 if not p.requires_grad:
-26 continue
-27
-28 state = self.state[p]
-29 g = p.grad
-30
-31 # initialize optimizer states for specific modes
-32 if self.mode in ['m', 'c'] and 'm' not in state:
-33 state['m'] = g.detach().clone()
-34 if self.mode in ['g', 'm'] and 'v' not in state:
-35 state['v'] = g.detach().square()
-36
-37 # decoupled weight decay or absorb in gradient
-38 if self.decouple_wd: # p := (1 - lr * wd) * p
-39 p.data.mul_(1 - lr * wd)
-40 else: # g := g + wd * p
-41 g.data.add_(p.data, alpha = wd)
-42
-43 if self.mode in ['m', 'c']:
-44 m = state['m']
-45 if self.mode == 'c': # conditional estimator
-46 beta_v = 1 - (1 - beta)**2
-47 g2 = g.detach().square()
-48 v = beta_v * m.square() + (1 - beta_v) * g2
-49 # update momentum
-50 m.mul_(beta).add_(g.detach(), alpha=1 - beta)
-51 d = m
-52 else:
-53 d = g.detach()
-54
-55 if self.mode in ['g', 'm']: # EMA estimator
-56 v = state['v']
-57 v.mul_(beta).add_(d.square(), alpha=1 - beta)
-58
-59 # BCOS update: p := p - lr * (d / (sqrt(v) + eps))
-60 p.data.add_(d.div(v.sqrt() + eps), alpha= - lr)
-Appendix B Appendix
-B.1 Aiming versus convexity
+Appendix A
+A.1 Aiming versus convexity
+Our aiming condition in Assumption 1 does not have an upper bound as in the classical condition (32), meaning that it does not require smoothness.
+Smoothness would give a better bound on the bias and variance of the second-moment estimator of the practical BCOS algorithm, but does not affect stability of the algorithm as long as Assumption 2 holds.
+Here we focus on the connection of the aiming condition with convexity.
+For ease of comparison, we consider the general block structure described in Section 1.3 with $d_{t}=\nabla f(x_{t},\xi_{t})$ and set $\lambda=0$.
+In this case, we have $\mathbf{E}[d_{t}]=\nabla F(x_{t})$ and
+$\sum_{k=1}^{m}\left\langle x_{k}-x_{\ast,k},~\frac{\nabla_{k}F(x)}{\sqrt{\mathbf{E}\bigl[\left\|\nabla_{k}f(x,\xi)\right\|^{2}\bigr]}}\right\rangle\geq 0,\qquad\forall\,x,$ (67)
+where $\nabla_{k}F(x)\in\mathbf{R}^{n_{k}}$ denotes the sub-vector of $\nabla F(x)$ corresponding to the $k$th block.
+In the specific case of a full-dimensional block ($m=1$), the denominator $\sqrt{\mathbf{E}\bigl[\|\nabla f(x,\xi)\|^{2}\bigr]}$ can be omitted without affecting the inequality and the aiming condition simplifies to:
+$\left\langle x-x_{\ast},\nabla F(x)\right\rangle\geq 0,\qquad\forall\,x.$ (68)
+This is a direct consequence of convexity when $x_{\ast}$ is a minimizer of $F$, which is characterized by (e.g., Nesterov04book):
+$\left\langle x-y,\nabla F(x)-\nabla F(y)\right\rangle\geq 0,\qquad\forall\,x,y.$
+To see the implication, simply substituting $y=x_{\ast}$ and $\nabla F(x_{\ast})=0$ into the above inequality gives the full-dimensional aiming condition.
+It is easy to come up with examples that satisfy the aiming condition but not convexity.
+For simplicity, we consider the 1-dimensional case and without loss of generality let $x^{*}=0$.
+Let $g(x):=\mathbf{E}_{\xi}[g(x,\xi)]$ be any continuous function that satisfies $g(x)>0$ for $x>0$ and $g(x)<0$ for $x<0$, which leads to
+$\langle x-x^{*},g(x)\rangle=x\cdot g(x)>0,$
+for all $x\neq 0$. For example, consider
+$g(x)=x(x-3)^{2}(x+2)^{2}+\lambda x,$
+which is a non-monotone polynomial, thus the gradient of a nonconvex polynomial. Yet, this function satisfies the aiming condition (32) for $x^{*}=0$ and for any $\lambda\geq 0$.
+However, other than the full-dimensional block case, the aiming condition (67) is not necessarily implied by convexity.
+As an example, consider the case of single coordinate blocks with uniform SiF, i.e., $\rho_{t,k}=\rho_{t,k^{\prime}}$ for all $k,k^{\prime}\in\{1,\ldots,n\}$.
+In light of the equivalent formulation using SiF in (34), the corresponding aiming condition becomes
+$\left\langle x-x_{\ast},\,\textup{sign}(\nabla F(x))\right\rangle\geq 0,\qquad\forall\,x.$ (69)
+Comparing this condition with convexity, it is easy to construct examples that satisfy one of them while failing the other.
 To illustrate the fundamental differences between the coordinate-wise aiming condition (69) and the standard convexity condition (68), we provide the following two counterexamples, each satisfying one condition while failing the other:
-  • Aiming but not convex: Let $f(x):=\log(x)$ with the optimal solution $x_{\ast}=0$.
+  • Aiming but not convex: Let $f(x):=\log(x)$ with the target solution $x_{\ast}=0$.
 On the domain of $\mathbf{R}_{+}$, the gradient is $f^{\prime}(x)=\frac{1}{x}$, and thus $\textup{sign}(f^{\prime}(x))=1$ for all $x>0$.
@@
     Thus, the aiming condition (69) does not hold at this point even though $f$ is convex.
-B.2 Proof of Lemma 5.5
-Proof.
-The proof leverages the Taylor expansion of $\phi(y):=1/\sqrt{y}$. Specifically, we have
+A.2 Proof of Lemma 9
+Proof The proof leverages the Taylor expansion of $\phi(y):=1/\sqrt{y}$. Specifically, we have
 $\phi(y+\delta)=\sum_{p=0}^{\infty}\frac{\phi^{(p)}(y)}{p!}\delta^{p},$
@@
 $=\frac{\tau+\mathcal{O}(\tau^{2})}{2}\left|\frac{\mathbf{E}_{t}[d_{t}]}{\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}}\right|+\mathcal{O}(\epsilon)$
-where the inequality (69) is a consequence of Assumption B. ∎
-B.3 Proof of Lemma 5.6
+where the inequality (70) is a consequence of Assumption 2.
+A.3 Proof of Lemma 10
 We first derive a useful approximation for a general smooth function.
-Lemma B.1.
+Lemma 15
 Let $\phi:\mathbf{R}^{n}\to\mathbf{R}$ be a smooth ($C^{\infty}$) function and $X\in\mathbf{R}^{n}$ a random variable. Then we have
 $\mathbf{E}[\phi(X)]=\phi(\mathbf{E}[X])+\frac{1}{2}\bigl\langle\nabla^{2}\phi(\mathbf{E}[X]),\textup{Cov}(X)\bigr\rangle+\sum_{|\bm{p}|=3}^{\infty}\!\frac{D^{(\bm{p})}\phi(\mathbf{E}[X])}{\bm{p}!}\mathbf{E}\bigl[(X\!-\!\mathbf{E}[X])^{(\bm{p})}\bigr],$ (71)
-[…]
+where $\langle\cdot,\cdot\rangle$ denotes matrix inner product, i.e, $\langle A,B\rangle=\textup{Tr}(A^{T}B)$.
+For the multi-variable derivatives, we denote $\bm{p}\in\{0,1,2,\ldots\}^{n}$, $|\bm{p}|=p_{1}+\cdots+p_{n}$, $\bm{p}!=p_{1}!\cdots p_{n}!$, and
 $D^{(\bm{p})}\phi$ $=\frac{\partial^{|p|}\phi}{\partial X^{p}}=\frac{\partial^{p_{1}+\cdots+p_{n}}\phi}{\partial X_{1}^{p_{1}}\cdots\partial X_{n}^{p_{n}}},$
 $(X-\mathbf{E}[X])^{(\bm{p})}$ $=(X_{1}-\mathbf{E}[X_{1}])^{p_{1}}\cdots(X_{n}-\mathbf{E}[X_{n}])^{p_{n}}.$
-Proof.
-Let $\delta:=X-\mathbf{E}[X]$ and consider the Taylor expansion of $\phi$ at $\mathbf{E}[X]$:
+Proof Let $\delta:=X-\mathbf{E}[X]$ and consider the Taylor expansion of $\phi$ at $\mathbf{E}[X]$:
 $\phi(X)$ $=\phi(\mathbf{E}[X])+\nabla\phi(\mathbf{E}[X])^{T}\delta+\frac{1}{2}\delta^{T}\nabla^{2}\phi(\mathbf{E}[X])\delta+\sum_{|\bm{p}|=3}^{\infty}\frac{D^{(\bm{p})}\phi(\mathbf{E}[X])}{\bm{p}!}\delta^{(\bm{p})}.$
@@
 $\mathbf{E}\bigl[\nabla\phi(\mathbf{E}[X])^{T}\delta\bigr]=\nabla\phi(\mathbf{E}[X])^{T}\mathbf{E}\bigl[\delta\bigr]=0,$
-which, together with the equation $\textup{Cov}(X)=\mathbf{E}[\delta\delta^{T}]$, yields the desired result. ∎
-Proof of Lemma 5.6.
-We apply Lemma B.1 with $X:=(Y,Z)$ and $\phi(x)=\phi(y,z):=\frac{y}{\sqrt{z}}$. First, the gradient and Hessian of $g$ can be calculated as
+which, together with the equation $\textup{Cov}(X)=\mathbf{E}[\delta\delta^{T}]$, yields the desired result.
+Proof [Proof of Lemma 10] We apply Lemma 15 with $X:=(Y,Z)$ and $\phi(x)=\phi(y,z):=\frac{y}{\sqrt{z}}$. First, the gradient and Hessian of $g$ can be calculated as
+$\nabla\phi(x)=\nabla\phi(y,z)=\begin{pmatrix}\displaystyle\frac{1}{z^{1/2}}\\[8.61108pt] \displaystyle-\frac{y}{2z^{3/2}}\end{pmatrix},\qquad\nabla^{2}\phi(x)=\nabla^{2}\phi(y,z)=\begin{bmatrix}\displaystyle 0,&\displaystyle-\frac{1}{2z^{3/2}}\\[8.61108pt] \displaystyle-\frac{1}{2z^{3/2}},&\displaystyle\frac{3y}{4z^{5/2}}\end{bmatrix}$.
+For general $p$-th partial derivative, we derive the following result for any $q\in\{0,\ldots,p\}$:
 […]
-For general $p$-th partial derivative, we derive the following result for any $q\in\{0,\ldots,p\}$:
-$\frac{\partial^{p}\phi}{\partial y^{q}\partial z^{p-q}}=\frac{\partial^{p-q}}{\partial z^{p-q}}\left(\frac{\partial^{q}\phi}{\partial y^{q}}\right)=\begin{cases}\displaystyle 0\quad&\text{if }q\geq 2,\\[8.61108pt] \displaystyle\frac{\partial^{p-1}}{\partial z^{p-1}}\frac{1}{\sqrt{z}}=(-1)^{p-1}\frac{(2p-2)!}{4^{p-1}(p-1)!}z^{-\frac{2p-1}{2}}\quad&\text{if }q=1,\\[8.61108pt] \displaystyle y\cdot\frac{\partial^{p}}{\partial z^{p}}\frac{1}{\sqrt{z}}=(-1)^{p}\frac{(2p)!}{4^{p}p!}yz^{-\frac{2p+1}{2}}\quad&\text{if }q=0.\\ \end{cases}$
 Substitute the Hessian and $p$-th order partial derivative into (71), we get
 $\mathbf{E}\left[\frac{Y}{\sqrt{Z}}\right]$ $=\frac{\mathbf{E}[Y]}{\sqrt{\mathbf{E}[Z]}}-\mathbf{E}\left[\frac{(Y-\mathbf{E}[Y])(Z-\mathbf{E}[Z])}{2\mathbf{E}[Z]^{3/2}}\right]+\mathbf{E}\left[\frac{3\mathbf{E}[Y](Z-\mathbf{E}[Z])^{2}}{8\mathbf{E}[Z]^{5/2}}\right]$
-$\quad+\sum_{p=3}^{\infty}\frac{1}{p!}\left(\frac{p(2p-2)!}{4^{p-1}(p-1)!}\mathbf{E}\left[\frac{(Y-\mathbf{E}[Y])(Z-\mathbf{E}[Z])^{p-1}}{\mathbf{E}[Z]^{\frac{2p-1}{2}}}\right]+(-1)^{p}\frac{(2p)!}{4^{p}p!}\mathbf{E}\left[\frac{\mathbf{E}[Y](Z-\mathbf{E}[Z])^{p}}{\mathbf{E}[Z]^{\frac{2p+1}{2}}}\right]\right)$
+$\quad+\sum_{p=3}^{\infty}\frac{1}{p!}(-1)^{p-1}\frac{p(2p-2)!}{4^{p-1}(p-1)!}\mathbf{E}\left[\frac{(Y-\mathbf{E}[Y])(Z-\mathbf{E}[Z])^{p-1}}{\mathbf{E}[Z]^{\frac{2p-1}{2}}}\right]$
+$\quad+\sum_{p=3}^{\infty}\frac{1}{p!}(-1)^{p}\frac{(2p)!}{4^{p}p!}\mathbf{E}\left[\frac{\mathbf{E}[Y](Z-\mathbf{E}[Z])^{p}}{\mathbf{E}[Z]^{\frac{2p+1}{2}}}\right]$
 $=\frac{\mathbf{E}[Y]}{\sqrt{\mathbf{E}[Z]}}-\frac{\textup{Cov}(Y,Z)}{2\mathbf{E}[Z]^{3/2}}+\frac{3\mathbf{E}[Y]\textup{Var}(Z)}{8\mathbf{E}[Z]^{5/2}}$
@@
 $\quad+\mathcal{O}\left(\frac{\mathbf{E}\bigl[(Y-\mathbf{E}[Y])(Z-\mathbf{E}[Z])^{p-1}\bigr]}{\mathbf{E}[Y]\mathbf{E}[Z]^{p-1}}\right)+\mathcal{O}\left(\frac{\mathbf{E}\bigl[(Z-\mathbf{E}[Z])^{p}\bigr]}{\mathbf{E}[Z]^{p}}\right)\Biggr),$
-which is the desired result. ∎
-B.4 Bias and variance calculation
+which is the desired result.
+A.4 Bias and variance calculation
 Here we derive the bias and variance expressions for some estimators listed in Section 5.4.
@@
     $=(1-\beta_{2})^{2}\textup{Var}_{t}\left(g_{t}^{2}\right).$
-  • The conditional estimator of BCOS-c. In this case, we have $d_{t}=m_{t}=\beta m_{t-1}+(1-\beta)g_{t}$ and $v_{t}=(1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$. Therefore,
+  • The conditional estimator of BCOS-c.
+In this case, we have $d_{t}=m_{t}=\beta m_{t-1}+(1-\beta)g_{t}$ and $v_{t}=\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}m_{t}+(1-\beta)^{2}g_{t}^{2}$.
+Therefore, we derive the bias and variance as follows
+    $\mathrm{Bias}$ $=|\mathbf{E}_{t}[v_{t}]-\mathbf{E}_{t}[d_{t}^{2}]|$
+    $=\Big|\mathbf{E}_{t}\left[\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}(\beta m_{t-1}+(1-\beta)g_{t})+(1-\beta)^{2}g_{t}^{2}\right]$
+    $\quad-\mathbf{E}_{t}\left[\left(\beta m_{t-1}+(1-\beta)g_{t}\right)^{2}\right]\Big|$
+    $=\Big|\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}(\beta m_{t-1}+(1-\beta)\mathbf{E}_{t}[g_{t}])+(1-\beta)^{2}\mathbf{E}_{t}\left[g_{t}^{2}\right]$
+    $\quad-\beta^{2}m^{2}_{t-1}-2\beta(1-\beta)m_{t-1}\mathbf{E}_{t}\left[g_{t}\right]-(1-\beta)^{2}\mathbf{E}_{t}\left[g_{t}^{2}\right]\Big|$
+    $=\left|2\beta(1-\beta)m_{t-1}(\beta m_{t-1}+(1-\beta)\mathbf{E}_{t}[g_{t}]-\mathbf{E}_{t}[g_{t}])\right|$
+    $=2\beta^{2}(1-\beta)\left|m_{t-1}\left(m_{t-1}-\mathbf{E}_{t}\left[g_{t}\right]\right)\right|,$
+    and
+    $\textup{Var}_{t}(v_{t})$ $=\mathbf{E}_{t}\Big[\Big(\beta^{2}m_{t-1}^{2}+2\beta(1-\beta)m_{t-1}(\beta m_{t-1}+(1-\beta)g_{t})+(1-\beta)^{2}g_{t}^{2}$
+    $\quad-\beta^{2}m_{t-1}^{2}-2\beta(1-\beta)m_{t-1}(\beta m_{t-1}+(1-\beta)\mathbf{E}_{t}[g_{t}])-(1-\beta)^{2}\mathbf{E}_{t}[g_{t}^{2}]\Big)^{2}\Big]$
+    $=\mathbf{E}_{t}\Big[\Big(2\beta(1-\beta)^{2}m_{t-1}(g_{t}-\mathbf{E}_{t}[g_{t}])+(1-\beta)^{2}(g_{t}^{2}-\mathbf{E}_{t}[g_{t}^{2}])\Big)^{2}\Big]$
+    $=4\beta^{2}(1-\beta)^{4}m_{t-1}^{2}\mathbf{E}_{t}[(g_{t}-\mathbf{E}_{t}[g_{t}])^{2}]+4\beta(1-\beta)^{4}m_{t-1}\mathbf{E}_{t}[(g_{t}-\mathbf{E}_{t}[g_{t}])(g_{t}^{2}-\mathbf{E}_{t}[g_{t}^{2}])]$
+    $\quad+(1-\beta)^{4}\mathbf{E}_{t}\left[\left(g_{t}^{2}-\mathbf{E}_{t}[g_{t}^{2}]\right)^{2}\right]$
+    $=4\beta^{2}(1-\beta)^{4}m_{t-1}^{2}\textup{Var}_{t}(g_{t})+4\beta(1-\beta)^{4}m_{t-1}\left(\mathbf{E}_{t}[g_{t}^{3}]-\mathbf{E}_{t}[g_{t}]\mathbf{E}_{t}[g_{t}^{2}]\right)+(1-\beta)^{4}\textup{Var}_{t}\left(g_{t}^{2}\right).$
+    The simple alternative $v_{t}=(1-(1-\beta)^{2})m_{t-1}^{2}+(1-\beta)^{2}g_{t}^{2}$ has bias and variance in simpler form:
     $\mathrm{Bias}$ $=|\mathbf{E}_{t}[v_{t}]-\mathbf{E}_{t}[d_{t}^{2}]|$